### v1

# Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India

Abstract.
Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, holds promise for empowering individuals and transforming healthcare delivery. However, barriers to its adoption remain underexplored in the Indian context. This study investigates PHI adoption among Indian users and stakeholders using a multi-method approach. An awareness survey (n = 87) examined the usage of wearables and general PHI engagement, followed by semi-structured interviews (n = 22) that explored motivations, usage patterns, and health information sources. Qualitative analysis revealed that while PHI is valued for health monitoring and shared/collective care, its adoption is hindered by factors such as low health literacy, usability challenges, and mistrust in digital health platforms. Further stakeholder interviews and co-design workshops informed the development of a Figma-based prototype, which was evaluated for usability. Based on these findings, we offer design recommendations for an integrated, user-controlled PHI platform featuring accessible analytics and verifiable health information. Our insights highlight the socio-technical challenges of PHI adoption in India and underscore the need for reliable, user-centric solutions to support proactive healthcare.
1. Introduction
Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling active self-management. Globally, PHI leverages technologies—from smartphone applications to wearable devices such as smartwatches and fitness trackers—facilitating continuous health monitoring (Shajari et al., 2023). Factors like high health literacy, technology familiarity, and a proactive healthcare ethos drive PHI adoption in the Western context (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024). Emerging evidence (Chmielewski and Meyer, 2024) in the Western context suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care—seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
In India, however, infrastructural, sociocultural, and technological challenges intersect and constrain PHI’s potential. Digital health tools, including wearables, are often perceived as fashion accessories or status symbols rather than integrated components of comprehensive health management (Misra et al., 2023). Moreover, cultural practices, family dynamics, and skepticism towards advanced technologies, such as AI-driven solutions, further influence PHI usage and adoption (Hoque and Bao, 2015; Zhang et al., 2022; Creaser et al., 2022; Patil et al., 2022; Reis et al., 2024; Kapania et al., 2022; Wisetsri and Vijai, 2021). Given the cultural dynamics of the Indian context, a significant gap in its adoption and use can be filled by examining and understanding how PHI is currently integrated and utilized in daily health practices in India and shared with family and loved ones.
To address these challenges, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach. We conducted a survey (n=87) to assess awareness and usage of smart wearables, followed by interviews (n=22) exploring user motivations, usage patterns, and health information sources. Additional interviews with key stakeholder pairs—couples, intergenerational family groups, and healthcare providers—and co-design workshops (n=12) employing card sorting examined shared care practices and data-sharing preferences. Further, we evaluated a figma prototype to understand its usability and get design recommendations from the participants on what, how, and when they would like to share their health information/analytics with different stakeholders.
Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India. Specifically, we ask: (1) How do Indians adopt, integrate, and perceive PHI in their health management? (2) What dictates PHI’s effective utilization and perceived utility, and how can we improve them?
Our findings offer a deeper understanding of PHI challenges and opportunities in India and provide a Proof-of-Concept evaluation to strengthen our recommendations further. Through this work, we lay the groundwork for a more reliable and integrated PHI system supporting personal and shared care management, culturally appropriate to the Indian context and the global south.
2. Related Work
The body of work on personal health informatics within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019). These factors have propelled research toward innovative health management solutions and broadened the adoption of PHI systems. The rapid increase in chronic health conditions and an expanding digital landscape make personal health informatics especially critical in India. This work aims to contextualize the evolution of PHI within the Indian healthcare landscape, addressing its unique challenges and opportunities. This section examines the existing scholarship through two critical lenses: (1) personal health informatics in HCI and digital health tracking and (2) the barriers and enablers of PHI in the Indian context. These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural health practices.
2.1. Personal Health Informatics in HCI
Personal health informatics (PHI) within HCI has evolved from individual tracking (Rooksby et al., 2014) to encompassing broader ecological models (Epstein et al., 2015) that account for social, structural, and cultural factors. Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts. These frameworks have been extensively applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022), emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing.
Wearable technologies and smartphone applications are at the forefront of PHI in various contexts, like sports, managing chronic conditions, and broader long-term health monitoring initiatives (Tholander and Nylander, 2015; Klasnja et al., 2011; Poole, 2013; Nunes et al., 2015; Zegura et al., 2018; Blandford, 2019; Gooch et al., 2021; Fritz et al., 2014). Prior research on wearable use for PHI has also examined why people use and abandon smart devices (Lazar et al., 2015) and how it differs among people with prior self-tracking experience (Rapp and Cena, 2016). Studies on PHI often highlight the continuous capture and analysis of biosignals for precise intervention, particularly in chronic disease management, such as multiple sclerosis, chronic fatigue, and hypertension (Stuart et al., 2022; Jones et al., 2021; Tong et al., 2019).
While proactive use of PHI systems exists for critical use cases, such as emergencies (Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024), fall detections(Delahoz and Labrador, 2014; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022), and medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), we are yet to see them used for a broader PHI in the Indian context.
PHI in HCI is not limited to wearables and trackers; it also uses other modalities to integrate into an individual’s wellbeing. These diverse modalities represent different engagement strategies, from gamification in AR for improving physical activity (Zhao et al., 2024) using virtual pets to fostering social interaction through community displays (Gooch et al., 2021). Beyond physical and social wellbeing, wearables increasingly support mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023). Despite this, the literature (Lee et al., 2022) often emphasizes specialized healthcare needs, such as technology readiness and e-health literacy, while general proactive health tracking remains underexplored in HCI, particularly in India. Our work adopts an ecological perspective to examine intrinsic and extrinsic factors influencing PHI adoption, bridging the gap between research on wearables and broader proactive health management strategies.
2.2. Barriers and Enablers of PHI in the Indian Context
Kabir and Wiese’s meta-synthesis provides a broader understanding of the general global barriers and facilitators for personal informatics systems (Kabir and Wiese, 2023). Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges. Limited digital and health literacy among marginalized populations poses significant challenges. Older adults often face steep learning curves when adopting digital health technologies (Cherian et al., 2024), while children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023). Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement (Thakkar et al., 2022; Okolo et al., 2024; Lee et al., 2024b; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap.
Efforts to improve engagement through public health information datafication have shown promise in enhancing accessibility and comprehension (Huang et al., 2022; Pai and Alathur, 2020). However, implementing these strategies in India demands policy-level changes to ensure equitable access and integration with local health systems. National healthcare initiatives—such as Ayushman Bharat Health Account (ABHA) and expanding telemedicine services—have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023). Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024). However, sustaining long-term engagement remains challenging, as users frequently cite usability issues and information overload as key reasons for abandonment (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024).
Apart from the Individual and Public/Infrastructural challenges, cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023). Collectivist norms and uncertainty avoidance play a significant role in shaping attitudes toward digital health technologies (Karnatak et al., 2023; Zhang et al., 2022). The perceived economic value of health technologies is often underappreciated unless they provide critical utility to users through financial implications, for example, in managing chronic conditions (Singh et al., 2022b). While trust and perception issues persist, privacy concerns, data accuracy doubts, and discomfort with wearables further inhibit adoption (Shajari et al., 2023; Seelam et al., 2024).
While these multi-faceted barriers exist in the Indian context, our study aims to reconfirm these, identify more, and develop strategies to improve health literacy, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization. Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies.
3. Methodology
To examine the factors influencing the adoption of PHI in India, we conducted a study starting in January 2024, following approval from the Institute’s IRB. Our investigation began with a thorough review of the existing literature to understand the PHI usage and adoption landscape. It focused on the barriers and drivers for its adoption in personal health management and how its use (or the lack of it) is integrated into health management practices. We then designed a mixed-methods approach comprising surveys, interviews, and co-design workshops to capture the diverse perspectives of users. After analyzing all participant data and their inputs, we evaluated a rapid figma prototype of sharing preferences to further reinforce a unique finding.
3.1. Participants
The study recruited participants from across India, aiming to capture a diverse demographic range in age, gender, and professional background. To ensure a representative sample (especially those who can purchase and use even affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale by Saleem et al. (Saleem and Jan, 2021). We employed a combination of purposive and snowball sampling methods to identify and engage participants. The sample included a variety of professions, such as IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society. This approach enabled us to gather insights from participants with differing levels of familiarity and engagement with wearable health technologies. Prior to participation, all individuals were provided with a clear explanation of the study’s purpose, and informed consent was obtained. Demographic details for the surveys and interview participants are summarized in Table 1, offering a comprehensive overview of the sample’s composition.
Study Component | Total | Gender | Age | Education | |||||||
Male | Female | Prefer Not | Range | Mean | High | Bachelors | Advanced | ||||
to Say | School | Degrees | |||||||||
PHI awareness survey | 87 | 50 | 36 | 1 | 18-83 | 32 | 6 | 31 | 50 | ||
34 | 20 | 14 | 0 | 18-62 | 32 | 1 | 16 | 17 | ||
interviews | 22 | 16 | 8 | 0 | 18-62 | 28 | 5 | 6 | 11 | ||
co-design workshops | 12 | 6 | 6 | 0 | 27-40 | 33 | 0 | 8 | 4 | ||
114 | 44 | 70 | 0 | 19-75 | 34 | 3 | 57 | 54 |
[The table listing participant statistics from various study components]Table 1 lists the ranges and mean of age, gender, and educational background for the three components of the study.
3.2. Study Design
3.2.1. Survey Design
The initial PHI awareness survey was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants. The survey included closed-ended and open-ended questions designed to capture a broad spectrum of wearable health technology use data. The survey comprised a total of 29 questions, including 8 questions on demographic information and 8 open-ended questions (6 for users and 2 for non-users). The closed-ended questions focused on collecting demographic information (age, gender, occupation), the type of wearable devices used for PHI, frequency of use, and specific health metrics tracked (such as steps, heart rate, and sleep patterns). The open-ended questions focused on aspects of trust and usage, such as (1)Please explain briefly if you have ever changed the health wearable you use and why. (2) What do you understand about the health data/analysis from your wearable? (3) Do you trust the accuracy of the wearable health data? (4) Has the history from your health wearable ever helped you with doctor visits? The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives. These responses were crucial for identifying recurring themes and patterns in the qualitative analysis. The survey remained open until saturation was observed in the responses.
The PHI awareness survey had a total of 99 responses, of which 87 were complete and were used for further analysis. The demographics had a mean age of 32 and a maximum age of 83, which included 50 males, 36 females, and 1 participant who preferred not to disclose their gender. Among the complete responders, there were 55 users and 32 non-users. The majority (85%) of the survey participants were between 18 and 44 years old. The diversity of participants in age, gender, profession, and education provided a robust dataset for analyzing broader adoption/use/behavior patterns and informed the qualitative exploration in the subsequent interviews.
The concluding survey on health-sharing features was designed to assess both awareness and utilization of these features among Android and iOS users. It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each subgroup (Apple users, Apple non-users, Android users, Android non-users). Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions. The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and—when applicable—the frequency of their usage. Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses. The survey was closed upon reaching saturation of responses.
The health sharing feature survey was completed by 114 participants, with a mean age of 34 and a maximum age of 75, which included 44 males and 70 females. Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed. This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms, especially with the Android users (who are % of users in India).
ID | Age | Gender | Education | Wearable Usage | Was Gifted | Abandonment | Health Analytics |
P1 | 19 | Male | High School | Health | Yes | Still Using | Advanced |
P2 | 25 | Female | Advanced Degree | Non-User | N/A | N/A | N/A |
P3 | 33 | Female | Advanced Degree | Health | Yes | Still Using | Advanced |
P4 | 33 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P5 | 29 | Female | Bachelors | Utility | No | Still Using | N/A |
P6 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P7 | 32 | Male | Doctorate | Health/Utility | No | Still Using | Basic |
P8 | 26 | Male | Advanced Degree | Health | No | Still Using | Basic |
P9 | 29 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
P10 | 62 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P11 | 34 | Male | Bachelors | Health/Utility | No | Stopped Using | Advanced |
P12 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
P13 | 29 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Advanced |
P14 | 20 | Male | High School | Non-User | N/A | N/A | N/A |
P15 | 25 | Female | Advanced Degree | Fashion/Health | Yes | Still Using* | Basic |
P16 | 26 | Female | Advanced Degree | Health/Fashion | No | Still Using | Basic |
P17 | 26 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Basic |
P18 | 18 | Male | High School | Health/Utility | No | Still Using | Advanced |
P19 | 36 | Male | Advanced Degree | Health/Utility | No | Stopped Using** | Advanced |
P20 | 29 | Male | Advanced Degree | Non-User | N/A | N/A | N/A |
P21 | 27 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
P22 | 27 | Female | Advanced Degree | Health/Utility | Yes | Stopped Using | Basic |
[The table listing participant demographics]Table 2 lists details of age, gender, education, usage/non-usage, level of health analytics, and Usage Abandonment for each participant by the uniquely assigned participant ID from the semi-structured Interviews.
3.2.2. Interview Design
The secondary component consisted of semi-structured interviews with 22 participants, all of whom were selected from the survey pool. The interviews were conducted based on the availability of the participants and were conducted until response saturation was reached on the emerging themes. Participants were chosen to reflect on various experiences with wearable technology for their PHI. The interview group included 16 males and 8 females, with a mean age of 28, from diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees. The detailed demographics are presented in Table 2. The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or stopped using their wearable devices. This diversity in the interview sample allowed for a comprehensive exploration of the different attitudes and behaviors toward wearable health technologies for PHI.
The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables. The interviews also explored participants’ broader health management practices and how wearables for PHI fit within this ”ecology of tracking.” This approach allowed the researchers to capture rich, contextual data that provided insights beyond what could be obtained from the survey alone.
ID | Age | Gender | Education | No. of Children |
CP1 | 34 | Female | Advanced Degree | 0 |
CP2 | 35 | Male | Bachelors | 0 |
CP3 | 33 | Female | Advanced Degree | 0 |
CP4 | 32 | Male | Bachelors | 0 |
CP5 | 27 | Female | Bachelors | 0 |
CP6 | 30 | Male | Bachelors | 0 |
CP7 | 33 | Male | Advanced Degree | 1 |
CP8 | 31 | Female | Advanced Degree | 1 |
CP9 | 29 | Female | Bachelors | 1 |
CP10 | 31 | Male | Bachelors | 1 |
CP11 | 38 | Female | Bachelors | 2 |
CP12 | 40 | Male | Bachelors | 2 |
[The table listing participant demographics of card sorting activity.]Table 3 lists details of age, gender, education, and number of children for each participant by the uniquely assigned participant ID from the co-design workshops.Advanced degree includes graduate degrees (like M.A., L.L.M., M.Sc., MS, MTech., M.Eng., MBA, CA, MD, MPhil, PhD, DPhil, etc.)
[Sample Card Sorting Activity Screens.]The screenshot shows the interface of the card sorting activity used by couples to group cards comprising of different health data (such as heart rate, SpO2, menstrual cycles, etc.) and aspects of sharing (such as control, privacy, and storage, etc.) to be sorted into categories of various stakeholder groups (such as parents, siblings, friends, doctors, etc.)
3.2.3. Co-design Workshops
One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information. This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI, such as Spouses, Parents/Adult Children, and Doctors, through the co-design workshops. To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6), Older Parent/Adult Children pairs (4), and doctors (4 from 3 specialties). To further identify the sharing preferences, we conducted a card sorting activity using kartSort (just using the couples as older participants struggled with the activity in the digital format and in-person activity was unfeasible as they lived across India) where they moved around different data/information cards to different stakeholders as categories. A sample screen from the card sorting activity is presented in Figure 1. We recruited six couples to reach saturation in understanding the variation of preferences among couples with and without a child (or children). The detailed demographics of the card sorting pairs are presented in Table 3.
[Sample Rapid Prototyping Screens.]The sample screenshot of the Figma prototype shows different screen flows in the prototype for selecting sharing preferences and approving requests for sharing among family members.
3.2.4. Rapid Prototyping and Evaluation
After the co-design workshops, we developed a medium fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members. This process helped us understand the ideal sharing preferences of individuals with various stakeholders involved. Figure 2 gives a sample screenshot of the prototype. The figma prototype was evaluated informally by sharing it with the participants and eliciting a qualitative evaluation. The unanimous positive takeaway from the participants’ feedback included its simplicity, modularity, and ease of control. Apart from these, there were individual features that the participant appreciated and provided further recommendations, which are discussed in the findings of this work. It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness, usage/non-usage of the same. The details of those observations are also presented in the findings, and the demographics were presented in 1.
Themes | Sub-Themes | Main Codes |
Collective health management | Use as a care device | Tracking heart rate & step count |
through sharing practices | Using because it was a gift | |
Track parents/partners health metrics | ||
Use as a health indicator | Use to monitor stress | |
Use to check critical conditions | ||
Step tracking as primary use | ||
Sleep tracking awareness | ||
Specific health tracking need | ||
(e.g., menstrual cycle) | ||
Digital PHI as a motivational tool | Perception on Wearable | Handy tool for tracking and utility |
for proactive health management | Useful motivational tool | |
Perception affecting degree of use | ||
Perceived low utility | ||
Trust in Data & Technology | Conditional trust in certain metrics | |
Acceptance of minor inaccuracies | ||
Useful as an approximate indicator | ||
Validation through comparison | ||
Low medical literacy on wearable data | ||
Navigating health information | Technical and Logistical Issues | Analytics not easy to understand |
in a fragmented digital ecosystem | Device management issues | |
Device incompatibilities | ||
Device change issues | ||
Buying a wearable is an investment | ||
Lifestyle and Physical Barriers | Lack of preferred modality of wearable | |
Rashes from the strap | ||
Minimal impact on doctor visits | ||
Physical discomfort on consistent use | ||
Preference for traditional methods | Trust on Body feeling | Will go to the doctor |
and perception of non-necessity | if I feel something wrong | |
Trust in health lifestyle/habits | ||
Perceptive lack of current need | ||
Reluctance to trust something new | ||
Trust in Doctors and | Will choose doctor for critical issues | |
digital PHI overwhelming | Do not know enough, doctors reliable | |
Wearable use overwhelming | ||
reactive care is enough | ||
Difference between body feeling | ||
and wearable metrics |
[Thematic Analysis table.]Table 4 lists the main themes, sub-themes, and the main codes obtained from the reflexive thematic analysis from the interview transcripts and the field notes taken during the interviews.
3.3. Data Analysis
We employed a combination of mixed methods to analyze the data. The survey responses were processed using descriptive statistics, where we calculated frequencies and percentages to assess patterns in the use and adoption of wearable health technologies for PHI, along with participants’ trust levels and perceived usefulness of these devices. To explore associations between demographic characteristics (such as age and gender) and usage patterns, we conducted cross-tabulations to identify significant relationships. We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, and field notes for the qualitative analysis. The analysis proceeded in several stages. Initially, the interview transcripts and field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns. Subsequently, preliminary codes were developed to capture key concepts and insights. These codes were refined and organized into broader, more comprehensive themes. From this iterative process, four principal themes emerged: (1) Collective health management through sharing practices; (2) navigating health information in a fragmented digital ecosystem; (3) digital PHI as a motivational tool for proactive health management; and (4) preference for traditional methods and perception of non-necessity. These themes highlighted how PHI is viewed and used within the context of personal health management among urban Indian participants. The thematic coding framework is detailed in Table 4, providing a structured overview of the identified themes and their subcategories.
3.4. Data Triangulation
To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature. This method enabled cross-verification and identification of patterns and discrepancies across data types, enhancing the validity of the results. For instance, survey and interview data revealed a lack of trust in wearable technologies due to perceived inaccuracies in advanced metrics. This notion was corroborated by existing literature that links health literacy and trust to digital health technologies (Bagalkot et al., 2020; Kreps, 2017). Our analysis of barriers to wearable adoption, such as a preference for traditional healthcare and the need for culturally tailored information, aligns with broader studies emphasizing context-specific approaches in non-Western settings (Ugargol and Bailey, 2018; Volpato et al., 2021). Further, we observed that participants often viewed wearables as supplementary rather than essential health tools for their PHI, reflecting a prevalent reactive approach to health management as shown by prior literature (Lee et al., 2022; Kang and Exworthy, 2022). By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings.
3.5. Positionality
The authors of this study acknowledge their positionality as researchers from the Indian cultural context. While some authors still live in India and bring an insider perspective, others live abroad with a background in distributed computing. This diverse positionality allowed for a comprehensive analysis considering local nuances and broader global trends in wearable technology use. The authors also recognize that their interpretations are influenced by their own experiences with using PHI and backgrounds, which may shape the analysis and conclusions drawn from the data. Throughout the research process, efforts were made to remain reflexive and mindful of potential biases, ensuring that the participants’ voices were accurately represented and that the analysis remained grounded in the data.
4. Findings
Using data triangulation from the various components of the study, we examined the perceptions, usage patterns, motivations, and barriers associated with PHI through an in-depth analysis. The thematic analysis revealed several key themes and sub-themes that offer valuable insights into how users interact with PHI, their trust in the technology, and their challenges/barriers to adoption.
4.1. Collective Health Management Through Sharing Practices
4.1.1. Health Seen as a Collective Responsibility
While major research has focused on individualistic healthcare models, the reflection of Indian practices indicates that health and wellbeing are shared/collective in nature. People often view the health of their family members as a shared responsibility, driving the need to keep on top of each other’s health statuses. For example, in our survey, over 60% (55/87) of participants were already tracking their health using a wearable, and in the interviews, about 45% (10/22) were gifted their tracking device to be used within families to track vital metrics, enabling a collaborative care model.
They viewed health management as a shared, collective process rather than an individual responsibility. They often rely on digital PHI not only to monitor their health but also to care for family members and close contacts. This shared approach fosters mutual support, accountability, and informed decision-making among trusted social circles. For instance, P8 disclosed that apart from getting a personal tracker, he also wanted to track his parents’ vitals to check on them even when living away.
”I also got a fitbit watch for my parents, and my parents wear them. So I can track heart rate and other metrics. So first, I bought for them, and then I got one for myself.” - P8 (M, 26)
Even among people who leave the country for work, their needs based on cultural values lead them to seek ways to track their parents’ health and support their health management remotely. This feeling was expressed through CP12’s personal experience of wanting to monitor his father’s health when he was living abroad a decade ago, but finding a lack of cost-effective means to do so.
”During my initial career days (living abroad), as my dad got surgery, I wanted to be able to monitor him regularly. But there was no cheap and reliable means to do so. So, I had to rely on relatives to inform me of his status. If there was an economical way to monitor my father continuously, it would have helped me a lot.” - CP12 (M, 40)
4.1.2. Understanding the Aspects of Sharing
Through the secondary interviews, co-design sessions on sharing preferences, and a Sharing Awareness Survey, we discovered that individuals are open to sharing their PHI with spouses and doctors without reservation, but are unaware of existing mechanisms to share. They had varied opinions on how they would like their data to be shared and used. Some couples (2/6) in the co-design session were open to their data being used by others as long as they were informed of the use case, while all the rest wanted control over not only sharing but also the storing and usage of the data. The participants’ interest in viewing all kinds of health information about themselves and their family members in a single place showcases a need for a unified system.
The level of openness with sharing decreased in the following order among the participants: Spouses Doctors Parents Siblings Friends Employer Insurance Companies. Some couples (2/6) also mentioned that they wanted to avoid sharing with employers and insurance companies if that was the option.
”So, I am open to sharing everything with my spouse. Maybe just an update on how I am doing with my parents maybe. I don’t think I’ll be sharing much with my siblings or friends except if there are any emergency alerts. I am also open to sharing everything with doctors as long as it makes sense for the doctor to be seeing those for the purpose of the visit.” - CP1 (F, 34)
When inquired about their knowledge and usage of any current sharing frameworks, most of the couples (5/6) were unaware of those already in the market through prominent brands (Apple Health and Samsung Health). This general lack of awareness and usage of such sharing frameworks was also reinforced through the outcome of the survey responses (94/114) from the health sharing feature survey. The aware couple said they were already using it (however sparingly) for things that can be tracked through their wearable. However, they indicated they are unhappy about the lack of control over what can be shared, how it can be shared, and when it can be shared. This observation hints that either the existing platforms are so rigid that users are unhappily using them as little as possible, or they are unaware of it because it is not popular in the urban Indian context. A couple even drew insights from their experience living abroad and the system they have seen being used.
”Sharing within spouses is OK, but what about the other stakeholders like employers or insurance companies? do we get a say on what data goes to them or how they use it? That is what I am looking for, if there is an application to seamless share my health information. See, for example, when we were in the UK, there was this app that shows all details concerning our GP, from appointments to medication. When we want to share the medication information with the chemist/pharmacist, we just need to hit share, and we can select what we want to share. That level of control would be very useful.” - CP6 (M, 30) & CP5 (F, 27)
Another insightful observation among participants interested in sharing was that people with children are interested in tracking and sharing PHI only when there is no overhead of pulling all things together by themselves. While the motivation to track using wearables differs individually (even within a couple), they all were motivated for the overall PHI for themselves and their children if they could do so using a single device.
”See, in this digital age, almost every data is online. Not only with each other but also with companies like Google or Apple (laughs). After this interview, I will probably get an Instagram reel on how couples should track data together (laughs). So, there is no need for stricter control within the couple on how/what/when to share. We currently use whatever means possible (mostly just WhatsApp each other on health documents and information), we are open to a dedicated platform only if we do not have to add everything, every time manually.” - CP7 (M, 33) & CP8 (F, 31)
After incorporating the feedback and inputs, when evaluating the rapid figma prototype, the participants liked the overall functionality of the interface and provided further suggestions. They felt that the prototype mimicked the discussion on their sharing preferences and appreciated the control over them. Some participants also suggested that further information or health tips could be added about the benefits of good sleep and (for women) steps to take when periods are late.
”Some family tracking and reward like a badge to the member who completes all goals, just to have encouraging competition in the family for better health would be nice.” - CP1 (F, 34)
”I like the daily, weekly, and monthly stats being shown for health features. But I would also like to have a graphical representation.” - (CP3, F, 33)
The responses from the Sharing Feature Awareness Survey also indicate a need for a reliable medium of sharing information among Android users and a better demonstration of capabilities among Apple/iPhone users.
”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage. If they can be stored all together in one place, that would be a helpful feature. ” - (F, 33)
”I thought sharing was about putting it out on social media platforms. I never knew about sharing features that send health updates to our family or health care provider.” - (M, 24)
Overall, participants’ motivations and needs indicate a need for a single dedicated application or a sharing mechanism where they can pool all kinds of health information about themselves and their trusted circles while maintaining full control over the preferences, such that they do not have additional cognitive load on managing it. While some prefer textual representation of this information, others are comfortable with graphical, while a very limited few are okay with numbers and medical jargon.
4.2. Navigating Health Information in a Fragmented Digital Ecosystem
4.2.1. Unreliable Tools and Systems
Users encounter a patchwork of digital health information, often leaving them uncertain about what to trust. While basic functionalities (like step counting and heart rate monitoring) are widely accepted, there is notable skepticism around advanced metrics. Participants frequently cross-check data with multiple devices, healthcare professionals, and even traditional methods to validate accuracy. This verification process underscores the challenge of navigating a fragmented digital health landscape.
”So I think, What is missing in current systems is having that authenticity or maybe accuracy? I don’t blindly believe them, like, OK, it’s saying something is wrong, so it must be, but I won’t ignore it either. If my tracker is giving some wrong data, I will check with another and see whether it’s me or it’s the sensor, but if 2 or 3 trackers have been giving you that data, I don’t think there is any need to doubt it. (laughs).” - (P4, F, 33)
Even the participants who have used various trackers throughout their evolution noted that the current capabilities of PHI systems are not on par with their current purpose or motivation. They emphasized the lack of reliable sources for health information or even the capability of systems to explain their health in understandable terms.
”So one of my first smartwatches was a Pebble smartwatch. It is very old. It started on Kickstarter (—) So that is how I got into it. (—) But more lately. It’s been to serve the purpose of. OK, how do I understand my body a little better? (—) So what it started with and what it is right now, it’s very different. The current systems cannot explain it well to me.” - P19 (M, 36)
4.2.2. Island of Ecosystems
Many of these PHI trackers have their own ecosystems that are specialized for a specific brand or product line, focused on boosting their unique functionalities to promote their line of products. This causes a lack of cross-functional capabilities. This was echoed during the interviews with doctors with different specialties, the common recurring theme was the existing use of specialized trackers for different chronic conditions like diabetes, digital pressure cuffs for hypertension, or even mood journaling for psychiatry. The doctors from endocrinology and medicine confirmed using existing specialized smart devices in practice, albeit their price factor limited it to a few patients. The high financial investment required for reliable devices further compounds these reservations for individuals as well.
”It’s kind of an investment, like it is at least for me. So buying a watch which which will cost 20,000/25,000 at the time of the release, it is kind of an expensive affair. So you cannot change these things with every update. ” - P3 (F, 33)
Wearables, an integral part of PHI for many users, had technical issues apart from their limitations of cross-functionality on what they could offer the user, causing further frustrations among participants. One such instance is inaccurate heart rate measurements needing recalibration during exercise. This observation shows that mechanisms such as comparison tools or validation indicators that help users assess the accuracy of their health data are currently lacking.
”It happens quite a lot, especially in the gym when I can feel that my heart is like pounding (laughs); like help!. OK, so I can feel my heart beating really fast. But when I see my watch, it will show readings like 75 or 76, which is kind of normal. (—) Then, then I have to again, again like reconfigure that (manually calibrate the reading). Then you put, then you press on that measure. So some kind of calibration is required.” - P3 (F, 33)
Participants noted the lack of reliable health information sources for them to verify their PHI and the reduced level of interoperability among available devices. They expressed that they do not have verified, reputable health information sources to compare their metrics and rely on doctor visits to make sense of them.
”I do not understand whether the heart rate is good for me or not, the tracker does not tell me that. When I had the time to go to my doctor that time also I was made to understand that it is in between 72 and 80, which was fine according to the doctor. Presently also, it is nearly like that. When I walk, it shows 78 to 82; when I am not, it shows something about 68 to 72. So, I am guessing it is OK, even though I am not 100% sure.” - (P10, M, 62)
Overall, the observations with participants’ struggles to make sense of the data indicate that the fragmented ecosystems are complex and lack inter-compatibility to use them all cohesively for PHI. Simplifying the complexity of digital health data through intuitive design and user-friendly explanations that boost both digital and health literacy could aid in their journey better, as noted by a few participants.
”So if you get a chance to kind of have that information to help you navigate through that (PHI), it will be great. Because, at the end of the day, being healthier is everyone’s priority. The lack of such information or even the means to get them is stopping people (from proactive health management), with other priorities getting ahead.” - (P9, F, 29)
4.3. Digital PHI as a Motivational Tool for Proactive Health Management
Digital PHI catalyzes behavior change by transforming passive data collection into proactive health management. Users appreciate features that provide real-time feedback, set personalized goals, and track progress over time. These tools encourage people to adopt healthier lifestyles, offering actionable insights and reminders. However, the motivational impact is closely tied to individuals’ priorities and preferences. For example, some participants were mindful of their lifestyle, like P16 (F, 26) and P7 (M,32), who track their steps and maintain an active lifestyle during sedentary workdays.
”Since I have a desk job, so I have to make sure that I’m doing at least, like so far that I come across the Internet, they do mention that you should be at least walking 5 kilometers if you are sitting all day and working. So that is one thing I keep in mind. That’s what makes me keep a check on my step count and how much I’m walking.” - (P16, F, 26)
”So I was walking very fast. Then it (the tracker) told me you are walking outdoors; you have walked so much distance. So then, I wish to walk more.” - (P7, M, 32)
Integrating reminders, gamified challenges, and progress-tracking features that motivate users to stay engaged with their health data was among the recommendations from participants in the interviews and rapid prototyping. They considered them resourceful motivators, a means to make healthy habits, or a means for a more personalized health management tool by delivering tailored advice or actionable insights that resonate with their individual and family health goals.
”So the (health) reminder thing, right, it can be a good guilt-tripping system so. (—) more like a motivator (to work out more). This would be a good, good thing to see, like a progress of myself.” - (P17, M, 26)
Participants were eager to view summarized stats over day, week, and month. However, they also wished for them to be presented alternately to increase their engagement. Participants noted that being handy could also be another reason for adopting digital PHI as a reliable tool. This finding indicates the potential of digital PHI as a motivational tool if different presentation modalities are provided.
4.4. Preference for Traditional Methods and Perception of Non-necessity
4.4.1. Traditional and Cultural Influence
Some participants remain cautious about fully adopting digital PHI, favoring traditional health management methods. Concerns about privacy, technological complexity, and physical discomfort (e.g., frequent charging needs and skin irritation) contribute to this resistance. Some users (5) trust their bodily awareness and conventional healthcare practices, finding digital tracking unnecessary or overwhelming.
”I do not prefer to constantly check my miband. I only use it when I run or perform any exercise. I do not find the need to check the history or track, as if something is wrong, I could feel it in my body.” - P21(M,27)
Similarly, P20 (M, 29) emphasized his belief that his body would alert him if something were wrong, reducing the perceived need for a smartwatch and confidence in his healthier lifestyle.
”Whenever you know I’m like whenever I’m sick, (laughs) maybe at the time, something is wrong and I don’t, as you said (tracking health metrics regularly) maybe I’m not much bothered about, the specifics. (—) If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad. So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
Some participants even recognized and reasoned about the disjointedness between the customs of their parents’ generation and the current one, indicating a need for better integrating digital PHI with traditional cultural practices to promote active, healthy living.
”If you think about it, our lifestyles has been completely different from our parents. These (PHI) trackers only work for us. Designers are completely ignoring the population that still leads a more self-reliant lifestyle. If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
While participants who seek a more body-over-technology approach exist, they also make recommendations in ways that can improve for a wider audience. This observation shows that the adoption is dictated by the value it offers, which does not meet their needs, and indicates room for improvement.
4.4.2. Barriers or Challenges Causing Digital PHI Resistance
Participants considered an assurance of precision in both basic and advanced features a necessity to maintain trust and encourage ongoing proactive engagement. They compared their childhood experience with traditional measuring devices as a frame of reference for building trust in digital PHI systems, indicating a need to get used to such digital systems.
”Maybe we are just used to a different kind of devices. For example, growing up, for the blood pressure machine (makes hand motions of the pumping used in a blood pressure cuff) (laughs), that thing, like you see something happening. (—) I think it’s one of the perceptive Issues. And also because I’m not used it a lot. I don’t have the perception, and its that.” - (P5, F, 29)
Comfort and lifestyle fit have also emerged as critical factors influencing the continued use of smart wearables as care devices, causing the aspects of using a wearable to become a barrier for a few. For example, P9 (F, 29), who received a smartwatch as a gift, eventually found the device’s frequent charging needs and complexities overwhelming, not integrating seamlessly into her life.
”So, so one, once I basically invested in this watch. I mean it was all fancy and all of it (—) What I realized is that this was too much of a task for me. The most painful thing is to kind of charge the device. (—) I wanted to do something to kind of keep my health in check, but it got too much for me.” - P9 (F,29)
For others, the physical discomfort or rashes due to wearable material can detract from their overall utility, and others prefer different modalities such as rings, bands, or other wearables to a watch so they can wear their traditional watches, which they enjoy wearing.
”For me, specifically, I do not prefer to wear it continuously because I get rashes from the strap if I wear it for too long regardless of the kind of strap.” - P22 (F,27)
”I’m also trying to go away from a watch format and more a band or a ring or something else of that sort. (—) Mostly because I also wanna like continue wearing my other watches.” - (P19, M, 36)
Overall, the findings indicate that digital PHI can offer options that allow for both proactive tracking and reactive monitoring, acknowledging users’ comfort with conventional health cues would be welcome among users. They also recommend cost-effective solutions to ease the transition from traditional to digital health management.
5. Discussion
Our findings shed light on how urban Indians adopt, use, and perceive health wearables for PHI in India, emphasizing the cultural, societal, and technological factors influencing user behavior. While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we uncover a broader gap: how these same technologies are—or could be—woven into routine health management practices. By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape PHI adoption. Below, we discuss three key themes—navigating a fragmented ecosystem, wearables as care devices, and culturally sensitive engagement —and offer design and policy recommendations to support a more culturally attuned PHI future.
5.1. Navigating a Fragmented Digital Health Ecosystem: Technology & Policy
India’s digital health landscape is marked by fragmentation and the absence of centralized, reliable information—conditions that contrast sharply with the integrated systems typical in Western contexts. This fragmentation hampers PHI usability, trust, and effectiveness.
Participants reported difficulty navigating multiple disjointed platforms, often defaulting to traditional health practices. Choosing a tracker was perceived as a significant investment, complicated by non-compatible ecosystems—a clear barrier (section 4.2). While PHI systems evolve to support personal health management, the IT infrastructure backing them remains outdated and fragmented (Jones et al., 2012). Proprietary ecosystems hinder interoperability, complicating and increasing the cost of decision-making. For example, a Samsung smartwatch user with an iPhone or non-Samsung Android device cannot access all features or rely on data accuracy, illustrating how such silos generate user confusion and disengagement. This calls for design strategies that reduce the cognitive load of choice and simplify use. As seen in our prototype evaluation, users need flexible interaction modes and wearable form factors tailored to their comfort levels (section 4.4.2). While some of these challenges can be overcome through software/interaction design, the rest require policy-level changes that enforce manufacturers to comply with an open data framework that ensures accessibility of health data regardless of the medium or proprietary ecosystem requirements of individual brands/companies.
Furthermore, the fragmented digital ecosystem and the accompanying mistrust raise ethical and privacy concerns. This aspect of PHI must be balanced with the need for integrated, actionable health information. While countries worldwide have policies regarding collecting and using health information (for example, HIPAA, PHIPA, GDPR, and APPs), India still uses the IT Act of 2000 even for medical and health information. The IT Act 2000 provides basic data protection by requiring data owners to be informed about its use, but it is not tailored to healthcare. The National Digital Health Blueprint (NDHB) recommends ISO 22600:2014 Health Informatics, but implementing these standards may be challenging for many healthcare providers in India, and ensuring compliance at the provider level could also be challenging for the government. Although similar healthcare-specific legislation was proposed in 2017 and 2022, it has not been implemented. While the Government of India has now published the “Digital Personal Data Protection Act, 2023” (DPDP 2023) covering all personal data (including health data), it remains unenforced and is not robust enough to meet the requirements posed by Indian users’ socially reflective shared caring models.
We propose a unified, open framework that aggregates PHI, doctor visits, lab reports, medications, and more into a single interface. This interface should support all smart wearables—focused on Android, which dominates the Indian market (94%)—and employs accessible language to reduce learning burdens and foster proactive health behaviors.
5.2. Wearables as Care Devices: Collective Care & Engagement Strategies
Our participants overwhelmingly described health wearables not merely as personal gadgets but as instruments of family‐centered care (findings 4.1.1). Unlike the individualist framing common in Western studies (Chmielewski and Meyer, 2024), Indian users envision sharing health metrics/analysis with parents, spouses, or children, turning data streams into collaborative decision‐making tools. We observed that users readily assume caregiver roles—monitoring a parent’s daily steps or a child’s sleep patterns —and express enthusiasm for co‐designed interfaces that let family dyads explore and interpret data together (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023).
To drive proactive engagement, gamification of PHI systems must capture activities familiar to Indian lifestyles—such as traditional labor and household tasks—rather than only Western‐style workouts to motivate them better (findings 4.1.2 & 4.3). At the same time, designs should include contextual alerts (for example, flagging sustained deviations in resting heart rate) that prompt timely medical check‐ins, moving users from reactive responses toward ongoing health management. Prior studies on intergenerational sharing of health data (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024a), women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a), and children-specific care for chronic ailments (Cha et al., 2024), reinforce that broadening the scope of shareable metrics can foster more comprehensive, preventive care—particularly where caregiving is a shared, familial endeavor.
Privacy concerns were minimal among our participants (finding 4.1.2), indicating openness to data exchange within trusted circles. This suggests potential for paired or multi‐user wearable configurations, provided that the system offers clear, in-app explanations of how data are collected, processed, and used. Policymakers can foster privacy-conscious behaviors by enforcing clear regulations around data protection, specifically in PHI wearables. As people adopt a more proactive approach, wearables’ role in caregiving could be significantly enhanced by providing continuous insights that allow early intervention before critical incidents. The ability to share comprehensive health data with family members could also strengthen the perceived value of these systems, encouraging consistent use and deeper integration into daily routines. Designers and developers must therefore prioritize features that integrate into users’ lives, facilitating comprehensive health monitoring and sharing. Such usability and transparency, coupled with policy measures mandating robust PHI protections, will be essential to transforming wearables from novelty gadgets into indispensable tools for early intervention and collaborative health management. Our observations showcase that the ”Personal” in PHI involves more than the individual in the Indian context. These family-centered insights directly inform our unified PHI platform vision in the next section.
5.3. Culturally Sensitive Engagement: Norms, Literacy, Misinformation, & Trust
In India, paper-and-analog practices remain deeply trusted, but users embrace digital PHI when it feels like an extension, not a replacement. Participants, for example, described “smart” digital pressure cuffs that look and behave like traditional sphygmomanometers yet auto-sync readings to a unified PHI dashboard (findings 4.4.2). We frame this as a modular hybrid PHI paradigm, where users toggle between digital and analog interfaces based on context and comfort. Such an approach acknowledges India’s healthcare realities and offers a pragmatic, modular pathway for digital transformation—one that integrates seamlessly with generic platforms.
In a society where talking openly about intimate or sexual health is still taboo, cultural norms around sharing vary by metric: intimate health data (e.g., sexual health) was shared only with spouses, whereas vitals like step counts traveled freely among parents, children, or extended family depending on trust (findings 4.1.1). Parents even leaned on devices recommended by their children, illustrating “derived trust.” To embrace these differences, we argue for a culturally informed data-governance model with user-controlled sharing settings and transparent, in-app analytics—so families can collaborate on care without over- or under-sharing. Future policy frameworks should incorporate these principles to foster an ecosystem where digital health innovations are ethically grounded and culturally resonant.
Low health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) further compound these challenges and impact adoption. Participants distrusted novel features when metrics seemed “off” or unfamiliar/unevaluated (section 4.2.2), and authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently. This disparity highlights the need for authentic, reliable, vetted, user-friendly health information platforms so Indians can access trustworthy guidance without skepticism. And since the pandemic has driven more health-conscious behaviors, implementing the long-overdue 2005 school-health curriculum recommendation (MoHFW, 2005) could foster health-conscious behavior from a young age (Kann et al., 2007; Pulimeno et al., 2020; Auld et al., 2020; Benes and Alperin, 2022; Sulz et al., 2024).
Trust hinges on data accuracy and medical validation. Wearables often function as “black boxes,” limiting lay users’ comprehension (Shin et al., 2019; Zhang and Shahriar, 2020). Sustained adoption, therefore, requires medical validation and clinical integration. Devices should earn formal endorsements, sync with electronic health records, and surface sensor-accuracy indicators (e.g., confidence intervals around SpO2 or heart-rate) to bolster user confidence (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023). Integrating wearables with formal healthcare systems will further bolster their credibility and sustained use.
These collective/shared care insights directly inform our integrated PHI platform that consolidates disparate data, embeds user-controlled sharing and transparent analytics, employs inclusive design methods (e.g., family workshops), and links seamlessly with clinical systems—so that as health literacy grows, users can critically weigh benefits against risks (Mackert et al., 2016; Meyers et al., 2020).
5.4. Future Work
While there are novel recommendations on personal, public, and cultural aspects of PHI adoption, the authors acknowledge several limitations in this study. First, the cross-sectional design captures user behavior and perceptions at a single point in time, limiting the ability to observe how these factors may evolve. Additionally, the focus on urban, middle- and upper-middle-class participants restricts the generalizability of the findings to rural and lower-income populations, where different adoption barriers, such as different lifestyle practices, natural remedies, and many more, may exist. Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their PHI usage, utility, and adoption.
Building upon our observations and the strategic direction outlined, we propose the design and development of a proof-of-concept application aimed at addressing general health, wellbeing, and medical inquiries. This system will also assist users in interpreting and contextualizing their personal health data aggregated from diverse sources, presenting the information in a format that is both comprehensible and actionable. Leveraging advancements in artificial intelligence, particularly finetuned open-source large language models (LLMs), the application will integrate data from clinical documents (e.g., physician reports, laboratory results, and medication records) alongside information from wearable health devices. This multimodal data synthesis will enable the generation of personalized, holistic insights, thereby enhancing user engagement and health awareness. The interface will employ a conversational, chatbot-style design to facilitate intuitive interaction, mediate family sharing norms, and surface literacy-appropriate explanations. Beyond helping users understand their personal health information (PHI), the system will also provide access to evidence-based health, medical, and wellness content. By sourcing validated information from reputable sources, the application seeks to address the pervasive issue of health misinformation and contribute to the improvement of public health literacy.
6. Conclusion
This study comprehensively explores how Indian users integrate PHI wearables for health management. Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of PHI in India. By doing so, we emphasize the influence of sociocultural factors, usability challenges, and trust issues in their consistent/long-term use of PHI for health management. While PHI and specifically wearables hold significant potential to shift personal health management from a reactive to a proactive model, their integration is hindered by cultural norms, limited health literacy, and skepticism around data accuracy. To address these challenges, we propose a dual focus on improving health literacy and building trust through transparent communication, medical validation, and user-centered design.
Two essential contributions of this research are identifying critical areas where designers and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants. The critical areas include expanding PHI functionalities to align with family-based caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information. As shown by the evaluated prototype, the HCI community, technology developers, and policymakers must work together to standardize core functionalities and integrate PHI into formal healthcare systems. By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive health management, supporting both individual users and familial care structures.
This work offers a framework for future research and industry efforts, reinforcing the transformative potential of PHI within diverse cultural contexts. By addressing technological and sociocultural barriers, PHI can aid in preventive healthcare, fostering long-term engagement and broader adoption across India and beyond.
Acknowledgements.
We would like to thank the student Harsh Pala for his involvement in the development of Figma prototypes used for evaluating an interactive interface of the sharing preferences identified by the card sorting activity.References
- (1)
- Alrowaily (2024) Majed Abdullah Alrowaily. 2024. Smart Emergency Alerting System: A Machine Learning Approach. In Proceedings of the 2024 2nd International Conference on Communications, Computing and Artificial Intelligence (CCCAI ’24). Association for Computing Machinery, New York, NY, USA, 8–15. https://doi.org/10.1145/3676581.3676583
- Amato et al. (2022) Flora Amato, Walter Balzano, and Giovanni Cozzolino. 2022. Design of a Wearable Healthcare Emergency Detection Device for Elder Persons. Applied Sciences 12, 5 (Jan. 2022), 2345. https://doi.org/10.3390/app12052345
- Anderson and Durstine (2019) Elizabeth Anderson and J. Larry Durstine. 2019. Physical Activity, Exercise, and Chronic Diseases: A Brief Review. Sports Medicine and Health Science 1, 1 (Dec. 2019), 3–10. https://doi.org/10.1016/j.smhs.2019.08.006
- Ankrah et al. (2023) Elizabeth A. Ankrah, Franceli L. Cibrian, Lucas M. Silva, Arya Tavakoulnia, Jesus A. Beltran, Sabrina E.b. Schuck, Kimberley D. Lakes, and Gillian R. Hayes. 2023. Me, My Health, and My Watch: How Children with ADHD Understand Smartwatch Health Data. ACM Trans. Comput.-Hum. Interact. 30, 4 (Sept. 2023), 59:1–59:25. https://doi.org/10.1145/3577008
- Auld et al. (2020) M. Elaine Auld, Marin P. Allen, Cicily Hampton, J. Henry Montes, Cherylee Sherry, Angela D. Mickalide, Robert A. Logan, Wilma Alvarado-Little, and Kim Parson. 2020. Health Literacy and Health Education in Schools: Collaboration for Action. NAM Perspectives 2020 (July 2020), 10.31478/202007b. https://doi.org/10.31478/202007b
- Bagalkot et al. (2020) Naveen Bagalkot, Nervo Verdezoto, Anushri Ghode, Shipra Purohit, Lakshmi Murthy, Nicola Mackintosh, and Paula Griffiths. 2020. Beyond Health Literacy: Navigating Boundaries and Relationships During High-risk Pregnancies: Challenges and Opportunities for Digital Health in North-West India. In Proceedings of the 11th Nordic Conference on Human-Computer Interaction: Shaping Experiences, Shaping Society (NordiCHI ’20). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3419249.3420126
- Bastico et al. (2022) Matteo Bastico, Verónica Ruiz Bejerano, and Alberto Belmonte-Hernández. 2022. Simultaneous Real-Time Human Fall Detection and Reidentification Based on Multisensors Data. In Proceedings of the 15th International Conference on PErvasive Technologies Related to Assistive Environments (PETRA ’22). Association for Computing Machinery, New York, NY, USA, 365–370. https://doi.org/10.1145/3529190.3534728
- Bat-Erdene and Saver (2021) Bat-Orgil Bat-Erdene and Jeffrey L. Saver. 2021. Automatic Acute Stroke Symptom Detection and Emergency Medical Systems Alerting by Mobile Health Technologies: A Review. Journal of Stroke and Cerebrovascular Diseases 30, 7 (July 2021), 105826. https://doi.org/10.1016/j.jstrokecerebrovasdis.2021.105826
- Benes and Alperin (2022) Sarah Benes and Holly Alperin. 2022. The Essentials of Teaching Health Education: Curriculum, Instruction, and Assessment. Human Kinetics, United States.
- Bhat et al. (2023) Karthik S. Bhat, Amanda K. Hall, Tiffany Kuo, and Neha Kumar. 2023. ”We Are Half-Doctors”: Family Caregivers as Boundary Actors in Chronic Disease Management. Proc. ACM Hum.-Comput. Interact. 7, CSCW1 (April 2023), 111:1–111:29. https://doi.org/10.1145/3579545
- Bhat and Kumar (2020) Karthik S. Bhat and Neha Kumar. 2020. Sociocultural Dimensions of Tracking Health and Taking Care. Proc. ACM Hum.-Comput. Interact. 4, CSCW2 (Oct. 2020), 129:1–129:24. https://doi.org/10.1145/3415200
- Binda et al. (2017) Jomara Binda, Hyehyun Park, John M. Carroll, Natalie Cope, Chien Wen (Tina) Yuan, and Eun Kyoung Choe. 2017. Intergenerational Sharing of Health Data among Family Members. In Proceedings of the 11th EAI International Conference on Pervasive Computing Technologies for Healthcare (PervasiveHealth ’17). Association for Computing Machinery, New York, NY, USA, 468–471. https://doi.org/10.1145/3154862.3154895
- Blandford (2019) Ann Blandford. 2019. HCI for Health and Wellbeing: Challenges and Opportunities. International Journal of Human-Computer Studies 131 (Nov. 2019), 41–51. https://doi.org/10.1016/j.ijhcs.2019.06.007
- Braun and Clarke (2006) Virginia Braun and Victoria Clarke. 2006. Using Thematic Analysis in Psychology. Qualitative Research in Psychology 3, 2 (Jan. 2006), 77–101. https://doi.org/10.1191/1478088706qp063oa
- Bronfenbrenner (1974) Urie Bronfenbrenner. 1974. Developmental Research, Public Policy, and the Ecology of Childhood. Child Development 45, 1 (1974), 1–5. https://doi.org/10.2307/1127743
- Brun and Häkkilä (2022) Damien Brun and Jonna Häkkilä. 2022. Exploration of Electrochromics for Calm Reminder: A Customizable Bracelet. In Proceedings of the 33rd Australian Conference on Human-Computer Interaction (OzCHI ’21). Association for Computing Machinery, New York, NY, USA, 269–272. https://doi.org/10.1145/3520495.3520509
- Cagiltay and Mutlu (2024) Bengisu Cagiltay and Bilge Mutlu. 2024. Toward Family-Robot Interactions: A Family-Centered Framework in HRI. In Proceedings of the 2024 ACM/IEEE International Conference on Human-Robot Interaction (HRI ’24). Association for Computing Machinery, New York, NY, USA, 76–85. https://doi.org/10.1145/3610977.3634976
- Carter et al. (2014) Nancy Carter, Denise Bryant-Lukosius, Alba DiCenso, Jennifer Blythe, and Alan J. Neville. 2014. The Use of Triangulation in Qualitative Research. Oncology Nursing Forum 41, 5 (Sept. 2014), 545–547. https://doi.org/10.1188/14.ONF.545-547
- Cha et al. (2024) Yoon Jeong Cha, Yasemin Gunal, Alice Wou, Joyce Lee, Mark W Newman, and Sun Young Park. 2024. Shared Responsibility in Collaborative Tracking for Children with Type 1 Diabetes and Their Parents. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–20. https://doi.org/10.1145/3613904.3642344
- Chandrasekaran et al. (2020) Ranganathan Chandrasekaran, Vipanchi Katthula, and Evangelos Moustakas. 2020. Patterns of Use and Key Predictors for the Use of Wearable Health Care Devices by US Adults: Insights from a National Survey. Journal of Medical Internet Research 22, 10 (Oct. 2020), e22443. https://doi.org/10.2196/22443
- Cherian et al. (2024) Josh Cherian, Samantha Ray, Thomas Mernar, Paul Taele, Helen Mach, Jung In Koh, Ping Ye, and Tracy Hammond. 2024. A Step Toward Better Care: Understanding What Caregivers and Residents in Assisted Living Facilities Value in Health Monitoring Systems. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 13:1–13:29. https://doi.org/10.1145/3637290
- Chmielewski and Meyer (2024) Marta Chmielewski and Matthew J. Meyer. 2024. Discussions With End Users to Inform the Vision for a Shared Care Record in Ontario: Qualitative Interview Study. Online Journal of Public Health Informatics 16, 1 (Feb. 2024), e51231. https://doi.org/10.2196/51231
- Costa Figueiredo and Chen (2021) Mayara Costa Figueiredo and Yunan Chen. 2021. Health Data in Fertility Care: An Ecological Perspective. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3411764.3445189
- Creaser et al. (2022) Amy V. Creaser, Jennifer Hall, Silvia Costa, Daniel D. Bingham, and Stacy A. Clemes. 2022. Exploring Families’ Acceptance of Wearable Activity Trackers: A Mixed-Methods Study. International Journal of Environmental Research and Public Health 19, 6 (March 2022), 3472. https://doi.org/10.3390/ijerph19063472
- Delahoz and Labrador (2014) Yueng Santiago Delahoz and Miguel Angel Labrador. 2014. Survey on Fall Detection and Fall Prevention Using Wearable and External Sensors. Sensors 14, 10 (Oct. 2014), 19806–19842. https://doi.org/10.3390/s141019806
- Egede et al. (2021) Joy Egede, Maria J. Galvez Trigo, Adrian Hazzard, Martin Porcheron, Edgar Bodiaj, Joel E. Fischer, Chris Greenhalgh, and Michel Valstar. 2021. Designing an Adaptive Embodied Conversational Agent for Health Literacy: A User Study. In Proceedings of the 21st ACM International Conference on Intelligent Virtual Agents (IVA ’21). Association for Computing Machinery, New York, NY, USA, 112–119. https://doi.org/10.1145/3472306.3478350
- Epstein et al. (2015) Daniel A. Epstein, An Ping, James Fogarty, and Sean A. Munson. 2015. A Lived Informatics Model of Personal Informatics. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 731–742. https://doi.org/10.1145/2750858.2804250
- Evans et al. (2022) Hayley Irene Evans, Catherine R Deeter, Jiawei Zhou, Kimberly Do, Andrew M Sherrill, and Rosa I. Arriaga. 2022. Perspectives on Integrating Trusted Other Feedback in Therapy for Veterans with PTSD. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3491102.3517513
- Fritz et al. (2014) Thomas Fritz, Elaine M. Huang, Gail C. Murphy, and Thomas Zimmermann. 2014. Persuasive Technology in the Real World: A Study of Long-Term Use of Activity Sensing Devices for Fitness. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’14). Association for Computing Machinery, New York, NY, USA, 487–496. https://doi.org/10.1145/2556288.2557383
- Ghosh and Das (2024) Manthan Ghosh and Deepa Das. 2024. Voice-Activated SOS: An AI-Enabled Wearable Device. In Impact of AI on Advancing Women’s Safety. IGI Global, Pennsylvania, USA, 251–277. https://doi.org/10.4018/979-8-3693-2679-4.ch016
- Gjoreski et al. (2016) Hristijan Gjoreski, Jani Bizjak, and Matja Gams. 2016. Using Smartwatch as Telecare and Fall Detection Device. In 2016 12th International Conference on Intelligent Environments (IE). IEEE, London, UK, 242–245. https://doi.org/10.1109/IE.2016.55
- Gooch et al. (2021) Daniel Gooch, Blaine A. Price, Anna Klis-Davies, and Julie Webb. 2021. A Design Exploration of Health-Related Community Displays. Proc. ACM Hum.-Comput. Interact. 5, CSCW1 (April 2021), 85:1–85:22. https://doi.org/10.1145/3449159
- Grimme et al. (2024) Sophie Grimme, Susanna Marie Spoerl, Susanne Boll, and Marion Koelle. 2024. My Data, My Choice, My Insights: Women’s Requirements When Collecting, Interpreting and Sharing Their Personal Health Data. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–18. https://doi.org/10.1145/3613904.3642851
- Gustafsson et al. (2022) Lena-Karin Gustafsson, Ildikó Asztalos Morell, Carl Johansson, and Santa De. 2022. Informal Caregiving from the Perspectives of Older People Living Alone in India. International Journal of Older People Nursing 17, 6 (Nov. 2022), e12468. https://doi.org/10.1111/opn.12468
- Haque et al. (2024) MD Romael Haque, Zeno Franco, Praveen Madiraju, Natalie D Baker, Sheikh Iqbal Ahamed, Otis Winstead, Robert Curry, and Sabirat Rubya. 2024. ”Butt Call Me Once You Get a Chance to Chat 🙂” : Designing Persuasive Reminders for Veterans to Facilitate Peer-Mentor Support. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3613904.3642962
- Hoque and Bao (2015) Md. Rakibul Hoque and Yukun Bao. 2015. Cultural Influence on Adoption and Use of E-Health: Evidence in Bangladesh. Telemedicine and e-Health 21, 10 (Oct. 2015), 845–851. https://doi.org/10.1089/tmj.2014.0128
- Huang et al. (2022) Yue Huang, Borke Obada-Obieh, Elissa M. Redmiles, Satya Lokam, and Konstantin Beznosov. 2022. COVID-19 Information-Tracking Solutions: A Qualitative Investigation of the Factors Influencing People’s Adoption Intention. In Proceedings of the 2022 Conference on Human Information Interaction and Retrieval (CHIIR ’22). Association for Computing Machinery, New York, NY, USA, 12–24. https://doi.org/10.1145/3498366.3505756
- Huhn et al. (2022) Sophie Huhn, Miriam Axt, Hanns-Christian Gunga, Martina Anna Maggioni, Stephen Munga, David Obor, Ali Sié, Valentin Boudo, Aditi Bunker, Rainer Sauerborn, Till Bärnighausen, and Sandra Barteit. 2022. The Impact of Wearable Technologies in Health Research: Scoping Review. JMIR mHealth and uHealth 10, 1 (Jan. 2022), e34384. https://doi.org/10.2196/34384
- Hwang et al. (2024) Angel Hsing-Chi Hwang, Dan Adler, Meir Friedenberg, and Qian Yang. 2024. Societal-Scale Human-AI Interaction Design? How Hospitals and Companies Are Integrating Pervasive Sensing into Mental Healthcare. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3613904.3642793
- Jones et al. (2012) Spencer Jones, Paul Heaton, Robert Rudin, and Eric Schneider. 2012. Unraveling the IT Productivity Paradox - Lessons for Health Care. The New England journal of medicine 366 (June 2012), 2243–5. https://doi.org/10.1056/NEJMp1204980
- Jones et al. (2021) Simon L. Jones, William Hue, Ryan M. Kelly, Rosemarie Barnett, Violet Henderson, and Raj Sengupta. 2021. Determinants of Longitudinal Adherence in Smartphone-Based Self-Tracking for Chronic Health Conditions: Evidence from Axial Spondyloarthritis. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 5, 1 (March 2021), 16:1–16:24. https://doi.org/10.1145/3448093
- Kabir and Wiese (2023) Kazi Sinthia Kabir and Jason Wiese. 2023. A Meta-Synthesis of the Barriers and Facilitators for Personal Informatics Systems. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 7, 3 (Sept. 2023), 103:1–103:35. https://doi.org/10.1145/3610893
- Kang and Exworthy (2022) Harjeevan Singh Kang and Mark Exworthy. 2022. Wearing the Future—Wearables to Empower Users to Take Greater Responsibility for Their Health and Care: Scoping Review. JMIR mHealth and uHealth 10, 7 (July 2022), e35684. https://doi.org/10.2196/35684
- Kann et al. (2007) Laura Kann, Susan K. Telljohann, and Susan F. Wooley. 2007. Health Education: Results From the School Health Policies and Programs Study 2006. Journal of School Health 77, 8 (2007), 408–434. https://doi.org/10.1111/j.1746-1561.2007.00228.x
- Kapania et al. (2022) Shivani Kapania, Oliver Siy, Gabe Clapper, Azhagu Meena SP, and Nithya Sambasivan. 2022. ”Because AI Is 100% Right and Safe”: User Attitudes and Sources of AI Authority in India. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–18. https://doi.org/10.1145/3491102.3517533
- Karnatak et al. (2023) Nimisha Karnatak, Brooke Loughrin, Tiffany Amy Kuo, Odeline Mateu-Silvernail, Indrani Medhi Thies, William Thies, and Mohit Jain. 2023. “Is It Even Giving the Correct Reading or Not?”: How Trust and Relationships Mediate Blood Pressure Management in India. ACM Trans. Comput.-Hum. Interact. 30, 6 (Sept. 2023), 90:1–90:27. https://doi.org/10.1145/3609327
- Karusala et al. (2023a) Naveena Karusala, Victoria G, Shirley Yan, and Richard Anderson. 2023a. Unsettling Care Infrastructures: From the Individual to the Structural in a Digital Maternal and Child Health Intervention. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581553
- Karusala et al. (2023b) Naveena Karusala, Shirley Yan, Nupoor Rajkumar, Victoria G, and Richard Anderson. 2023b. Speculating with Care: Worker-centered Perspectives on Scale in a Chat-based Health Information Service. Proc. ACM Hum.-Comput. Interact. 7, CSCW2 (Oct. 2023), 361:1–361:26. https://doi.org/10.1145/3610210
- Kemp et al. (2021) Emma Kemp, Joshua Trigg, Lisa Beatty, Chris Christensen, Haryana M. Dhillon, Anthony Maeder, Patricia A. H. Williams, and Bogda Koczwara. 2021. Health Literacy, Digital Health Literacy and the Implementation of Digital Health Technologies in Cancer Care: The Need for a Strategic Approach. Health Promotion Journal of Australia 32, S1 (2021), 104–114. https://doi.org/10.1002/hpja.387
- Kim et al. (2022) Jeong-Kyun Kim, Da-Som Oh, Kangbok Lee, and Sang Gi Hong. 2022. Fall Detection Based on Interpretation of Important Features with Wrist-Wearable Sensors. In Proceedings of the 28th Annual International Conference on Mobile Computing And Networking (MobiCom ’22). Association for Computing Machinery, New York, NY, USA, 823–825. https://doi.org/10.1145/3495243.3558250
- Klasnja et al. (2011) Predrag Klasnja, Sunny Consolvo, and Wanda Pratt. 2011. How to Evaluate Technologies for Health Behavior Change in HCI Research. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’11). Association for Computing Machinery, New York, NY, USA, 3063–3072. https://doi.org/10.1145/1978942.1979396
- Kreps (2017) Gary L. Kreps. 2017. The Relevance of Health Literacy to mHealth. Information Services & Use 37, 2 (Jan. 2017), 123–130. https://doi.org/10.3233/ISU-170828
- Kruzan et al. (2023) Kaylee Payne Kruzan, Ada Ng, Colleen Stiles-Shields, Emily G Lattie, David C. Mohr, and Madhu Reddy. 2023. The Perceived Utility of Smartphone and Wearable Sensor Data in Digital Self-tracking Technologies for Mental Health. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3544548.3581209
- Kyytsönen et al. (2023) Maiju Kyytsönen, Tuulikki Vehko, Heidi Anttila, and Jonna Ikonen. 2023. Factors Associated with Use of Wearable Technology to Support Activity, Well-Being, or a Healthy Lifestyle in the Adult Population and among Older Adults. PLOS Digital Health 2, 5 (May 2023), e0000245. https://doi.org/10.1371/journal.pdig.0000245
- Lazar et al. (2015) Amanda Lazar, Christian Koehler, Theresa Jean Tanenbaum, and David H. Nguyen. 2015. Why We Use and Abandon Smart Devices. In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (UbiComp ’15). Association for Computing Machinery, New York, NY, USA, 635–646. https://doi.org/10.1145/2750858.2804288
- Lee et al. (2024a) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee. 2024a. FamilyScope: Visualizing Affective Aspects of Family Social Interactions Using Passive Sensor Data. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 57:1–57:27. https://doi.org/10.1145/3637334
- Lee et al. (2024b) Jarrett G.W. Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe. 2024b. Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 8, 2 (May 2024), 51:1–51:21. https://doi.org/10.1145/3659585
- Lee et al. (2022) Wan Ling Lee, Zi Jing Lim, Li Yoong Tang, Nor Aziyan Yahya, Kasturi Dewi Varathan, and Salizar Mohamed Ludin. 2022. Patients’ Technology Readiness and eHealth Literacy: Implications for Adoption and Deployment of eHealth in the COVID-19 Era and Beyond. CIN: Computers, Informatics, Nursing 40, 4 (April 2022), 244. https://doi.org/10.1097/CIN.0000000000000854
- Lieber et al. (2020) Judith Lieber, Lynda Clarke, Ian M. Timæus, Poppy Alice Carson Mallinson, and Sanjay Kinra. 2020. Changing Family Structures and Self-Rated Health of India’s Older Population (1995-96 to 2014). SSM - Population Health 11 (Aug. 2020), 15. https://doi.org/10.1016/j.ssmph.2020.100572
- Lu et al. (2024) Yuqian Lu, Zihan Zhou, and Xuelian Li. 2024. Study on the Relationship between Technology, Data and Body Isomorphism—The Example of Smartwatches. Telematics and Informatics Reports 13 (March 2024), 100118. https://doi.org/10.1016/j.teler.2024.100118
- Luz and Masoodian (2024) Saturnino Luz and Masood Masoodian. 2024. Managing Personal Health and Well-being by Integrating Data Streams Using Interactive Visualizations. In Proceedings of the 2024 International Conference on Advanced Visual Interfaces (AVI ’24). Association for Computing Machinery, New York, NY, USA, 1–3. https://doi.org/10.1145/3656650.3656743
- Ma and Ren (2023) Yumeng Ma and Jiahao Ren. 2023. ProactiveAgent: Personalized Context-Aware Reminder System. In Adjunct Proceedings of the 36th Annual ACM Symposium on User Interface Software and Technology (UIST ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, 1–3. https://doi.org/10.1145/3586182.3625115
- Mackert et al. (2016) Michael Mackert, Amanda Mabry-Flynn, Sara Champlin, Erin E Donovan, and Kathrynn Pounders. 2016. Health Literacy and Health Information Technology Adoption: The Potential for a New Digital Divide. Journal of Medical Internet Research 18, 10 (Oct. 2016), e264. https://doi.org/10.2196/jmir.6349
- Meyers et al. (2020) Nicole Meyers, Alexander F. Glick, Alan L. Mendelsohn, Ruth M. Parker, Lee M. Sanders, Michael S. Wolf, Stacy Bailey, Benard P. Dreyer, Jessica J. Velazquez, and H. Shonna Yin. 2020. Parents’ Use of Technologies for Health Management: A Health Literacy Perspective. Academic Pediatrics 20, 1 (Jan. 2020), 23–30. https://doi.org/10.1016/j.acap.2019.01.008
- Misra et al. (2023) Saumya Misra, Rachana Adtani, Yuvraj Singh, Simran Singh, and Devanshu Thakkar. 2023. Exploring the Factors Affecting Behavioral Intention to Adopt Wearable Devices. Clinical Epidemiology and Global Health 24 (Nov. 2023), 101428. https://doi.org/10.1016/j.cegh.2023.101428
- MoHFW (2005) MoHFW. 2005. School Health Programme.
- Murnane et al. (2018) Elizabeth L. Murnane, Tara G. Walker, Beck Tench, Stephen Voida, and Jaime Snyder. 2018. Personal Informatics in Interpersonal Contexts: Towards the Design of Technology That Supports the Social Ecologies of Long-Term Mental Health Management. Proc. ACM Hum.-Comput. Interact. 2, CSCW (Nov. 2018), 127:1–127:27. https://doi.org/10.1145/3274396
- Napolitano et al. (2023) Enea Vincenzo Napolitano, Simona Fioretto, Elio Masciari, and Arianna Anniciello. 2023. How Pandemic Affected the Adoption of E-Health Systems. In Proceedings of the 27th International Database Engineered Applications Symposium (IDEAS ’23). Association for Computing Machinery, New York, NY, USA, 94–98. https://doi.org/10.1145/3589462.3589496
- Ning et al. (2023) Emma Ning, Andrea T. Cladek, Mindy K. Ross, Sarah Kabir, Amruta Barve, Ellyn Kennelly, Faraz Hussain, Jennifer Duffecy, Scott L. Langenecker, Theresa Nguyen, Theja Tulabandhula, John Zulueta, Olusola A. Ajilore, Alexander P. Demos, and Alex Leow. 2023. Smartphone-Derived Virtual Keyboard Dynamics Coupled with Accelerometer Data as a Window into Understanding Brain Health: Smartphone Keyboard and Accelerometer as Window into Brain Health. In Proceedings of the 2023 CHI Conference on Human Factors in Computing Systems (CHI ’23). Association for Computing Machinery, New York, NY, USA, 1–15. https://doi.org/10.1145/3544548.3580906
- Noblin et al. (2012) Alice M. Noblin, Thomas T. H. Wan, and Myron Fottler. 2012. The Impact of Health Literacy on a Patient’s Decision to Adopt a Personal Health Record. Perspectives in Health Information Management / AHIMA, American Health Information Management Association 9, Fall (Oct. 2012), 1e.
- Nunes et al. (2015) Francisco Nunes, Nervo Verdezoto, Geraldine Fitzpatrick, Morten Kyng, Erik Grönvall, and Cristiano Storni. 2015. Self-Care Technologies in HCI: Trends, Tensions, and Opportunities. ACM Trans. Comput.-Hum. Interact. 22, 6 (Dec. 2015), 33:1–33:45. https://doi.org/10.1145/2803173
- Okolo et al. (2024) Chinasa T. Okolo, Dhruv Agarwal, Nicola Dell, and Aditya Vashistha. 2024. ”If It Is Easy to Understand Then It Will Have Value”: Examining Perceptions of Explainable AI with Community Health Workers in Rural India. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 71:1–71:28. https://doi.org/10.1145/3637348
- Pai and Alathur (2020) Rajesh R. Pai and Sreejith Alathur. 2020. Determinants of Mobile Health Application Awareness and Use in India: An Empirical Analysis. In Proceedings of the 13th International Conference on Theory and Practice of Electronic Governance (ICEGOV ’20). Association for Computing Machinery, New York, NY, USA, 576–584. https://doi.org/10.1145/3428502.3428589
- Pandya et al. (2020) Bhavesh Pandya, Amir Pourabdollah, and Ahmad Lotfi. 2020. Fuzzy Logic Web Services for Real-Time Fall Detection Using Wearable Accelerometer and Gyroscope Sensors. In Proceedings of the 13th ACM International Conference on PErvasive Technologies Related to Assistive Environments (PETRA ’20). Association for Computing Machinery, New York, NY, USA, 1–7. https://doi.org/10.1145/3389189.3397989
- Pasi (2024) Suyash Pasi. 2024. The Impact of Medical Misinformation on Health Rights in India. https://www.humanrightsresearch.org/post/the-impact-of-medical-misinformation-on-health-rights-in-india.
- Patil et al. (2022) Vathsala Patil, Deepak Kumar Singhal, Nithesh Naik, B. M. Zeeshan Hameed, Milap J. Shah, Sufyan Ibrahim, Komal Smriti, Gaurav Chatterjee, Ameya Kale, Anshika Sharma, Rahul Paul, Piotr Chłosta, and Bhaskar K. Somani. 2022. Factors Affecting the Usage of Wearable Device Technology for Healthcare among Indian Adults: A Cross-Sectional Study. Journal of Clinical Medicine 11, 23 (Nov. 2022), 7019. https://doi.org/10.3390/jcm11237019
- Poole (2013) Erika S. Poole. 2013. HCI and Mobile Health Interventions. Translational Behavioral Medicine 3, 4 (Dec. 2013), 402–405. https://doi.org/10.1007/s13142-013-0214-3
- Pulimeno et al. (2020) Manuela Pulimeno, Prisco Piscitelli, Salvatore Colazzo, Annamaria Colao, and Alessandro Miani. 2020. School as Ideal Setting to Promote Health and Wellbeing among Young People. Health Promotion Perspectives 10, 4 (Nov. 2020), 316–324. https://doi.org/10.34172/hpp.2020.50
- Rapp and Cena (2016) Amon Rapp and Federica Cena. 2016. Personal Informatics for Everyday Life: How Users without Prior Self-Tracking Experience Engage with Personal Data. International Journal of Human-Computer Studies 94 (Oct. 2016), 1–17. https://doi.org/10.1016/j.ijhcs.2016.05.006
- Reis et al. (2024) Moritz Reis, Florian Reis, and Wilfried Kunde. 2024. Influence of Believed AI Involvement on the Perception of Digital Medical Advice. Nature Medicine 30, 7 (July 2024), 1–3. https://doi.org/10.1038/s41591-024-03180-7
- Rooksby et al. (2014) John Rooksby, Mattias Rost, Alistair Morrison, and Matthew Chalmers. 2014. Personal Tracking as Lived Informatics. In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (CHI ’14). Association for Computing Machinery, New York, NY, USA, 1163–1172. https://doi.org/10.1145/2556288.2557039
- Saleem and Jan (2021) Sheikh Mohd Saleem and Shah Sumaya Jan. 2021. Modified Kuppuswamy Socioeconomic Scale Updated for the Year 2021. Indian Journal of Forensic and Community Medicine 8, 1 (March 2021), 1–3. https://doi.org/10.18231/j.ijfcm.2021.001
- Samhale (2022) Khalid Samhale. 2022. The Impact of Trust in the Internet of Things for Health on User Engagement. Digital Business 2, 1 (March 2022), 100021. https://doi.org/10.1016/j.digbus.2022.100021
- Scroggs and Love (2022) Barrett Scroggs and Heather A. Love. 2022. Understanding Sexual and Gender Minority Substance Use through Latent Profiles of Ecological Systems. Substance Abuse 43, 1 (2022), 640–648. https://doi.org/10.1080/08897077.2021.1986769
- Seelam et al. (2024) Ananya Seelam, Arnab Paul Choudhury, Connie Liu, Miyuki Goay, Kalika Bali, and Aditya Vashistha. 2024. ”Fact-checks Are for the Top 0.1%”: Examining Reach, Awareness, and Relevance of Fact-Checking in Rural India. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 56:1–56:34. https://doi.org/10.1145/3637333
- Shajari et al. (2023) Shaghayegh Shajari, Kirankumar Kuruvinashetti, Amin Komeili, and Uttandaraman Sundararaj. 2023. The Emergence of AI-Based Wearable Sensors for Digital Health Technology: A Review. Sensors 23, 23 (Jan. 2023), 9498. https://doi.org/10.3390/s23239498
- Shin et al. (2019) Grace Shin, Mohammad Hossein Jarrahi, Yu Fei, Amir Karami, Nicci Gafinowitz, Ahjung Byun, and Xiaopeng Lu. 2019. Wearable Activity Trackers, Accuracy, Adoption, Acceptance and Health Impact: A Systematic Literature Review. Journal of Biomedical Informatics 93 (May 2019), 103153. https://doi.org/10.1016/j.jbi.2019.103153
- Singh et al. (2022a) Nidhi Singh, Richa Misra, Sonali Singh, Nripendra P. Rana, and Sangeeta Khorana. 2022a. Assessing the Factors That Influence the Adoption of Healthcare Wearables by the Older Population Using an Extended PMT Model. Technology in Society 71 (Nov. 2022), 102126. https://doi.org/10.1016/j.techsoc.2022.102126
- Singh et al. (2022b) Nihaal Singh, Rushikesh Shukla, Sourya Acharya, and Samarth Shukla. 2022b. The Past, Present, and Future of Health Economics in India. Journal of Family Medicine and Primary Care 11, 12 (Dec. 2022), 7513–7516. https://doi.org/10.4103/jfmpc.jfmpc_2266_21
- Stuart et al. (2022) Tucker Stuart, Jessica Hanna, and Philipp Gutruf. 2022. Wearable Devices for Continuous Monitoring of Biosignals: Challenges and Opportunities. APL Bioengineering 6, 2 (April 2022), 021502. https://doi.org/10.1063/5.0086935
- Su (2023) Zhaoyuan Su. 2023. Caring for Children’s Health and Wellbeing Through Understanding and Designing for Health Data Literacy. In Proceedings of the 22nd Annual ACM Interaction Design and Children Conference (IDC ’23). Association for Computing Machinery, New York, NY, USA, 785–788. https://doi.org/10.1145/3585088.3593922
- Sulz et al. (2024) Lauren Sulz, Daniel B. Robinson, Hayley Morrison, Josh Read, Ashley Johnson, Lucinda Johnston, and Kim Frail. 2024. A Scoping Review of K–12 Health Education in Canada: Understanding School Stakeholders’ Perceptions. Curriculum Studies in Health and Physical Education 0, 0 (Jan. 2024), 1–23. https://doi.org/10.1080/25742981.2024.2311113
- Sun et al. (2023) Shuyi Sun, Alejandra Ruiz, Sima Pirmoradi, and Katia Vega. 2023. BioSparks: Jewelry as Electrochemical Sweat Biosensors with Modular, Repurposing and Interchangeable Approaches. In Adjunct Proceedings of the 2023 ACM International Joint Conference on Pervasive and Ubiquitous Computing & the 2023 ACM International Symposium on Wearable Computing (UbiComp/ISWC ’23 Adjunct). Association for Computing Machinery, New York, NY, USA, 315–320. https://doi.org/10.1145/3594739.3610787
- Tachtler et al. (2021) Franziska Tachtler, Reem Talhouk, Toni Michel, Petr Slovak, and Geraldine Fitzpatrick. 2021. Unaccompanied Migrant Youth and Mental Health Technologies: A Social-Ecological Approach to Understanding and Designing. In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (CHI ’21). Association for Computing Machinery, New York, NY, USA, 1–19. https://doi.org/10.1145/3411764.3445470
- Teixeira Lopes and Henriques (2024) Carla Teixeira Lopes and Mariana Henriques. 2024. Unveiling Health Literacy through Web Search Behavior: A Classification-Based Analysis of User Interactions. In Proceedings of the 2024 Conference on Human Information Interaction and Retrieval (CHIIR ’24). Association for Computing Machinery, New York, NY, USA, 1–11. https://doi.org/10.1145/3627508.3638318
- Teixeira Lopes and Ramos (2020) Carla Teixeira Lopes and Edgar Ramos. 2020. Studying How Health Literacy Influences Attention during Online Information Seeking. In Proceedings of the 2020 Conference on Human Information Interaction and Retrieval (CHIIR ’20). Association for Computing Machinery, New York, NY, USA, 283–291. https://doi.org/10.1145/3343413.3377966
- Thakkar et al. (2022) Divy Thakkar, Azra Ismail, Pratyush Kumar, Alex Hanna, Nithya Sambasivan, and Neha Kumar. 2022. When Is Machine Learning Data Good?: Valuing in Public Health Datafication. In Proceedings of the 2022 CHI Conference on Human Factors in Computing Systems (CHI ’22). Association for Computing Machinery, New York, NY, USA, 1–16. https://doi.org/10.1145/3491102.3501868
- Thapa et al. (2023) Sanjit Thapa, Abubakar Bello, Alana Maurushat, and Farnaz Farid. 2023. Security Risks and User Perception towards Adopting Wearable Internet of Medical Things. International Journal of Environmental Research and Public Health 20, 8 (April 2023), 5519. https://doi.org/10.3390/ijerph20085519
- Tholander and Nylander (2015) Jakob Tholander and Stina Nylander. 2015. Snot, Sweat, Pain, Mud, and Snow: Performance and Experience in the Use of Sports Watches. In Proceedings of the 33rd Annual ACM Conference on Human Factors in Computing Systems (CHI ’15). Association for Computing Machinery, New York, NY, USA, 2913–2922. https://doi.org/10.1145/2702123.2702482
- Thulasi et al. (2024) K Thulasi, Sumathi Balakrishnan, Lu Yue, Rattana Monpreeyadee, Syeda Rahmathunnisa Fatima, Mohammad Izhar Shafiq Bin Rosli, Giok Qi Wei, and MB Malarvili. 2024. IoT-Enabled Wearable Prototype: Detecting Signs of Depression. In Proceedings of the 2023 7th International Conference on Computational Biology and Bioinformatics (ICCBB ’23). Association for Computing Machinery, New York, NY, USA, 78–84. https://doi.org/10.1145/3638569.3638580
- Tong et al. (2019) Catherine Tong, Matthew Craner, Matthieu Vegreville, and Nicholas D. Lane. 2019. Tracking Fatigue and Health State in Multiple Sclerosis Patients Using Connnected Wellness Devices. Proc. ACM Interact. Mob. Wearable Ubiquitous Technol. 3, 3 (Sept. 2019), 106:1–106:19. https://doi.org/10.1145/3351264
- Ugargol and Bailey (2018) Allen Prabhaker Ugargol and Ajay Bailey. 2018. Family Caregiving for Older Adults: Gendered Roles and Caregiver Burden in Emigrant Households of Kerala, India. Asian Population Studies 14, 2 (May 2018), 194–210. https://doi.org/10.1080/17441730.2017.1412593
- Upadhyay et al. (2023) Nitin Upadhyay, Aakash Kamble, and Ajinkya Navare. 2023. Virtual Healthcare in the New Normal: Indian Healthcare Consumers Adoption of Electronic Government Telemedicine Service. Government Information Quarterly 40, 2 (April 2023), 101800. https://doi.org/10.1016/j.giq.2022.101800
- Volpato et al. (2021) Lucia Volpato, María del Río Carral, Nicolas Senn, and Marie Santiago Delefosse. 2021. General Practitioners’ Perceptions of the Use of Wearable Electronic Health Monitoring Devices: Qualitative Analysis of Risks and Benefits. JMIR mHealth and uHealth 9, 8 (Aug. 2021), e23896. https://doi.org/10.2196/23896
- Wang et al. (2019) Collin Wang, Carolyn Pang, Karyn Moffatt, Rock Leung, and Joanna McGrenere. 2019. Supporting Older Adults in the Use of Smart Devices for Personal Health Management. In Proceedings of the 21st International ACM SIGACCESS Conference on Computers and Accessibility (ASSETS ’19). Association for Computing Machinery, New York, NY, USA, 615–617. https://doi.org/10.1145/3308561.3354628
- Wani et al. (2024) Asra Sakeen Wani, Ishika Joshi, Nadia Ishfaq Nahvi, and Pushpendra Singh. 2024. ”Unrest and Trauma Stays with You!”: Navigating Mental Health and Professional Service-Seeking in Kashmir. In Proceedings of the CHI Conference on Human Factors in Computing Systems (CHI ’24). Association for Computing Machinery, New York, NY, USA, 1–17. https://doi.org/10.1145/3613904.3642507
- Wisetsri and Vijai (2021) Worakamol Wisetsri and C. Vijai. 2021. Rise of Artificial Intelligence in Healthcare Startups in India. Advances in Management 14, 1 (March 2021), 48–52.
- Wozniak-O’Connor et al. (2024) Vaughan Wozniak-O’Connor, Megan C. Rose, and Deborah Lupton. 2024. Health Information Ecologies: An Arts-based and Multisensory Approach. In Proceedings of the 35th Australian Computer-Human Interaction Conference (OzCHI ’23). Association for Computing Machinery, New York, NY, USA, 706–711. https://doi.org/10.1145/3638380.3638436
- Zaher et al. (2024) Mahmoud Zaher, Nabil Eldakhly, and Yahia Hassan. 2024. The Emerging Role of Wearable Health Technologies in Proactive Disease Prevention. International Journal of Wireless and Ad Hoc Communication 8 (Jan. 2024), 40–50. https://doi.org/10.54216/IJWAC.080105
- Zegura et al. (2018) Ellen Zegura, Carl DiSalvo, and Amanda Meng. 2018. Care and the Practice of Data Science for Social Good. In Proceedings of the 1st ACM SIGCAS Conference on Computing and Sustainable Societies (COMPASS ’18). Association for Computing Machinery, New York, NY, USA, 1–9. https://doi.org/10.1145/3209811.3209877
- Zhan et al. (2024) Xiao Zhan, Noura Abdi, William Seymour, and Jose Such. 2024. Healthcare Voice AI Assistants: Factors Influencing Trust and Intention to Use. Proc. ACM Hum.-Comput. Interact. 8, CSCW1 (April 2024), 62:1–62:37. https://doi.org/10.1145/3637339
- Zhang and Shahriar (2020) Chi Zhang and Hossain Shahriar. 2020. The Adoption, Issues, and Challenges of Wearable Healthcare Technology for the Elderly. In Proceedings of the 21st Annual Conference on Information Technology Education (SIGITE ’20). Association for Computing Machinery, New York, NY, USA, 50–53. https://doi.org/10.1145/3368308.3415454
- Zhang et al. (2022) Zhenming Zhang, Enjun Xia, and Jieping Huang. 2022. Impact of the Moderating Effect of National Culture on Adoption Intention in Wearable Health Care Devices: Meta-analysis. JMIR mHealth and uHealth 10, 6 (June 2022), e30960. https://doi.org/10.2196/30960
- Zhao et al. (2024) Yuchen Zhao, Tulika Banerjee, Na Liu, and Jennifer G Kim. 2024. Grow With Me: Exploring the Integration of Augmented Reality and Health Tracking Technologies to Promote Physical Activity. In Extended Abstracts of the 2024 CHI Conference on Human Factors in Computing Systems (CHI EA ’24). Association for Computing Machinery, New York, NY, USA, 1–9. https://doi.org/10.1145/3613905.3650907

### v2

@@
-Unpacking Personal(?!) Health Informatics: An Investigation of Awareness, Understanding, & Leveraged Utility in India
+Unpacking Personal(?!) Health Informatics for Proactive Collective Care in India
 Abstract.
-Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, holds promise for empowering individuals and transforming healthcare delivery.
-However, barriers to its adoption remain underexplored in the Indian context.
-This study investigates PHI adoption among Indian users and stakeholders using a multi-method approach.
-An awareness survey (n = 87) examined the usage of wearables and general PHI engagement, followed by semi-structured interviews (n = 22) that explored motivations, usage patterns, and health information sources.
-Qualitative analysis revealed that while PHI is valued for health monitoring and shared/collective care, its adoption is hindered by factors such as low health literacy, usability challenges, and mistrust in digital health platforms.
-Further stakeholder interviews and co-design workshops informed the development of a Figma-based prototype, which was evaluated for usability.
-Based on these findings, we offer design recommendations for an integrated, user-controlled PHI platform featuring accessible analytics and verifiable health information.
-Our insights highlight the socio-technical challenges of PHI adoption in India and underscore the need for reliable, user-centric solutions to support proactive healthcare.
+Personal Health Informatics (PHI), which leverages digital tools and information systems to support health assessment and self-care, promises more proactive, user-centered care, yet adoption and meaningful utilization barriers in India remain underexplored.
+Through a sequential mixed-methods study in urban India (Initial survey (n=87) and semi-structured interviews (n=22), follow-up survey = 116, and co-design workshops (n=6)), we surface practices, perceptions, and behaviors to identify ways PHI can be better utilized for proactive care in the Indian context.
+We find that PHI is valued for monitoring and enabling collective care; however, its adoption is constrained by low health and technology literacy, usability and integration issues, fragmented and costly technology ecosystems, and mistrust of digital health platforms.
+From triangulated evidence, we derive concrete design requirements, including user-controlled sharing, accessible analytics, and verifiable health information, and present a culturally grounded design vision for an integrated platform for collective care through design and evaluation of a figma prototype.
+The prototype evaluation provides further directions for design and development to better orient PHI for proactive care through the PHI-Proact operational map, which involves agency, elicitation, and engagement.
+Finally, using PHI-Proact, we conclude with concrete recommendations for designing and responsibly deploying PHI systems for proactive collective care in emerging contexts, which differ socially, culturally, infrastructurally, and technologically from WEIRD contexts.
 1. Introduction
-Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling active self-management.
-Globally, PHI leverages technologies—from smartphone applications to wearable devices such as smartwatches and fitness trackers—facilitating continuous health monitoring (Shajari et al., 2023).
-Factors like high health literacy, technology familiarity, and a proactive healthcare ethos drive PHI adoption in the Western context (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024).
-Emerging evidence (Chmielewski and Meyer, 2024) in the Western context suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care—seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
-In India, however, infrastructural, sociocultural, and technological challenges intersect and constrain PHI’s potential.
+Personal health informatics (PHI) involves using information systems and tools to assess health and wellbeing and holds significant promise for empowering individuals and transforming healthcare by enabling proactive self-management.
+Globally, PHI leverages technologies (from smartphone applications to wearable devices such as smartwatches, fitness trackers, and other sensors), often facilitating continuous health monitoring (Shajari et al., 2023).
+Factors such as high health literacy, familiarity with technology, and a proactive healthcare ethos drive PHI adoption in Western contexts (Kyytsönen et al., 2023; Chandrasekaran et al., 2020; Zhang and Shahriar, 2020; Huang et al., 2022; Zhan et al., 2024; Zaher et al., 2024).
+Emerging evidence (Chmielewski and Meyer, 2024) in the Western context further suggests that healthcare providers and users are beginning to view PHI as a medium for a more engaged care, seeking personal insights and opportunities to share data with healthcare providers for more proactive management.
+In India, however, infrastructural, sociocultural, and technological challenges intersect and further constrain PHI’s potential for proactive care.
 Digital health tools, including wearables, are often perceived as fashion accessories or status symbols rather than integrated components of comprehensive health management (Misra et al., 2023).
 Moreover, cultural practices, family dynamics, and skepticism towards advanced technologies, such as AI-driven solutions, further influence PHI usage and adoption (Hoque and Bao, 2015; Zhang et al., 2022; Creaser et al., 2022; Patil et al., 2022; Reis et al., 2024; Kapania et al., 2022; Wisetsri and Vijai, 2021).
-Given the cultural dynamics of the Indian context, a significant gap in its adoption and use can be filled by examining and understanding how PHI is currently integrated and utilized in daily health practices in India and shared with family and loved ones.
-To address these challenges, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach.
-We conducted a survey (n=87) to assess awareness and usage of smart wearables, followed by interviews (n=22) exploring user motivations, usage patterns, and health information sources.
-Additional interviews with key stakeholder pairs—couples, intergenerational family groups, and healthcare providers—and co-design workshops (n=12) employing card sorting examined shared care practices and data-sharing preferences.
+Given the cultural dynamics of the Indian context, a significant gap in the adoption and use of PHI can be filled by examining and understanding how it is currently integrated and utilized in daily care practices in India and shared with family and loved ones.
+Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India for general wellbeing.
+Specifically, we ask (RQ1)How (if) do Indians adopt, integrate, and perceive PHI for proactive care?
+Based on the analyses of the initial components of the study, we further ask (RQ2) How can technologies be better aligned with lived practices to support proactive collective care?
+To answer these questions, we investigated the adoption, integration, and perception of PHI among Indian users and stakeholders using a multi-method approach.
+We conducted a survey (n = 87) to assess awareness and usage of smart wearables, followed by interviews (n = 22) that explored user motivations, usage patterns, and sources of health information.
+After analyzing the collected data, our findings indicated a lack of awareness and access to tools that support their lived realities.
+To identify means to support that gap, we conducted additional interviews with key stakeholder pairs (couples, intergenerational family groups, and healthcare providers) and co-design workshops (n = 12) that employed card sorting, examining shared care practices and data-sharing preferences.
 Further, we evaluated a figma prototype to understand its usability and get design recommendations from the participants on what, how, and when they would like to share their health information/analytics with different stakeholders.
-Guided by Murnane et al.’s ecological model of PHI (Murnane et al., 2018) and Bronfenbrenner’s Ecological Systems Theory (Bronfenbrenner, 1974), we investigate both intrinsic and extrinsic factors influencing PHI adoption in India.
-Specifically, we ask: (1) How do Indians adopt, integrate, and perceive PHI in their health management?
-(2) What dictates PHI’s effective utilization and perceived utility, and how can we improve them?
-Our findings offer a deeper understanding of PHI challenges and opportunities in India and provide a Proof-of-Concept evaluation to strengthen our recommendations further.
-Through this work, we lay the groundwork for a more reliable and integrated PHI system supporting personal and shared care management, culturally appropriate to the Indian context and the global south.
+Our findings provide a deeper understanding of PHI challenges and opportunities for proactive collective care in India, offering a Proof-of-Concept evaluation to further strengthen our recommendations for designers and developers.
+Through this work, we lay the groundwork for a more reliable and integrated PHI system that supports both personal and shared care practices and is culturally appropriate for non-WEIRD contexts, such as India.
 2. Related Work
-The body of work on personal health informatics within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019).
+The body of work on PHI within the broader HCI community, though established, has seen significant expansion in recent years, driven by the rising prevalence of chronic diseases associated with sedentary lifestyles and advancements in digital health technologies (Anderson and Durstine, 2019).
 These factors have propelled research toward innovative health management solutions and broadened the adoption of PHI systems.
-The rapid increase in chronic health conditions and an expanding digital landscape make personal health informatics especially critical in India.
-This work aims to contextualize the evolution of PHI within the Indian healthcare landscape, addressing its unique challenges and opportunities.
-This section examines the existing scholarship through two critical lenses: (1) personal health informatics in HCI and digital health tracking and (2) the barriers and enablers of PHI in the Indian context.
-These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural health practices.
-2.1. Personal Health Informatics in HCI
-Personal health informatics (PHI) within HCI has evolved from individual tracking (Rooksby et al., 2014) to encompassing broader ecological models (Epstein et al., 2015) that account for social, structural, and cultural factors.
-Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts.
-These frameworks have been extensively applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022), emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing.
-Wearable technologies and smartphone applications are at the forefront of PHI in various contexts, like sports, managing chronic conditions, and broader long-term health monitoring initiatives (Tholander and Nylander, 2015; Klasnja et al., 2011; Poole, 2013; Nunes et al., 2015; Zegura et al., 2018; Blandford, 2019; Gooch et al., 2021; Fritz et al., 2014).
-Prior research on wearable use for PHI has also examined why people use and abandon smart devices (Lazar et al., 2015) and how it differs among people with prior self-tracking experience (Rapp and Cena, 2016).
-Studies on PHI often highlight the continuous capture and analysis of biosignals for precise intervention, particularly in chronic disease management, such as multiple sclerosis, chronic fatigue, and hypertension (Stuart et al., 2022; Jones et al., 2021; Tong et al., 2019).
-While proactive use of PHI systems exists for critical use cases, such as emergencies (Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024), fall detections(Delahoz and Labrador, 2014; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022), and medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), we are yet to see them used for a broader PHI in the Indian context.
-PHI in HCI is not limited to wearables and trackers; it also uses other modalities to integrate into an individual’s wellbeing.
-These diverse modalities represent different engagement strategies, from gamification in AR for improving physical activity (Zhao et al., 2024) using virtual pets to fostering social interaction through community displays (Gooch et al., 2021).
-Beyond physical and social wellbeing, wearables increasingly support mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023).
-Despite this, the literature (Lee et al., 2022) often emphasizes specialized healthcare needs, such as technology readiness and e-health literacy, while general proactive health tracking remains underexplored in HCI, particularly in India.
-Our work adopts an ecological perspective to examine intrinsic and extrinsic factors influencing PHI adoption, bridging the gap between research on wearables and broader proactive health management strategies.
-2.2. Barriers and Enablers of PHI in the Indian Context
+The rapid increase in chronic health conditions and an expanding digital landscape make PHI especially critical in India.
+This work aims to contextualize the evolution of PHI towards collective care, addressing its unique challenges and opportunities.
+For the entirety of this work, we define collective care as the culturally grounded, trust-based system of shared responsibility for health and wellbeing that shapes information flow, decision-making, and support within a trusted circle of stakeholders (spouse, family, friends).
+This section examines the existing scholarship through two critical lenses: (1) the evolution of modeling and use cases of PHI in HCI, and (2) the barriers and enablers of PHI for Collective Care.
+These lenses offer complementary perspectives, highlighting technological and user-centric design considerations while incorporating them into India’s sociocultural care practices.
+2.1. Evolution of Modeling and Use Cases of PHI in HCI
+Personal health informatics (PHI) models within HCI have evolved from individual-centric models (Rooksby et al., 2014; Zegura et al., 2018; Blandford, 2019; Kirchner et al., 2021) to stage-based models (Li et al., 2010) to broader models (Epstein et al., 2015; Pina et al., 2017) that account for social, structural, and cultural factors.
+While Murnane et al.’s adaptation of Bronfenbrenner’s Ecological Systems Theory (Murnane et al., 2018; Bronfenbrenner, 1974) highlights the importance of examining PHI within a network of health information ecologies, integrating formal health systems and informal social contexts for mental health, it also translates well to general health and wellbeing.
+These frameworks have been inherently or explicitly applied to domains like mental health (Tachtler et al., 2021), chronic conditions (Bhat and Kumar, 2020), PTSD/trauma (Evans et al., 2022; Wani et al., 2024), fertility (Costa Figueiredo and Chen, 2021), and substance abuse (Scroggs and Love, 2022) emphasizing how ecological approaches can inform technology design for managing physical, mental, and psychological aspects of an individual’s health and wellbeing.
+While these models support the individual even from an ecological perspective, there is a gap in understanding how all stakeholders can be supported to better utilize PHI for collective care contexts.
+Moving on to applications, studies on PHI often highlight the usefulness of continuous capture and analysis of biosignals for precise intervention, particularly for specialized use cases.
+Earlier PHI systems developed using mobile phones and specialized sensors/technologies started with supporting older adults and people with movement disabilities during emergencies and fall detection (Ketabdar and Polzehl, 2009; Ko et al., 2010; Delahoz and Labrador, 2014; Robinson et al., 2015; Gjoreski et al., 2016; Pandya et al., 2020; Bastico et al., 2022; Kim et al., 2022; Amato et al., 2022; Alrowaily, 2024; Thulasi et al., 2024; Chang et al., 2024).
+However, over time, with advancements in wearable technology, its use has been expanded to provide medical reminders (Gjoreski et al., 2016; Brun and Häkkilä, 2022; Ma and Ren, 2023; Haque et al., 2024), activity monitoring/tracking/recognition (Shen et al., 2025; Chen et al., 2024; Le et al., 2024), and tracking behavior change (Intille, 2004; Klasnja et al., 2011; Sullivan and Lachman, 2017).
+Beyond these, PHI has evolved further to support chronic conditions (Raj et al., 2019; Tong et al., 2019; Stuart et al., 2022) monitoring continuous physiological signals to detect critical changes and also for other interventions with long-term impact goals such as fitness/self-improvment (Fritz et al., 2014; Lu et al., 2024; Wang et al., 2024), sports rehabilitation (Tholander and Nylander, 2015), and workplace wellness (Chung et al., 2017).
+These prior works provide a comprehensive understanding of how to model, comprehend, and utilize PHI, primarily in Western, Educated, Industrialized, Rich, and Democratic (WEIRD) contexts.
+However, a gap still exists in mapping the factors affecting its adoption in non-WEIRD collective care contexts.
+While most of the examples given so far indicate the role of PHI through mobile and sensing-based approaches, PHI in HCI has evolved to support other modalities.
+These diverse modalities represent different engagement strategies: gamification in AR for improving physical activity (Zhao et al., 2024), utilizing virtual pets to foster social interaction through community displays (Gooch et al., 2021), and employing generative AI to provide personalized support tailored to the individual (Chopra et al., 2025).
+Beyond physical and social wellbeing, PHI (through wearables) increasingly supports mental health interventions through personalized, context-aware features (Ning et al., 2023; Kruzan et al., 2023; Jiang et al., 2023), thereby expanding the scope of PHI ever so slightly.
+Over the years, literature surveys (Nunes et al., 2015; Epstein et al., 2020; Motahar and Wiese, 2022) and case studies (Lin et al., 2025; Rey et al., 2023) have further added to this knowledge by keeping these critical factors and strategies for better utilization updated with the advances in technology.
+The work by Niess et al.
+(Niess et al., 2021) to examine beyond WEIRD user groups, and the work by Choube et al.
+on open-ended sensemaking of passive sensing (Choube et al., 2025) indicate a scope for tracking broader general wellbeing, there exists a gap in supporting various stakeholders in proactive collective care.
+Our work employs an ecological perspective to examine the intrinsic and extrinsic factors influencing PHI adoption, use, perception, and the stakeholders involved, with the aim of modeling it more effectively for proactive collective care.
+This approach bridges a gap in the research, particularly in non-Western, educated, industrialized, rich, and democratic (WEIRD) contexts, such as India.
+2.2. Barriers and Enablers of PHI for Collective Care
 Kabir and Wiese’s meta-synthesis provides a broader understanding of the general global barriers and facilitators for personal informatics systems (Kabir and Wiese, 2023).
-Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges.
-Limited digital and health literacy among marginalized populations poses significant challenges.
-Older adults often face steep learning curves when adopting digital health technologies (Cherian et al., 2024), while children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023).
-Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement (Thakkar et al., 2022; Okolo et al., 2024; Lee et al., 2024b; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap.
-Efforts to improve engagement through public health information datafication have shown promise in enhancing accessibility and comprehension (Huang et al., 2022; Pai and Alathur, 2020).
-However, implementing these strategies in India demands policy-level changes to ensure equitable access and integration with local health systems.
-National healthcare initiatives—such as Ayushman Bharat Health Account (ABHA) and expanding telemedicine services—have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023).
-Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024).
-However, sustaining long-term engagement remains challenging, as users frequently cite usability issues and information overload as key reasons for abandonment (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024).
-Apart from the Individual and Public/Infrastructural challenges, cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023).
+Furthermore, despite advancements, there is an emphasis on technology readiness and e-health literacy for better adoption and utilization (Lee et al., 2022), as limited digital and health literacy, especially among marginalized populations, pose significant challenges.
+Older adults and specially-abled Individuals (such as blind or deaf individuals) often face steep learning curves, accessibility issues, privacy concerns, and more in their journey with personal health technologies (Zhang and Shahriar, 2020; Cherian et al., 2024; Lee et al., 2024c), and this barrier is amplified due to the lack of resources to support them in non-WEIRD contexts.
+Similarly, children with ADHD struggle to interpret data from wearable devices, limiting meaningful engagement (Ankrah et al., 2023).
+Although simplified interfaces and accessible visualization techniques have been proposed to improve user engagement and support self-compassion (Thakkar et al., 2022; Okolo et al., 2024; Loerakker et al., 2024; Lee et al., 2024d; Zhan et al., 2024; Wang et al., 2019; Mackert et al., 2016), additional supporting tools remain necessary to bridge the information gap where resources, knowledge, and infrastructure constraints amplify these barriers further in non-WEIRD contexts.
+Beyond these barriers, prior research on PHI use has further examined causes for disengagement (Park and Lee, 2023), bias in PHI systems (Yfantidou et al., 2023), adoption and usability (Lu et al., 2020), the lack of ”lifestyle fit” with individuals (Yang et al., 2015), why people use and abandon smart devices for self-tracking (Lazar et al., 2015; Jones et al., 2021), how it differs among people with prior self-tracking experience (Rapp and Cena, 2016), and other workarounds/customization for contextualization (Harrison et al., 2015; Feustel et al., 2018) to engage with PHI resourcefully.
+While understanding some of these barriers is resourceful for non-WEIRD contexts as well, they primarily support an individual and not the other stakeholders involved in collective care contexts.
+Through this work, we explicitly explore India’s unique sociocultural and infrastructural challenges towards supporting various actors involved in proactive collective care.
+Beyond the individual, broader efforts to improve engagement through the use of public health information datafication have also shown promise in enhancing accessibility and comprehension of health information in critical instances (Huang et al., 2022; Pai and Alathur, 2020).
+However, shared smartphone usage (as noted by prior studies (Gupta et al., 2022; Kaur and Singh, 2025; Aneja et al., 2025) among emergent Indian users) deters its application, and policy-level changes are required to ensure equitable access and integration with local health systems.
+National healthcare initiatives, such as Ayushman Bharat Health Account (ABHA) and the expansion of telemedicine services, have gradually increased accessibility and awareness of PHI in India (Hwang et al., 2024; Napolitano et al., 2023).
+Engagement strategies through governmental schemes and NGO organizations, such as interactive, chat-based health information services and augmented displays, have shown promise in supporting medical professionals and older adults, respectively, in managing their health (Karusala et al., 2023b; Sun et al., 2023; Luz and Masoodian, 2024).
+However, sustaining long-term engagement and implementing these strategies supporting multiple stakeholders in collective care remains challenging.
+Cultural and economic factors also play a pivotal role in shaping PHI adoption in India (Chandrasekaran et al., 2020; Upadhyay et al., 2023).
 Collectivist norms and uncertainty avoidance play a significant role in shaping attitudes toward digital health technologies (Karnatak et al., 2023; Zhang et al., 2022).
 The perceived economic value of health technologies is often underappreciated unless they provide critical utility to users through financial implications, for example, in managing chronic conditions (Singh et al., 2022b).
-While trust and perception issues persist, privacy concerns, data accuracy doubts, and discomfort with wearables further inhibit adoption (Shajari et al., 2023; Seelam et al., 2024).
-While these multi-faceted barriers exist in the Indian context, our study aims to reconfirm these, identify more, and develop strategies to improve health literacy, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization.
-Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies.
+While trust and perception issues persist, privacy concerns, doubts about data accuracy, and discomfort with wearables further inhibit adoption (Seelam et al., 2024).
+While these multifaceted barriers exist in the Indian context, our study aims to reconfirm these, identify additional ones, and develop strategies to improve health literacy/overcome the literacy gap, enhance PHI understanding and adoption, and sustain engagement by examining user trust, cultural awareness, and design personalization.
+Understanding and addressing these challenges will contribute to a more inclusive and sustainable adoption of PHI technologies for proactive collective care.
 3. Methodology
-To examine the factors influencing the adoption of PHI in India, we conducted a study starting in January 2024, following approval from the Institute’s IRB.
-Our investigation began with a thorough review of the existing literature to understand the PHI usage and adoption landscape.
-It focused on the barriers and drivers for its adoption in personal health management and how its use (or the lack of it) is integrated into health management practices.
-We then designed a mixed-methods approach comprising surveys, interviews, and co-design workshops to capture the diverse perspectives of users.
-After analyzing all participant data and their inputs, we evaluated a rapid figma prototype of sharing preferences to further reinforce a unique finding.
+We adopted a sequential exploratory mixed-methods design.
+We started the study in January 2024, following approval from the Institute’s IRB.
+RQ1 (survey + interviews) was exploratory and aimed to discover how urban Indian users understand, use, and leverage PHI for proactive care.
+Themes that emerged particularly around health sharing and collective care motivated a follow-up design phase.
+To investigate design responses to these emergent needs, we conducted co-design workshops, stakeholder interviews, prototype evaluation (RQ2), and a concluding survey.
+The co-design and prototype evaluation, therefore, functioned as a targeted, theory-informed design probe used to (a) generate design concepts grounded in participants’ existing practices, and (b) collect usability and preference feedback to refine those concepts.
+The workflow of the overall study is presented in Figure 1.
 3.1. Participants
 The study recruited participants from across India, aiming to capture a diverse demographic range in age, gender, and professional background.
-To ensure a representative sample (especially those who can purchase and use even affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale by Saleem et al.
+To ensure a representative sample (especially those who can purchase and use at least affordable smartphones and/or wearables), we focused on individuals from middle to upper-middle socioeconomic strata, categorized using the updated Kuppuswamy scale (the latest peer-reviewed SES scale widely used for India) by Saleem et al.
 (Saleem and Jan, 2021).
 We employed a combination of purposive and snowball sampling methods to identify and engage participants.
-The sample included a variety of professions, such as IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society.
+The sample comprised a diverse range of professions, including IT professionals, healthcare workers, students, entrepreneurs, and retirees, reflecting the broad spectrum of urban Indian society.
 This approach enabled us to gather insights from participants with differing levels of familiarity and engagement with wearable health technologies.
@@
 Demographic details for the surveys and interview participants are summarized in Table 1, offering a comprehensive overview of the sample’s composition.
-Study Component | Total | Gender | Age | Education | |||||||
-Male | Female | Prefer Not | Range | Mean | High | Bachelors | Advanced | ||||
-to Say | School | Degrees | |||||||||
-PHI awareness survey | 87 | 50 | 36 | 1 | 18-83 | 32 | 6 | 31 | 50 | ||
-34 | 20 | 14 | 0 | 18-62 | 32 | 1 | 16 | 17 | ||
-interviews | 22 | 16 | 8 | 0 | 18-62 | 28 | 5 | 6 | 11 | ||
-co-design workshops | 12 | 6 | 6 | 0 | 27-40 | 33 | 0 | 8 | 4 | ||
-114 | 44 | 70 | 0 | 19-75 | 34 | 3 | 57 | 54 |
-[The table listing participant statistics from various study components]Table 1 lists the ranges and mean of age, gender, and educational background for the three components of the study.
+Study Component Total Gender Age Education Male Female Prefer Not Range Mean High Bachelors Advanced to Say School Degrees Survey (RQ1) 87 50 36 1 18-83 32 6 31 50 open-ended questions (RQ1) 34 20 14 0 18-62 32 1 16 17 Interviews (RQ1) 22 16 8 0 18-62 28 5 6 11 co-design workshops (RQ2) 12 6 6 0 27-40 33 0 8 4 Survey (RQ2) 114 44 70 0 19-75 34 3 57 54
 3.2. Study Design
-3.2.1. Survey Design
-The initial PHI awareness survey was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants.
-The survey included closed-ended and open-ended questions designed to capture a broad spectrum of wearable health technology use data.
+3.2.1. Survey Design (RQ1)
+The initial survey (examining RQ1) was administered using SurveyMonkey and targeted a diverse demographic of urban Indian participants.
+The survey included both closed-ended and open-ended questions, designed to capture a broad spectrum of data on wearable health technology use.
 The survey comprised a total of 29 questions, including 8 questions on demographic information and 8 open-ended questions (6 for users and 2 for non-users).
@@
 (4) Has the history from your health wearable ever helped you with doctor visits?
-The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives.
+The detailed survey questionnaire is presented in Appendix B.
+The open-ended questions of the survey allowed participants to elaborate on their experiences with wearable devices as part of their PHI, providing deeper insights into their motivations, the challenges they faced, and their overall perceptions of the utility of these devices in their daily lives for proactive care.
 These responses were crucial for identifying recurring themes and patterns in the qualitative analysis.
 The survey remained open until saturation was observed in the responses.
-The PHI awareness survey had a total of 99 responses, of which 87 were complete and were used for further analysis.
-The demographics had a mean age of 32 and a maximum age of 83, which included 50 males, 36 females, and 1 participant who preferred not to disclose their gender.
+The initial survey (RQ1) had a total of 99 responses, of which 87 were complete and were used for further analysis.
+The demographics consisted of a mean age of 32 and a maximum age of 83, with 50 males, 36 females, and 1 participant who preferred not to disclose their gender.
 Among the complete responders, there were 55 users and 32 non-users.
 The majority (85%) of the survey participants were between 18 and 44 years old.
-The diversity of participants in age, gender, profession, and education provided a robust dataset for analyzing broader adoption/use/behavior patterns and informed the qualitative exploration in the subsequent interviews.
-The concluding survey on health-sharing features was designed to assess both awareness and utilization of these features among Android and iOS users.
-It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each subgroup (Apple users, Apple non-users, Android users, Android non-users).
-Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions.
-The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and—when applicable—the frequency of their usage.
-Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses.
-The survey was closed upon reaching saturation of responses.
-The health sharing feature survey was completed by 114 participants, with a mean age of 34 and a maximum age of 75, which included 44 males and 70 females.
-Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed.
-This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms, especially with the Android users (who are % of users in India).
-ID | Age | Gender | Education | Wearable Usage | Was Gifted | Abandonment | Health Analytics |
-P1 | 19 | Male | High School | Health | Yes | Still Using | Advanced |
-P2 | 25 | Female | Advanced Degree | Non-User | N/A | N/A | N/A |
-P3 | 33 | Female | Advanced Degree | Health | Yes | Still Using | Advanced |
-P4 | 33 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
-P5 | 29 | Female | Bachelors | Utility | No | Still Using | N/A |
-P6 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
-P7 | 32 | Male | Doctorate | Health/Utility | No | Still Using | Basic |
-P8 | 26 | Male | Advanced Degree | Health | No | Still Using | Basic |
-P9 | 29 | Female | Bachelors | Health/Utility | Yes | Stopped Using | Advanced |
-P10 | 62 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
-P11 | 34 | Male | Bachelors | Health/Utility | No | Stopped Using | Advanced |
-P12 | 20 | Male | High School | Fashion | No | Stopped Using | N/A |
-P13 | 29 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Advanced |
-P14 | 20 | Male | High School | Non-User | N/A | N/A | N/A |
-P15 | 25 | Female | Advanced Degree | Fashion/Health | Yes | Still Using* | Basic |
-P16 | 26 | Female | Advanced Degree | Health/Fashion | No | Still Using | Basic |
-P17 | 26 | Male | Advanced Degree | Health/Utility | Yes | Still Using | Basic |
-P18 | 18 | Male | High School | Health/Utility | No | Still Using | Advanced |
-P19 | 36 | Male | Advanced Degree | Health/Utility | No | Stopped Using** | Advanced |
-P20 | 29 | Male | Advanced Degree | Non-User | N/A | N/A | N/A |
-P21 | 27 | Male | Bachelors | Health/Utility | Yes | Still Using | Basic |
-P22 | 27 | Female | Advanced Degree | Health/Utility | Yes | Stopped Using | Basic |
-[The table listing participant demographics]Table 2 lists details of age, gender, education, usage/non-usage, level of health analytics, and Usage Abandonment for each participant by the uniquely assigned participant ID from the semi-structured Interviews.
+The diversity of participants, in terms of age, gender, profession, and education, provided a robust dataset for analyzing broader adoption/use/behavior patterns, and informed the qualitative exploration in the subsequent interviews.
+ID Age Gender Education Wearable Usage Was Gifted Abandonment Health Analytics P1 19 Male High School Health Yes Still Using Advanced P2 25 Female Advanced Degree Non-User N/A N/A N/A P3 33 Female Advanced Degree Health Yes Still Using Advanced P4 33 Female Bachelors Health/Utility Yes Stopped Using Advanced P5 29 Female Bachelors Utility No Still Using N/A P6 20 Male High School Fashion No Stopped Using N/A P7 32 Male Doctorate Health/Utility No Still Using Basic P8 26 Male Advanced Degree Health No Still Using Basic P9 29 Female Bachelors Health/Utility Yes Stopped Using Advanced P10 62 Male Bachelors Health/Utility Yes Still Using Basic P11 34 Male Bachelors Health/Utility No Stopped Using Advanced P12 20 Male High School Fashion No Stopped Using N/A P13 29 Male Advanced Degree Health/Utility Yes Still Using Advanced P14 20 Male High School Non-User N/A N/A N/A P15 25 Female Advanced Degree Fashion/Health Yes Still Using* Basic P16 26 Female Advanced Degree Health/Fashion No Still Using Basic P17 26 Male Advanced Degree Health/Utility Yes Still Using Basic P18 18 Male High School Health/Utility No Still Using Advanced P19 36 Male Advanced Degree Health/Utility No Stopped Using** Advanced P20 29 Male Advanced Degree Non-User N/A N/A N/A P21 27 Male Bachelors Health/Utility Yes Still Using Basic P22 27 Female Advanced Degree Health/Utility Yes Stopped Using Basic
 3.2.2. Interview Design
 The secondary component consisted of semi-structured interviews with 22 participants, all of whom were selected from the survey pool.
-The interviews were conducted based on the availability of the participants and were conducted until response saturation was reached on the emerging themes.
-Participants were chosen to reflect on various experiences with wearable technology for their PHI.
-The interview group included 16 males and 8 females, with a mean age of 28, from diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees.
+The interviews were conducted based on the availability of the participants and continued until response saturation was reached regarding the emerging themes.
+Participants were selected to reflect on their diverse experiences with wearable technology for their PHI.
+The interview group consisted of 16 males and 8 females, with a mean age of 28, representing diverse educational backgrounds, including those with up to high school education, bachelor’s degrees, and advanced degrees.
 The detailed demographics are presented in Table 2.
-The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or stopped using their wearable devices.
+The interview group included twelve participants who actively tracked their health using wearables; the rest had never used a smart wearable for PHI or had stopped using their wearable devices.
 This diversity in the interview sample allowed for a comprehensive exploration of the different attitudes and behaviors toward wearable health technologies for PHI.
-The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables.
-The interviews also explored participants’ broader health management practices and how wearables for PHI fit within this ”ecology of tracking.” This approach allowed the researchers to capture rich, contextual data that provided insights beyond what could be obtained from the survey alone.
-ID | Age | Gender | Education | No. of Children |
+The interview questions were designed to delve deeper into themes that emerged from the survey data, such as the influence of cultural practices on wearable adoption, perceptions of data accuracy, and usability challenges associated with wearables for proactive care.
+While the initial interview guide was limited to users, based on reviewing the survey responses, we reoriented the study design to provide a deeper examination of the non-user perspective.
+This provided us with vital information on the factors that prevented them from becoming users, some of which helped us understand how PHI can be better aligned.
+The interviews also explored participants’ broader health management practices and how PHI usage for proactive care fit within this ”ecology of tracking.” This approach enabled the researchers to capture rich, contextual data, providing insights that went beyond what could be obtained from the survey alone.
+The detailed interview guide is presented in Appendix C.
+| ID | Age | Gender | Education | Child(ren) |
 | CP1 | 34 | Female | Advanced Degree | 0 |
@@
 | CP12 | 40 | Male | Bachelors | 2 |
-[The table listing participant demographics of card sorting activity.]Table 3 lists details of age, gender, education, and number of children for each participant by the uniquely assigned participant ID from the co-design workshops.Advanced degree includes graduate degrees (like M.A., L.L.M., M.Sc., MS, MTech., M.Eng., MBA, CA, MD, MPhil, PhD, DPhil, etc.)
-[Sample Card Sorting Activity Screens.]The screenshot shows the interface of the card sorting activity used by couples to group cards comprising of different health data (such as heart rate, SpO2, menstrual cycles, etc.) and aspects of sharing (such as control, privacy, and storage, etc.) to be sorted into categories of various stakeholder groups (such as parents, siblings, friends, doctors, etc.)
-3.2.3. Co-design Workshops
-One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information.
-This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI, such as Spouses, Parents/Adult Children, and Doctors, through the co-design workshops.
-To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6), Older Parent/Adult Children pairs (4), and doctors (4 from 3 specialties).
-To further identify the sharing preferences, we conducted a card sorting activity using kartSort (just using the couples as older participants struggled with the activity in the digital format and in-person activity was unfeasible as they lived across India) where they moved around different data/information cards to different stakeholders as categories.
+3.2.3. Co-design Workshops and Stakeholder Interviews
+One of the critical findings from the surveys and interviews was the participants’ desire to share health informatics with different stakeholders and the request for reliable tools to understand health information for proactive care.
+This led us to consider ideating the interaction for sharing data with stakeholders of an individual’s PHI.
+To understand how, what, why, and when they would like to share, we conducted interviews with Spousal Pairs/Couples (6) as part of the co-design workshops.
+As part of the workshops, we conducted a card sorting activity using kartSort, where participants moved different data/information cards to various categories for different stakeholders.
 A sample screen from the card sorting activity is presented in Figure 2.
-We recruited six couples to reach saturation in understanding the variation of preferences among couples with and without a child (or children).
+The different data/information cards could be duplicated, and the couple pairs could create new categories if needed for further granularity.
+The stakeholder interviews were conducted with older parent-adult child pairs (n = 4) (to avoid digital burden and unavailability for in-person workshops) and doctors (n = 4 from 3 specialties) (as they preferred interviews over workshops).
+Details of the stakeholder interview guide are presented in Appendix D.
+The workshops were conducted until saturation for understanding the variation of preferences among couples with and without a child (or children).
 The detailed demographics of the card sorting pairs are presented in Table 3.
-[Sample Rapid Prototyping Screens.]The sample screenshot of the Figma prototype shows different screen flows in the prototype for selecting sharing preferences and approving requests for sharing among family members.
 3.2.4. Rapid Prototyping and Evaluation
-After the co-design workshops, we developed a medium fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members.
-This process helped us understand the ideal sharing preferences of individuals with various stakeholders involved.
+After the co-design workshops, we developed a medium-fidelity prototype using Figma and evaluated its interaction with sharing preferences among couples and family members.
+This process helped us understand the ideal sharing preferences of individuals, involving various stakeholders.
 Figure 3 gives a sample screenshot of the prototype.
@@
 The unanimous positive takeaway from the participants’ feedback included its simplicity, modularity, and ease of control.
-Apart from these, there were individual features that the participant appreciated and provided further recommendations, which are discussed in the findings of this work.
-It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness, usage/non-usage of the same.
+Apart from these, the participant also appreciated individual features and provided further recommendations, which are discussed in the findings of this work.
+It was also noteworthy that participants from both Interviews and workshops generally lacked awareness of the health data sharing functionalities integrated into both Android and Apple smartphone applications, which indicated a need for another survey to gauge awareness and usage/non-usage of these functionalities.
 The details of those observations are also presented in the findings, and the demographics were presented in 1.
-Themes | Sub-Themes | Main Codes |
-Collective health management | Use as a care device | Tracking heart rate & step count |
-through sharing practices | Using because it was a gift | |
-Track parents/partners health metrics | ||
-Use as a health indicator | Use to monitor stress | |
-Use to check critical conditions | ||
-Step tracking as primary use | ||
-Sleep tracking awareness | ||
-Specific health tracking need | ||
-(e.g., menstrual cycle) | ||
-Digital PHI as a motivational tool | Perception on Wearable | Handy tool for tracking and utility |
-for proactive health management | Useful motivational tool | |
-Perception affecting degree of use | ||
-Perceived low utility | ||
-Trust in Data & Technology | Conditional trust in certain metrics | |
-Acceptance of minor inaccuracies | ||
-Useful as an approximate indicator | ||
-Validation through comparison | ||
-Low medical literacy on wearable data | ||
-Navigating health information | Technical and Logistical Issues | Analytics not easy to understand |
-in a fragmented digital ecosystem | Device management issues | |
-Device incompatibilities | ||
-Device change issues | ||
-Buying a wearable is an investment | ||
-Lifestyle and Physical Barriers | Lack of preferred modality of wearable | |
-Rashes from the strap | ||
-Minimal impact on doctor visits | ||
-Physical discomfort on consistent use | ||
-Preference for traditional methods | Trust on Body feeling | Will go to the doctor |
-and perception of non-necessity | if I feel something wrong | |
-Trust in health lifestyle/habits | ||
-Perceptive lack of current need | ||
-Reluctance to trust something new | ||
-Trust in Doctors and | Will choose doctor for critical issues | |
-digital PHI overwhelming | Do not know enough, doctors reliable | |
-Wearable use overwhelming | ||
-reactive care is enough | ||
-Difference between body feeling | ||
-and wearable metrics |
-[Thematic Analysis table.]Table 4 lists the main themes, sub-themes, and the main codes obtained from the reflexive thematic analysis from the interview transcripts and the field notes taken during the interviews.
+3.2.5. Survey Design (RQ2)
+The concluding survey (examining RQ2) on perception, awareness, and utilization of health-sharing features was designed to assess these features among Android and iOS users, based on a critical finding on the willingness to share PHI data among stakeholders.
+It comprised 26 items in total: five demographic questions followed by seven feature-specific questions tailored to each of the four subgroups (Apple users, non-Apple users, Android users, and non-Android users).
+Consequently, participants who actively used the relevant health application responded to six closed-ended questions and one open-ended question, while non-users completed five closed-ended questions and two open-ended questions.
+The closed-ended items measured participants’ familiarity with the sharing function within their health-tracking app, their understanding of its capabilities, and the frequency of their usage (when applicable).
+Open-ended questions invited respondents to elaborate on the reasoning behind their selected responses.
+The detailed survey questionnaire is presented in Appendix E.
+The survey was closed once the desired number of responses was reached.
+This survey (RQ2) was completed by 114 participants, with a mean age of 34 and a maximum age of 75, including 44 males and 70 females.
+Among the responders, only 20 were users of the health sharing feature, 54 were non-users, and 40 did not even know such a feature existed.
+This final survey helped reinforce the findings of the interviews and co-design workshops on the lack of such sharing mechanisms (or the lack of awareness of such features), especially among Android users (who account for % of users in India).
 3.3. Data Analysis
 We employed a combination of mixed methods to analyze the data.
-The survey responses were processed using descriptive statistics, where we calculated frequencies and percentages to assess patterns in the use and adoption of wearable health technologies for PHI, along with participants’ trust levels and perceived usefulness of these devices.
+The survey responses were analyzed using descriptive statistics, which involved calculating frequencies and percentages to identify patterns in the use and adoption of wearable health technologies for PHI, as well as participants’ trust levels and perceived usefulness of these devices.
 To explore associations between demographic characteristics (such as age and gender) and usage patterns, we conducted cross-tabulations to identify significant relationships.
-We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, and field notes for the qualitative analysis.
+We followed Braun and Clarke’s (Braun and Clarke, 2006) thematic analysis approach to comprehensively examine the interview data, open-ended survey responses, field notes, and the co-design session data for the qualitative analysis.
 The analysis proceeded in several stages.
-Initially, the interview transcripts and field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns.
+Initially, the interview transcripts and co-design field notes were reviewed multiple times to immerse the researchers in the data and identify recurring patterns.
 Subsequently, preliminary codes were developed to capture key concepts and insights.
 These codes were refined and organized into broader, more comprehensive themes.
-From this iterative process, four principal themes emerged: (1) Collective health management through sharing practices; (2) navigating health information in a fragmented digital ecosystem; (3) digital PHI as a motivational tool for proactive health management; and (4) preference for traditional methods and perception of non-necessity.
-These themes highlighted how PHI is viewed and used within the context of personal health management among urban Indian participants.
-The thematic coding framework is detailed in Table 4, providing a structured overview of the identified themes and their subcategories.
+From this iterative process, two principal themes emerged: (1) Barriers to Proactivity and (2) Enablers of Proactivity: Behaviours & Triggers.
+These themes highlighted how PHI is viewed and used within the context of proactive care among urban Indian participants.
+The thematic coding framework is detailed in Table A in the appendix, providing a structured overview of the identified themes, subthemes/codes, along with their definitions, evidence (corresponding component of the study), and design implications.
 3.4. Data Triangulation
-To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature.
+To ensure the robustness and credibility of our findings, we employed data triangulation(Carter et al., 2014) by integrating data from surveys, semi-structured interviews, co-design workshops, and existing literature towards the recommendations and discussion.
 This method enabled cross-verification and identification of patterns and discrepancies across data types, enhancing the validity of the results.
@@
 Further, we observed that participants often viewed wearables as supplementary rather than essential health tools for their PHI, reflecting a prevalent reactive approach to health management as shown by prior literature (Lee et al., 2022; Kang and Exworthy, 2022).
-By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings.
+By integrating multiple data sources, our study offers a comprehensive understanding of the adoption, use, and perceptions of wearable health technologies among urban Indian users, thereby enhancing the reliability and depth of our findings for proactive care use.
 3.5. Positionality
 The authors of this study acknowledge their positionality as researchers from the Indian cultural context.
-While some authors still live in India and bring an insider perspective, others live abroad with a background in distributed computing.
-This diverse positionality allowed for a comprehensive analysis considering local nuances and broader global trends in wearable technology use.
+While some authors still reside in India, bringing an insider’s perspective, others live abroad with families in India.
+This diverse positionality enabled a comprehensive analysis that considered both local nuances and broader global trends in the use of PHI for proactive care.
 The authors also recognize that their interpretations are influenced by their own experiences with using PHI and backgrounds, which may shape the analysis and conclusions drawn from the data.
@@
 4. Findings
-Using data triangulation from the various components of the study, we examined the perceptions, usage patterns, motivations, and barriers associated with PHI through an in-depth analysis.
-The thematic analysis revealed several key themes and sub-themes that offer valuable insights into how users interact with PHI, their trust in the technology, and their challenges/barriers to adoption.
-4.1. Collective Health Management Through Sharing Practices
-4.1.1. Health Seen as a Collective Responsibility
-While major research has focused on individualistic healthcare models, the reflection of Indian practices indicates that health and wellbeing are shared/collective in nature.
-People often view the health of their family members as a shared responsibility, driving the need to keep on top of each other’s health statuses.
-For example, in our survey, over 60% (55/87) of participants were already tracking their health using a wearable, and in the interviews, about 45% (10/22) were gifted their tracking device to be used within families to track vital metrics, enabling a collaborative care model.
-They viewed health management as a shared, collective process rather than an individual responsibility.
-They often rely on digital PHI not only to monitor their health but also to care for family members and close contacts.
-This shared approach fosters mutual support, accountability, and informed decision-making among trusted social circles.
-For instance, P8 disclosed that apart from getting a personal tracker, he also wanted to track his parents’ vitals to check on them even when living away.
-”I also got a fitbit watch for my parents, and my parents wear them. So I can track heart rate and other metrics. So first, I bought for them, and then I got one for myself.” - P8 (M, 26)
-Even among people who leave the country for work, their needs based on cultural values lead them to seek ways to track their parents’ health and support their health management remotely.
-This feeling was expressed through CP12’s personal experience of wanting to monitor his father’s health when he was living abroad a decade ago, but finding a lack of cost-effective means to do so.
-”During my initial career days (living abroad), as my dad got surgery, I wanted to be able to monitor him regularly.
-But there was no cheap and reliable means to do so.
-So, I had to rely on relatives to inform me of his status.
-If there was an economical way to monitor my father continuously, it would have helped me a lot.” - CP12 (M, 40)
-4.1.2. Understanding the Aspects of Sharing
-Through the secondary interviews, co-design sessions on sharing preferences, and a Sharing Awareness Survey, we discovered that individuals are open to sharing their PHI with spouses and doctors without reservation, but are unaware of existing mechanisms to share.
-They had varied opinions on how they would like their data to be shared and used.
-Some couples (2/6) in the co-design session were open to their data being used by others as long as they were informed of the use case, while all the rest wanted control over not only sharing but also the storing and usage of the data.
-The participants’ interest in viewing all kinds of health information about themselves and their family members in a single place showcases a need for a unified system.
-The level of openness with sharing decreased in the following order among the participants: Spouses Doctors Parents Siblings Friends Employer Insurance Companies.
-Some couples (2/6) also mentioned that they wanted to avoid sharing with employers and insurance companies if that was the option.
-”So, I am open to sharing everything with my spouse.
-Maybe just an update on how I am doing with my parents maybe.
-I don’t think I’ll be sharing much with my siblings or friends except if there are any emergency alerts.
-I am also open to sharing everything with doctors as long as it makes sense for the doctor to be seeing those for the purpose of the visit.” - CP1 (F, 34)
-When inquired about their knowledge and usage of any current sharing frameworks, most of the couples (5/6) were unaware of those already in the market through prominent brands (Apple Health and Samsung Health).
-This general lack of awareness and usage of such sharing frameworks was also reinforced through the outcome of the survey responses (94/114) from the health sharing feature survey.
-The aware couple said they were already using it (however sparingly) for things that can be tracked through their wearable.
-However, they indicated they are unhappy about the lack of control over what can be shared, how it can be shared, and when it can be shared.
-This observation hints that either the existing platforms are so rigid that users are unhappily using them as little as possible, or they are unaware of it because it is not popular in the urban Indian context.
-A couple even drew insights from their experience living abroad and the system they have seen being used.
+We examined the perceptions, usage patterns, motivations, and barriers associated with through an in-depth analysis.
+The thematic analysis revealed several key themes and sub-themes that offer valuable insights into their trust in PHI for proactive care, their barriers/enablers to proactivity, and how users interact with PHI for proactive care.
+4.1. Barriers to Proactivity
+Through the study’s various components, participants identified five recurring barriers at different levels of an individual’s ecosystem that prevent PHI from enabling earlier action and restricting its adoption or prolonged use.
+Each barrier reduces either the perceived usefulness of continuous sensing or the user’s ability to interpret and act on signals.
+4.1.1. Ecosystem fragmentation and interoperability
+At the system level, the silo of ecosystems made the process of choosing the tools for PHI arduous.
+Participants described PHI as scattered across multiple vendor apps and formats.
+Clinical documents and device data are not consolidated into a single longitudinal view.
+This fragmentation increases cognitive load and erodes the perceived value of monitoring, since users cannot easily aggregate trends or produce easy-to-read summaries.
+Several participants explicitly requested unified timelines and export templates (e.g., EHR compatibility) that would enable the actionable use of PHI in clinical encounters.
+Another byproduct of fragmentation is also the burden of purchasing wearables being an investment.
+As each ecosystems have its own compatibilities, any individual user needs to invest effort, time, and money to figure out which ones to adopt to make full use of proactive action.
+This is evident from P3’s experience, as she explains the process of buying a smartwatch as akin to a risky investment.
+If a user decides to purchase a higher-cost wearable, there is no guarantee that it will work with their smartphone, as ecosystems have unique compatibilities due to ”proprietary conditions”, and replacing wearables becomes an expensive affair.
+”It’s kind of an investment, like it is at least for me.
+So buying a watch which which will cost 20,000/25,000 at the time of the release, it is kind of an expensive affair.
+So you cannot change these things with every update.
+” - P3 (F, 33)
+Our interview data further supports this observation, as 10/22 participants only used their PHI wearable because it was gifted, and 3/7 participants who stopped using their PHI wearable did so due to incompatible ecosystems.
+As technology advances, it is clear that a unified framework for data storage and processing of PHI is possible; however, the current state of affairs indicates that factors unrelated to feasibility (market share and profitability) hinder this progress.
+Another aspect of fragmentation that affects user adoption is the lack of reliable systems supporting collective care.
+In India, care (especially for health and wellbeing) is enacted not individually or by self, but as a collective.
+This was observed from the participants’ motivation/purpose for using PHI, as noted in both the interviews and the survey.
+Even those with a focus on self-improvement did so with the need to care for their loved ones in mind.
+This further indicates that even if a unified ecosystem is developed, it will be lacking in support for care realities in India unless it encompasses all stakeholders involved.
+”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage.
+If they can be stored all together in one place, that would be a helpful feature.
+” - Survey Respondent (F, 33)
+4.1.2. Low digital health literacy
+At the individual level, a few participants found their level of digital health literacy to be a barrier to sustained usage of PHI.
+Raw metrics and clinical jargon were frequently reported as obstacles to understanding.
+Users often receive numbers without clear next steps.
+While numbers were sufficient for simpler metrics, the difficulty in taking action using derived advanced metrics proved to be a hurdle.
+Sometimes, users were not even aware of the possible features due to a lack of understanding or because the features were less popular among those without experience in self-tracking.
+Low digital health literacy, therefore, blocks interpretation and action.
+For instance, participant P10 admitted that he did not understand whether a certain value of heart rate was beneficial for him or not; however, he trusted the recommendations from doctor visits, which suggested what was good and what was observed during the visit.
+”I do not understand whether the heart rate is good for me or not, the tracker does not tell me that.
+When I had the time to go to my doctor that time also I was made to understand that it is in between 72 and 80, which was fine according to the doctor.
+Presently also, it is nearly like that.
+When I walk, it shows 78 to 82; when I am not, it shows something about 68 to 72.
+So, I am guessing it is OK, even though I am not 100% sure.” - (P10, M, 62)
+Participants asked for plain-language explanations and threshold-based “what to do” guidance tied to their readings, as P9 noted through her personal experience.
+These needs emerged repeatedly throughout our study and motivate the development of literacy-appropriate microcopy and tooltips that are actionable not only to the user but also to the stakeholders involved.
+”So if you get a chance to kind of have that information to help you navigate through that (PHI), it will be great.
+Because, at the end of the day, being healthier is everyone’s priority.
+The lack of such information or even the means to get them is stopping people (from proactive health management), with other priorities getting ahead.” - (P9, F, 29)
+4.1.3. Trust and sensor-confidence (cross-device checks)
+As the other side of the coin of digital health literacy, lack of confidence in sensors emerged as a critical barrier.
+Participants routinely cross-checked device readings and sought confirmation from clinicians or trusted experts before taking action.
+Conditioned trust was common, treating wearable outputs as indicative but not definitive.
+Some attributed this lack of inherent trust also to a lack of familiarity with using these new-age devices, as they had grown up with different devices (such as pressure cuffs, as observed by P5).
+”Maybe we are just used to a different kind of devices.
+For example, growing up, for the blood pressure machine (makes hand motions of the pumping used in a blood pressure cuff) (laughs), that thing, like you see something happening.
+(—) I think it’s one of the perceptive Issues.
+And also because I’m not used it a lot.
+I don’t have the perception, and its that.” - (P5, F, 29)
+Doubts about the accuracy of derived metrics (e.g., advanced sleep stages or spO2 estimates) led to hesitancy in taking preventive action, and participants suggested that the lack of confidence indicators or provenance metadata contributed to their reluctance to act on device signals.
+This is evident from P4’s notion of ”not blindly believing” in PHI systems when they indicate that something has gone wrong.
+”So I think, What is missing in current systems is having that authenticity or maybe accuracy?
+I don’t blindly believe them, like, OK, it’s saying something is wrong, so it must be, but I won’t ignore it either.
+If my tracker is giving some wrong data, I will check with another and see whether it’s me or it’s the sensor, but if 2 or 3 trackers have been giving you that data, I don’t think there is any need to doubt it.
+(laughs).” - (P4, F, 33)
+While this observation suggests that improving digital health literacy may overcome this barrier, the current state-of-the-art PHI systems used in Indian contexts still rely on black-box models for explaining readings and do not provide provenance for the insights delivered.
+The lack of infrastructure to support this logging, combined with the limited digital health education programs in India, further exacerbates this barrier.
+4.1.4. Device/form-factor friction
+While fragmentation, low health literacy, and a lack of trust were prevalent, physical discomfort (skin irritation), battery demands, and simple inconvenience led other participants to abandon continuous wearing.
+For example, while smartwatches are the most common form of PHI wearables, P22’s discomfort with wearables stems from her allergic reaction to straps being worn for long durations.
+While other form factors (in the form of rings and pendants) have started to emerge in the market, the lack of economic feasibility limits the utilization of such PHI systems.
+Form-factor constraints, therefore, interrupt the continuity of sensing and undermine trend detection.
+Several participants expressed a preference for alternatives (phone-based sensing, rings, or non-wrist wearables) and suggested design changes (battery-life indicators, low-sampling “always-on” modes) to reduce attrition or to continue wearing other watches (as in the case of P19).
+4.1.5. Preference for embodied and traditional cues (reactive norms)
+While the previous barriers had external influences, this one is particularly dependent on the users’ state of mind.
+A strong normative tendency to rely on felt symptoms and clinic visits reduces engagement with preventive PHI.
+For example, P20’s notion of taking medication or precaution if he felt something was wrong indicates his strong trust in self-perception about his wellbeing.
+”If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad.
+So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
+For others, embodied cues and clinician validation remain the primary triggers for action; continuous sensor outputs are viewed as supplemental rather than decisive.
+For example, P15 emphasized the need to integrate cultural cues into understanding the state of the body, rather than relying solely on continuous monitoring of vital signs.
+This cultural and behavioral preference helps explain why PHI often remains reactive in practice.
+”If you think about it, our lifestyles has been completely different from our parents.
+These (PHI) trackers only work for us.
+If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
+As Indians slowly shift towards a proactive mindset, the tools to support them are lacking, as most are designed with the mindset of supporting WEIRD contexts.
+The lack of equitable access and poor alignment with cultural and societal norms remain critical barriers to the widespread adoption of PHI for proactive collective care.
+4.2. Enablers of Proactivity: Behaviours & Triggers
+While barriers focused on why PHI was not being used effectively for proactive care, participants also highlighted several behaviors and enabling mechanisms that encourage earlier action and sustained monitoring.
+4.2.1. Stakeholder specific sharing
+In our sample, the family served as the primary channel for the introduction, mediation, and prompting of PHI.
+Younger family members often set up devices for older adults.
+For example, P8 wanted to keep an eye on his parents, who lived far away from him, and decided to monitor them through Fitbits.
+Even during the co-design workshops and stakeholder interviews, family requests or approvals were repeatedly cited as immediate triggers for checking metrics or seeking care.
+Role-based sharing (care circles with roles such as monitor, nudge, and emergency) was widely seen as useful, provided that sharing is granular and consented.
+These findings underscore the importance of family workflows as both adoption and proactive levers.
+For example, a couple (CP7 and CP8) noted that as long as they did not have to manually add every granular detail, a dedicated platform for family-based sharing was beneficial.
+”See, in this digital age, almost every data is online.
+Not only with each other but also with companies like Google or Apple (laughs).
+After this interview, I will probably get an Instagram reel on how couples should track data together (laughs).
+So, there is no need for stricter control within the couple on how/what/when to share.
+We currently use whatever means possible (mostly just WhatsApp each other on health documents and information), we are open to a dedicated platform only if we do not have to add everything, every time manually.” - CP7 (M, 33) & CP8 (F, 31)
+Beyond sharing with family (and friends in some cases), participants expressed a strong desire for concise, clinician-ready summaries that could be shown at visits.
+The prospect of clinician engagement increased the perceived credibility of PHI and would motivate sustained monitoring.
+Several respondents said they would act earlier if a clinician requested or validated specific metrics.
+In the co-design workshop, another couple (CP6 and CP5) discussed their experience from the UK, using examples from a management-centric application with their GP, and mentioned that having control over sharing summaries with stakeholders, similarly based on their needs, would improve engagement.
 ”Sharing within spouses is OK, but what about the other stakeholders like employers or insurance companies?
@@
 That level of control would be very useful.” - CP6 (M, 30) & CP5 (F, 27)
-Another insightful observation among participants interested in sharing was that people with children are interested in tracking and sharing PHI only when there is no overhead of pulling all things together by themselves.
-While the motivation to track using wearables differs individually (even within a couple), they all were motivated for the overall PHI for themselves and their children if they could do so using a single device.
-”See, in this digital age, almost every data is online.
-Not only with each other but also with companies like Google or Apple (laughs).
-After this interview, I will probably get an Instagram reel on how couples should track data together (laughs).
-So, there is no need for stricter control within the couple on how/what/when to share.
-We currently use whatever means possible (mostly just WhatsApp each other on health documents and information), we are open to a dedicated platform only if we do not have to add everything, every time manually.” - CP7 (M, 33) & CP8 (F, 31)
-After incorporating the feedback and inputs, when evaluating the rapid figma prototype, the participants liked the overall functionality of the interface and provided further suggestions.
-They felt that the prototype mimicked the discussion on their sharing preferences and appreciated the control over them.
-Some participants also suggested that further information or health tips could be added about the benefits of good sleep and (for women) steps to take when periods are late.
+While literature and current practices suggest that such sharing mechanisms exist specifically for each hospital to share information, such as clinical reports or lab test results, the ability to share PHI summaries with doctors or other critical stakeholders remains lacking.
+4.2.2. Actionable explanations and engagement
+Beyond the aspect of sharing information with stakeholders, plain-language recommendations, threshold-based guidance, and short “next-step” prompts (e.g., “re-measure/share with clinician/rest & re-check in 30 min”) were widely requested.
+Actionable microcopy converts numbers into behavior, reducing interpretation burden and increasing the likelihood of timely response.
+For example, participant P7’s experience of how notifications from his watch motivated him to walk more indicates that actionable insights are useful motivators.
+Participants welcomed light-touch social features as mechanisms to sustain routine monitoring.
+This included family challenges, badges, and progress nudges, as suggested by CP1, as well as viewing statistics on a routine basis in a graphical format, as suggested by CP3.
 ”Some family tracking and reward like a badge to the member who completes all goals, just to have encouraging competition in the family for better health would be nice.” - CP1 (F, 34)
 ”I like the daily, weekly, and monthly stats being shown for health features. But I would also like to have a graphical representation.” - (CP3, F, 33)
-The responses from the Sharing Feature Awareness Survey also indicate a need for a reliable medium of sharing information among Android users and a better demonstration of capabilities among Apple/iPhone users.
-”I usually share health reports with family members as and when we get health checkups done, however, they usually get scattered all over in phone storage.
-If they can be stored all together in one place, that would be a helpful feature.
-” - (F, 33)
-”I thought sharing was about putting it out on social media platforms. I never knew about sharing features that send health updates to our family or health care provider.” - (M, 24)
-Overall, participants’ motivations and needs indicate a need for a single dedicated application or a sharing mechanism where they can pool all kinds of health information about themselves and their trusted circles while maintaining full control over the preferences, such that they do not have additional cognitive load on managing it.
-While some prefer textual representation of this information, others are comfortable with graphical, while a very limited few are okay with numbers and medical jargon.
-4.2. Navigating Health Information in a Fragmented Digital Ecosystem
-4.2.1. Unreliable Tools and Systems
-Users encounter a patchwork of digital health information, often leaving them uncertain about what to trust.
-While basic functionalities (like step counting and heart rate monitoring) are widely accepted, there is notable skepticism around advanced metrics.
-Participants frequently cross-check data with multiple devices, healthcare professionals, and even traditional methods to validate accuracy.
-This verification process underscores the challenge of navigating a fragmented digital health landscape.
-”So I think, What is missing in current systems is having that authenticity or maybe accuracy?
-I don’t blindly believe them, like, OK, it’s saying something is wrong, so it must be, but I won’t ignore it either.
-If my tracker is giving some wrong data, I will check with another and see whether it’s me or it’s the sensor, but if 2 or 3 trackers have been giving you that data, I don’t think there is any need to doubt it.
-(laughs).” - (P4, F, 33)
-Even the participants who have used various trackers throughout their evolution noted that the current capabilities of PHI systems are not on par with their current purpose or motivation.
-They emphasized the lack of reliable sources for health information or even the capability of systems to explain their health in understandable terms.
-”So one of my first smartwatches was a Pebble smartwatch.
-It is very old.
-It started on Kickstarter (—) So that is how I got into it.
-(—) But more lately.
-It’s been to serve the purpose of.
-OK, how do I understand my body a little better?
-(—) So what it started with and what it is right now, it’s very different.
-The current systems cannot explain it well to me.” - P19 (M, 36)
-4.2.2. Island of Ecosystems
-Many of these PHI trackers have their own ecosystems that are specialized for a specific brand or product line, focused on boosting their unique functionalities to promote their line of products.
-This causes a lack of cross-functional capabilities.
-This was echoed during the interviews with doctors with different specialties, the common recurring theme was the existing use of specialized trackers for different chronic conditions like diabetes, digital pressure cuffs for hypertension, or even mood journaling for psychiatry.
-The doctors from endocrinology and medicine confirmed using existing specialized smart devices in practice, albeit their price factor limited it to a few patients.
-The high financial investment required for reliable devices further compounds these reservations for individuals as well.
-”It’s kind of an investment, like it is at least for me.
-So buying a watch which which will cost 20,000/25,000 at the time of the release, it is kind of an expensive affair.
-So you cannot change these things with every update.
-” - P3 (F, 33)
-Wearables, an integral part of PHI for many users, had technical issues apart from their limitations of cross-functionality on what they could offer the user, causing further frustrations among participants.
-One such instance is inaccurate heart rate measurements needing recalibration during exercise.
-This observation shows that mechanisms such as comparison tools or validation indicators that help users assess the accuracy of their health data are currently lacking.
-”It happens quite a lot, especially in the gym when I can feel that my heart is like pounding (laughs); like help!.
-OK, so I can feel my heart beating really fast.
-But when I see my watch, it will show readings like 75 or 76, which is kind of normal.
-(—) Then, then I have to again, again like reconfigure that (manually calibrate the reading).
-Then you put, then you press on that measure.
-So some kind of calibration is required.” - P3 (F, 33)
-Participants noted the lack of reliable health information sources for them to verify their PHI and the reduced level of interoperability among available devices.
-They expressed that they do not have verified, reputable health information sources to compare their metrics and rely on doctor visits to make sense of them.
-”I do not understand whether the heart rate is good for me or not, the tracker does not tell me that.
-When I had the time to go to my doctor that time also I was made to understand that it is in between 72 and 80, which was fine according to the doctor.
-Presently also, it is nearly like that.
-When I walk, it shows 78 to 82; when I am not, it shows something about 68 to 72.
-So, I am guessing it is OK, even though I am not 100% sure.” - (P10, M, 62)
-Overall, the observations with participants’ struggles to make sense of the data indicate that the fragmented ecosystems are complex and lack inter-compatibility to use them all cohesively for PHI.
-Simplifying the complexity of digital health data through intuitive design and user-friendly explanations that boost both digital and health literacy could aid in their journey better, as noted by a few participants.
-”So if you get a chance to kind of have that information to help you navigate through that (PHI), it will be great.
-Because, at the end of the day, being healthier is everyone’s priority.
-The lack of such information or even the means to get them is stopping people (from proactive health management), with other priorities getting ahead.” - (P9, F, 29)
-4.3. Digital PHI as a Motivational Tool for Proactive Health Management
-Digital PHI catalyzes behavior change by transforming passive data collection into proactive health management.
-Users appreciate features that provide real-time feedback, set personalized goals, and track progress over time.
-These tools encourage people to adopt healthier lifestyles, offering actionable insights and reminders.
-However, the motivational impact is closely tied to individuals’ priorities and preferences.
-For example, some participants were mindful of their lifestyle, like P16 (F, 26) and P7 (M,32), who track their steps and maintain an active lifestyle during sedentary workdays.
+Participants noted that having a healthy competition among family members would push each other to strive for a healthier lifestyle, indicating that such features serve both as positive engagement and action-triggering functions.
+They create repeated opportunities for attention and for family-led prompting that can convert sensor signals to action towards a proactive, healthy life.
+4.2.3. Proactive Monitoring vs Reactive Verifying
+With the broader analysis, participants’ engagements with PHI fell broadly along a proactive–reactive axis.
+A subset of respondents described routine, preventive behaviors such as regular step-counting (as in P16), sleep checks, or goal-driven monitoring, which we label as proactive monitoring.
 ”Since I have a desk job, so I have to make sure that I’m doing at least, like so far that I come across the Internet, they do mention that you should be at least walking 5 kilometers if you are sitting all day and working.
@@
 That’s what makes me keep a check on my step count and how much I’m walking.” - (P16, F, 26)
-”So I was walking very fast. Then it (the tracker) told me you are walking outdoors; you have walked so much distance. So then, I wish to walk more.” - (P7, M, 32)
-Integrating reminders, gamified challenges, and progress-tracking features that motivate users to stay engaged with their health data was among the recommendations from participants in the interviews and rapid prototyping.
-They considered them resourceful motivators, a means to make healthy habits, or a means for a more personalized health management tool by delivering tailored advice or actionable insights that resonate with their individual and family health goals.
-”So the (health) reminder thing, right, it can be a good guilt-tripping system so.
-(—) more like a motivator (to work out more).
-This would be a good, good thing to see, like a progress of myself.” - (P17, M, 26)
-Participants were eager to view summarized stats over day, week, and month.
-However, they also wished for them to be presented alternately to increase their engagement.
-Participants noted that being handy could also be another reason for adopting digital PHI as a reliable tool.
-This finding indicates the potential of digital PHI as a motivational tool if different presentation modalities are provided.
-4.4. Preference for Traditional Methods and Perception of Non-necessity
-4.4.1. Traditional and Cultural Influence
-Some participants remain cautious about fully adopting digital PHI, favoring traditional health management methods.
-Concerns about privacy, technological complexity, and physical discomfort (e.g., frequent charging needs and skin irritation) contribute to this resistance.
-Some users (5) trust their bodily awareness and conventional healthcare practices, finding digital tracking unnecessary or overwhelming.
-”I do not prefer to constantly check my miband.
-I only use it when I run or perform any exercise.
-I do not find the need to check the history or track, as if something is wrong, I could feel it in my body.” - P21(M,27)
-Similarly, P20 (M, 29) emphasized his belief that his body would alert him if something were wrong, reducing the perceived need for a smartwatch and confidence in his healthier lifestyle.
-”Whenever you know I’m like whenever I’m sick, (laughs) maybe at the time, something is wrong and I don’t, as you said (tracking health metrics regularly) maybe I’m not much bothered about, the specifics.
-(—) If there is something wrong with my body, maybe, I know or maybe my body will let me know if I’m feeling something bad.
-So in that case I will take medication or precaution or something, whatever that is needed at that time.” - P20 (M,29)
-Some participants even recognized and reasoned about the disjointedness between the customs of their parents’ generation and the current one, indicating a need for better integrating digital PHI with traditional cultural practices to promote active, healthy living.
-”If you think about it, our lifestyles has been completely different from our parents.
-These (PHI) trackers only work for us.
-Designers are completely ignoring the population that still leads a more self-reliant lifestyle.
-If they are able to integrate regular daily activities (like cleaning the house, washing clothes, gardening) as part of measuring other metrics for health markers, I believe it can improve the adoption and create a more proactive health management mindset.” - P15 (F,25)
-While participants who seek a more body-over-technology approach exist, they also make recommendations in ways that can improve for a wider audience.
-This observation shows that the adoption is dictated by the value it offers, which does not meet their needs, and indicates room for improvement.
-4.4.2. Barriers or Challenges Causing Digital PHI Resistance
-Participants considered an assurance of precision in both basic and advanced features a necessity to maintain trust and encourage ongoing proactive engagement.
-They compared their childhood experience with traditional measuring devices as a frame of reference for building trust in digital PHI systems, indicating a need to get used to such digital systems.
-”Maybe we are just used to a different kind of devices.
-For example, growing up, for the blood pressure machine (makes hand motions of the pumping used in a blood pressure cuff) (laughs), that thing, like you see something happening.
-(—) I think it’s one of the perceptive Issues.
-And also because I’m not used it a lot.
-I don’t have the perception, and its that.” - (P5, F, 29)
-Comfort and lifestyle fit have also emerged as critical factors influencing the continued use of smart wearables as care devices, causing the aspects of using a wearable to become a barrier for a few.
-For example, P9 (F, 29), who received a smartwatch as a gift, eventually found the device’s frequent charging needs and complexities overwhelming, not integrating seamlessly into her life.
-”So, so one, once I basically invested in this watch.
-I mean it was all fancy and all of it (—) What I realized is that this was too much of a task for me.
-The most painful thing is to kind of charge the device.
-(—) I wanted to do something to kind of keep my health in check, but it got too much for me.” - P9 (F,29)
-For others, the physical discomfort or rashes due to wearable material can detract from their overall utility, and others prefer different modalities such as rings, bands, or other wearables to a watch so they can wear their traditional watches, which they enjoy wearing.
-”For me, specifically, I do not prefer to wear it continuously because I get rashes from the strap if I wear it for too long regardless of the kind of strap.” - P22 (F,27)
-”I’m also trying to go away from a watch format and more a band or a ring or something else of that sort. (—) Mostly because I also wanna like continue wearing my other watches.” - (P19, M, 36)
-Overall, the findings indicate that digital PHI can offer options that allow for both proactive tracking and reactive monitoring, acknowledging users’ comfort with conventional health cues would be welcome among users.
-They also recommend cost-effective solutions to ease the transition from traditional to digital health management.
-5. Discussion
-Our findings shed light on how urban Indians adopt, use, and perceive health wearables for PHI in India, emphasizing the cultural, societal, and technological factors influencing user behavior.
-While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we uncover a broader gap: how these same technologies are—or could be—woven into routine health management practices.
-By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape PHI adoption.
-Below, we discuss three key themes—navigating a fragmented ecosystem, wearables as care devices, and culturally sensitive engagement —and offer design and policy recommendations to support a more culturally attuned PHI future.
-5.1. Navigating a Fragmented Digital Health Ecosystem: Technology & Policy
-India’s digital health landscape is marked by fragmentation and the absence of centralized, reliable information—conditions that contrast sharply with the integrated systems typical in Western contexts.
-This fragmentation hampers PHI usability, trust, and effectiveness.
-Participants reported difficulty navigating multiple disjointed platforms, often defaulting to traditional health practices.
-Choosing a tracker was perceived as a significant investment, complicated by non-compatible ecosystems—a clear barrier (section 4.2).
-While PHI systems evolve to support personal health management, the IT infrastructure backing them remains outdated and fragmented (Jones et al., 2012).
-Proprietary ecosystems hinder interoperability, complicating and increasing the cost of decision-making.
-For example, a Samsung smartwatch user with an iPhone or non-Samsung Android device cannot access all features or rely on data accuracy, illustrating how such silos generate user confusion and disengagement.
-This calls for design strategies that reduce the cognitive load of choice and simplify use.
-As seen in our prototype evaluation, users need flexible interaction modes and wearable form factors tailored to their comfort levels (section 4.4.2).
-While some of these challenges can be overcome through software/interaction design, the rest require policy-level changes that enforce manufacturers to comply with an open data framework that ensures accessibility of health data regardless of the medium or proprietary ecosystem requirements of individual brands/companies.
-Furthermore, the fragmented digital ecosystem and the accompanying mistrust raise ethical and privacy concerns.
-This aspect of PHI must be balanced with the need for integrated, actionable health information.
-While countries worldwide have policies regarding collecting and using health information (for example, HIPAA, PHIPA, GDPR, and APPs), India still uses the IT Act of 2000 even for medical and health information.
-The IT Act 2000 provides basic data protection by requiring data owners to be informed about its use, but it is not tailored to healthcare.
-The National Digital Health Blueprint (NDHB) recommends ISO 22600:2014 Health Informatics, but implementing these standards may be challenging for many healthcare providers in India, and ensuring compliance at the provider level could also be challenging for the government.
-Although similar healthcare-specific legislation was proposed in 2017 and 2022, it has not been implemented.
-While the Government of India has now published the “Digital Personal Data Protection Act, 2023” (DPDP 2023) covering all personal data (including health data), it remains unenforced and is not robust enough to meet the requirements posed by Indian users’ socially reflective shared caring models.
-We propose a unified, open framework that aggregates PHI, doctor visits, lab reports, medications, and more into a single interface.
-This interface should support all smart wearables—focused on Android, which dominates the Indian market (94%)—and employs accessible language to reduce learning burdens and foster proactive health behaviors.
-5.2. Wearables as Care Devices: Collective Care & Engagement Strategies
-Our participants overwhelmingly described health wearables not merely as personal gadgets but as instruments of family‐centered care (findings 4.1.1).
-Unlike the individualist framing common in Western studies (Chmielewski and Meyer, 2024), Indian users envision sharing health metrics/analysis with parents, spouses, or children, turning data streams into collaborative decision‐making tools.
-We observed that users readily assume caregiver roles—monitoring a parent’s daily steps or a child’s sleep patterns —and express enthusiasm for co‐designed interfaces that let family dyads explore and interpret data together (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023).
-To drive proactive engagement, gamification of PHI systems must capture activities familiar to Indian lifestyles—such as traditional labor and household tasks—rather than only Western‐style workouts to motivate them better (findings 4.1.2 & 4.3).
-At the same time, designs should include contextual alerts (for example, flagging sustained deviations in resting heart rate) that prompt timely medical check‐ins, moving users from reactive responses toward ongoing health management.
-Prior studies on intergenerational sharing of health data (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024a), women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a), and children-specific care for chronic ailments (Cha et al., 2024), reinforce that broadening the scope of shareable metrics can foster more comprehensive, preventive care—particularly where caregiving is a shared, familial endeavor.
-Privacy concerns were minimal among our participants (finding 4.1.2), indicating openness to data exchange within trusted circles.
-This suggests potential for paired or multi‐user wearable configurations, provided that the system offers clear, in-app explanations of how data are collected, processed, and used.
-Policymakers can foster privacy-conscious behaviors by enforcing clear regulations around data protection, specifically in PHI wearables.
-As people adopt a more proactive approach, wearables’ role in caregiving could be significantly enhanced by providing continuous insights that allow early intervention before critical incidents.
-The ability to share comprehensive health data with family members could also strengthen the perceived value of these systems, encouraging consistent use and deeper integration into daily routines.
-Designers and developers must therefore prioritize features that integrate into users’ lives, facilitating comprehensive health monitoring and sharing.
-Such usability and transparency, coupled with policy measures mandating robust PHI protections, will be essential to transforming wearables from novelty gadgets into indispensable tools for early intervention and collaborative health management.
-Our observations showcase that the ”Personal” in PHI involves more than the individual in the Indian context.
-These family-centered insights directly inform our unified PHI platform vision in the next section.
-5.3. Culturally Sensitive Engagement: Norms, Literacy, Misinformation, & Trust
-In India, paper-and-analog practices remain deeply trusted, but users embrace digital PHI when it feels like an extension, not a replacement.
-Participants, for example, described “smart” digital pressure cuffs that look and behave like traditional sphygmomanometers yet auto-sync readings to a unified PHI dashboard (findings 4.4.2).
-We frame this as a modular hybrid PHI paradigm, where users toggle between digital and analog interfaces based on context and comfort.
-Such an approach acknowledges India’s healthcare realities and offers a pragmatic, modular pathway for digital transformation—one that integrates seamlessly with generic platforms.
-In a society where talking openly about intimate or sexual health is still taboo, cultural norms around sharing vary by metric: intimate health data (e.g., sexual health) was shared only with spouses, whereas vitals like step counts traveled freely among parents, children, or extended family depending on trust (findings 4.1.1).
-Parents even leaned on devices recommended by their children, illustrating “derived trust.” To embrace these differences, we argue for a culturally informed data-governance model with user-controlled sharing settings and transparent, in-app analytics—so families can collaborate on care without over- or under-sharing.
-Future policy frameworks should incorporate these principles to foster an ecosystem where digital health innovations are ethically grounded and culturally resonant.
-Low health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) further compound these challenges and impact adoption.
-Participants distrusted novel features when metrics seemed “off” or unfamiliar/unevaluated (section 4.2.2), and authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently.
-This disparity highlights the need for authentic, reliable, vetted, user-friendly health information platforms so Indians can access trustworthy guidance without skepticism.
+Others reported episodic use of PHI, typically in response to symptoms or clinical prompts, which we label reactive use.
+These two modes coexist.
+Many participants maintained short-term, goal-oriented monitoring while reserving clinician contact for symptomatic events.
+This pattern explains why PHI often fails to translate into early clinical action despite widespread adoption of sensing devices in the sample.
+Participants described several triggers that convert passive sensing into action.
+Broadly, from the co-design sessions, social prompts (family requests and shared alerts), clinician-initiated requests, and interface-level cues (approval flows, explicit “ask to share” prompts) with granular user controls (as suggested by CP1) were repeatedly reported as immediate motivators for re-checking or seeking care.
+These observations motivated our operational focus.
+A PHI system must simultaneously enable agency (who controls the data), manage elicitation (confidence indicators and escalation paths) to shift use toward proactivity, and provide engagement (actionable explanations, social workflows).
+4.3. Prototyping & Evaluation
+Based on all the data from the previous components of the study, we designed and evaluated a medium-fidelity prototype focused on sharing preferences, family controls, and summary panels.
+Participants evaluated the prototype’s modular sharing UI, approval flows, and visualization options.
+The primary screen’s UI was designed to provide contextualized, actionable advice along with common physiological values.
+The sharing preferences were designed to be granular and customizable per stakeholder involved.
+Utilizing the critical elements discussed during the interviews and co-design workshops, which focused on the balance between control and ease of access, the buttons were designed to share health features among care circle members.
+Selected screens are shown in Figure 3.
+The takeaways from the prototype evaluation are listed in Table 4.
+| Likes | Respondents appreciated the simplicity and modularity of the “Data & Sharing” center, the ability to craft templates (e.g., “weekly summary to family; alerts to clinician only”), and the “ask to share” approval flow.
+Graphical daily/weekly/monthly summaries were widely preferred for trend recognition.
+|
+| Dislikes/change requests | Participants sought clearer, plain-language explanations (what does this sleep score mean?), clinician export templates, and EHR export options.
+They also requested confidence indicators explaining why a reading should (or should not) be trusted.
+|
+| Perceived Benefits | Several participants stated that the prototype’s granular sharing and prompt features would likely cause them to act earlier (e.g., prompting a parent to visit a clinician after a flagged trend).
+Thus, while not a field evaluation, the prototype functioned as a proof-of-concept, validating several design implications (agency controls, confidence badges, clinician summaries).
+|
+5. Discussion and Design Implications
+Our findings shed light on how urban Indians adopt, use, and perceive PHI for proactive care in India, highlighting the cultural, societal, and technological factors that influence user behavior.
+While existing literature has focused on specialized applications of wearables, such as fall detection or acute event alerts (Delahoz and Labrador, 2014; Gjoreski et al., 2016; Bat-Erdene and Saver, 2021; Amato et al., 2022; Ghosh and Das, 2024), we identify a broader gap: how these same technologies can be (or could be) integrated into routine health management practices.
+By situating our findings within both ecological systems theory (Bronfenbrenner, 1974) and Wozniak et al.’s concept of Health Information Ecologies (Wozniak-O’Connor et al., 2024), we illustrate how India’s collectivist familial networks, cultural expectations, and evolving technology landscapes jointly shape the adoption of PHI for proactive care.
+Our empirical work answers two linked questions.
+RQ1 asked how Indians adopt and perceive PHI.
+We recognize that uptake is shaped by both individual motivations (goal-driven tracking similar to western observations (Chmielewski and Meyer, 2024)) and collective caregiving practices (family-mediated monitoring and context-specific observations reflecting prior work’s findings such as supporting older adults in daily living/patients with chronic conditions/diabetic children with health management and so on (Lieber et al., 2020; Ugargol and Bailey, 2018; Gustafsson et al., 2022; Bhat et al., 2023; Cha et al., 2024)).
+Cultural norms that privilege embodied cues and clinician confirmation often mean that PHI is engaged with reactively, unless mediated by trusted social or clinical actors.
+RQ2 inquired about the conditions that promote effective use of PHI.
+We determine that effective utilization requires (a) agency: granular, transparent control of what is shared and with whom; (b) elicitation: trustworthy signal presentation (confidence badges) and multi-mode prompting strategies; and (c) engagement: literacy-appropriate explanations and family/clinician workflows that translate signals into action.
+Together, these capabilities form the operational core of PHI-Proact (a conceptual operational map for effective utilization of PHI for proactive care).
+5.1. PHI-Proact: Agency, Elicitation, & Engagement
+Through a deeper analysis of our findings, we propose PHI-Proact, an operational map that translates ecological PHI insights into prescriptive design levers for proactivity.
+PHI-Proact foregrounds three interdependent domains:
+Agency.
+Explicit, granular controls for data types, sharing recipients, temporal windows, and role templates (trusted proxies).
+Agency expands Bronfenbrenner’s (Bronfenbrenner, 1974) actor layer into concrete UI primitives that reflect family and clinician roles in India.
+Elicitation.
+Design patterns for when and how to surface signals: confidence badges, escalation paths (user → family → clinician), scheduled vs context-aware prompts, and clear action pathways.
+This domain addresses technical trustworthiness and the behavioral mechanics of prompting.
+Engagement.
+Literacy-appropriate explanations, culturally contextual motivators, family dashboards, and clinician-ready summaries that scaffold interpretation and action.
+This domain operationalizes the “support” layer of ecological models into design patterns.
+PHI-Proact therefore extends prior ecological PHI accounts by converting nested socio-technical influences into a concise set of actionable design levers and evaluation metrics suited to non-WEIRD contexts where family dynamics, device heterogeneity, and literacy constraints critically shape PHI adoption.
+Our findings confirm established HCI observations (Lazar et al., 2015; Zhang and Shahriar, 2020; Lu et al., 2024) regarding abandonment driven by usability friction and information overload, aligning with prior evidence on the importance of social support and clinician integration for sustained digital health engagement.
+At the same time, our work diverges from many Western studies that foreground the individual, quantified self-subject.
+In our Indian sample, family-mediated workflows and system-level interoperability (e.g., EHR compatibility) emerge as primary, context-specific requirements for proactivity.
+These contrasts suggest that PHI designs should prioritize collective caregiving and policy-level interoperability in non-WEIRD settings.
+5.2. Recommendations for Designers and Developers
+Based on a critical analysis using our PHI-Proact mapping, we derive ten prescriptive recommendations, each tied to our thematic findings and, where applicable, to prototype validation, and present them below in Table LABEL:Table_5.
+Each of these recommendations includes one or more of the three aspects of the PHI-Proact mapping.
+Recommendations R1, R2, R3, and R4 address various aspects of the Agency from the user’s perspective, as well as the stakeholders involved.
+Recommendations R2, R4, R5, and R8 aim to augment sensing and prompting strategies.
+Finally, R6, R7, R8, and R9 are focused on revamping engagement and support for all stakeholders involved.
+Beyond these nine recommendations, utilizing all three aspects of the PHI-Proact mapping, we also provide a final recommendation to ensure that designers and developers employ appropriate evaluation and feedback mechanisms to provide better-aligned support for proactive collective care.
+| ††: The table listing recommendations for designers and developers. | |
+| R1 | Explicit consent and granular controls: Designs should strike a balance between fine-grained controls and ease of access for the users, allowing users to specify what is shared, who receives it, and for how long, without needing to navigate dense settings.
+Our participants described care as a negotiated, family-mediated practice and highlighted the need for quick-use templates and audit logs that facilitate predictable and reversible sharing of information.
+Implementing per-metric toggles, role/situation templates (e.g., monitor/nudge/emergency (Binda et al., 2018)), and an exportable consent log addresses practical needs.
+Such controls also lower the social cost of sharing (users can hand off only what is needed to a spouse or clinician) and simplify onboarding for less technical family members.
+|
+| R2 | Multi-mode prompting and stakeholder-specific escalation: Prompting must be layered depending on the escalation of the situation.
+While gentle, scheduled reminders, context-aware nudges, and configurable escalation paths are already available for the individual, it is essential to modify these features to extend them to chosen family members and, if necessary, to clinicians.
+Participants in our workshops described social prompts and clinician requests as the most effective triggers for action.
+A system that supports quiet hours for not just the user but all stakeholders involved in their care, snooze/acknowledge responses, and an explicit escalation chain both reduces nuisance alerts and creates clear pathways to care when thresholds persist.
+Empirical work (Orzikulova et al., 2024) on adaptive JITAI-style prompting and explainable timing supports the development of escalation policies that are both behaviorally effective and intelligible to users, and it is important to extend them to other stakeholders involved.
+|
+| R3 | Privacy-preserving defaults and ethical safeguards: Systems must default to minimal sharing, offer clear and plain-language privacy summaries, provide straightforward retention controls (such as auto-delete after a specified number of days), and include revoke capabilities.
+Where PHI use carries social risk, governance by local ethics or community advisory boards should guide deployment.
+Our data show that although privacy concerns were limited and stakeholder-specific, trust concerns were a significant concern.
+Privacy-preserving defaults and visible, example-based consent dialogues reduce surprise and help users make informed choices (as observed through (Mackert et al., 2016; Meyers et al., 2020)).
+Recent research (Lee et al., 2024a) on privacy awareness tools and consent interfaces offers practical patterns for surfacing downstream uses and defaults.
+|
+| R4 | Trusted-circle workflows to overcome adoption barriers: Because family networks were a primary adoption channel in our sample, systems should provide low-friction “care circles” with role-based permissions (ensuring sensitive information is shared with appropriate stakeholders while respecting the boundaries), simple invitation and verification flows, and time- or event-limited sharing.
+Participants valued the ability to set role templates that remove the need to micromanage every share.
+Embedding onboarding that explains each role’s responsibilities further reduces confusion and social friction.
+Prior work on family fitness (Li et al., 2021), intergenerational sharing (Binda et al., 2017; Cagiltay and Mutlu, 2024; Lee et al., 2024b), and women’s data ecosystems (Grimme et al., 2024; Karusala et al., 2023a) corroborates that role-sensitive sharing affordances increase both engagement and perceived safety.
+|
+| R5 | Culturally-tailored motivation and contextual goals: Engagement mechanisms and goal framings should connect to locally meaningful activities, such as household tasks, festival periods, or family milestones, rather than focusing solely on gym-centric targets.
+Participants told us that culturally relevant framing and local language content make routine monitoring feel relevant and social.
+Embedding goal templates and coaching snippets that reference everyday practices (for example, festival-timed challenges or choreography-linked step goals) aligns PHI with lived routines and increases the likelihood that the system will be used as intended.
+Prior HCI work (Cha et al., 2025) on family and community-centered tracking supports localized, socially grounded motivators.
+|
+| R6 | Clinician-ready summaries & one-click exports: Existing sharing mechanisms from management-type applications already provide a blueprint for promoting better cohesion among stakeholders.
+To translate PHI into clinical action, provide concise, one-page clinician summaries that highlight time-windowed trends, the most recent abnormal events, and provenance notes, along with export templates (PDF/SMART-on-FHIR/HL7 snippets) that cleanly map into clinical workflows.
+Participants explicitly framed such exports as enablers of earlier care.
+Making a single “clinician view” and export button available at the point of sharing reduces friction during visits and increases the likelihood clinicians will consider PHI.
+Existing implementations (Mandl et al., 2020) and SMART/HL7 guidance show feasible export patterns for embedding wearable summaries into care.
+|
+| R7 | Supporting digital health literacy: Rather than exposing raw values and remaining black boxes (as observed in prior work (Shin et al., 2019; Zhang and Shahriar, 2020)), interfaces should translate detected patterns into concise, prescriptive “what to do” guidance, accompanied by optional clinical details, behind a progressive-disclosure affordance.
+Our respondents requested micro-guidance (e.g., “If X Y for Z days re-measure/share with clinician”) and examples that map metrics to everyday, embodied cues.
+Presenting single-sentence action steps upfront, along with a link to clinician-level notes, reduces the interpretation burden while preserving rigor for expert users.
+Prior literature on health literacy (Bagalkot et al., 2020; Su, 2023; Teixeira Lopes and Ramos, 2020; Egede et al., 2021; Teixeira Lopes and Henriques, 2024; Noblin et al., 2012; Kreps, 2017; Lee et al., 2022; Kemp et al., 2021) and rampant medical misinformation (Pasi, 2024) are critical indicators for the poor adoption of health technologies.
+HCI work (Wolf, 2019) on explainability and patient-facing explanations indicates that scenario-based, plain-language explanations materially improve comprehension and trust.
+Moreover, authoritative Indian health association articles rank poorly in search results from Indian IPs, often supplanted by vague content, whereas Western IP searches surface reliable sources prominently.
+This disparity underscores the need for authentic, reliable, and vetted user-friendly health information platforms, enabling Indians to access trustworthy guidance without hesitation.
 And since the pandemic has driven more health-conscious behaviors, implementing the long-overdue 2005 school-health curriculum recommendation (MoHFW, 2005) could foster health-conscious behavior from a young age (Kann et al., 2007; Pulimeno et al., 2020; Auld et al., 2020; Benes and Alperin, 2022; Sulz et al., 2024).
-Trust hinges on data accuracy and medical validation.
-Wearables often function as “black boxes,” limiting lay users’ comprehension (Shin et al., 2019; Zhang and Shahriar, 2020).
-Sustained adoption, therefore, requires medical validation and clinical integration.
-Devices should earn formal endorsements, sync with electronic health records, and surface sensor-accuracy indicators (e.g., confidence intervals around SpO2 or heart-rate) to bolster user confidence (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023).
-Integrating wearables with formal healthcare systems will further bolster their credibility and sustained use.
-These collective/shared care insights directly inform our integrated PHI platform that consolidates disparate data, embeds user-controlled sharing and transparent analytics, employs inclusive design methods (e.g., family workshops), and links seamlessly with clinical systems—so that as health literacy grows, users can critically weigh benefits against risks (Mackert et al., 2016; Meyers et al., 2020).
-5.4. Future Work
+|
+| R8 | Tracking Provenance for reliability: Each derived metric should surface a compact provenance and confidence statement (for example, “Estimated from wrist PPG; confidence: medium; last sync 2h ago”), so users know how much to rely on a reading and whether to corroborate it with another device or a clinician.
+Study participants described cross-checking devices before acting.
+Lightweight badges and an optional provenance modal (utilizing sampling rate, sensor type, and last calibration, among other factors) reduce the cognitive friction of validation, helping users decide when to act versus when to re-measure.
+Technical and HCI literature (Cho et al., 2021) on wearable uncertainty and person-generated data quality and literature on user perception and trust (Huhn et al., 2022; Volpato et al., 2021; Kang and Exworthy, 2022; Singh et al., 2022a; Samhale, 2022; Zhang et al., 2022; Karnatak et al., 2023; Thapa et al., 2023) supports surfacing these dimensions to users, critical for holistic support.
+|
+| R9 | Interoperability and data portability: Design the data layer around standard import/export formats (FHIR/CSV, OAuth connectors), source tagging, and deduplication/time-alignment so users and clinicians can assemble coherent longitudinal views from heterogeneous devices.
+Our participants framed ecosystem fragmentation as a barrier to meaningful histories.
+Supporting connector flows and clearly labelling provenance reduces the manual effort of assembling past data and enables portability across apps and EHRs.
+While prior work (Jones et al., 2012) indicates the IT infrastructure is still broken and fragmented, recent HCI and health-informatics work (Bathelt et al., 2022) emphasizes SMART-on-FHIR patterns for wearable integration as practical and interoperable.
+|
+| R10 | Measure proactivity—evaluation & feedback loops: Finally, teams should instrument and report operational metrics that capture proactivity, such as the Proactivity Rate (the percentage of flagged events that result in contact within 7 days), Alert-to-Action latency, clinician-confirmed early detections, and other relevant metrics.
+Our participants linked specific features (family nudges, clinician summaries, confidence badges) to earlier action.
+Tracking these metrics in pilots and using them to pilot-test design choices produces the evidence needed to iterate responsibly and demonstrate clinical value.
+Work (Cho et al., 2021) on person-generated data quality and wearable evaluation offers validated measurement constructs that can seed these dashboards.
+|
+Each recommendation is grounded in participant evidence and validated by the co-design prototype as a proof of concept (where relevant).
+This work contributes to PHI theory by translating ecological PHI (Murnane/Bronfenbrenner) into an operational PHI-for-proactivity map.
+Rather than a descriptive nesting of contexts alone, PHI-Proact prescribes concrete UI/UX and system levers (agency templates, confidence indicators, clinician export formats) that operationalize ecological constructs.
+In non-WEIRD settings, where family caregiving and systemic fragmentation are salient, this mapping demonstrates how ecological theory can directly inform design and evaluation to support proactive collective care, concretely answering our RQ2.
+5.3. Limitations and Future Work
 While there are novel recommendations on personal, public, and cultural aspects of PHI adoption, the authors acknowledge several limitations in this study.
-First, the cross-sectional design captures user behavior and perceptions at a single point in time, limiting the ability to observe how these factors may evolve.
-Additionally, the focus on urban, middle- and upper-middle-class participants restricts the generalizability of the findings to rural and lower-income populations, where different adoption barriers, such as different lifestyle practices, natural remedies, and many more, may exist.
-Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their PHI usage, utility, and adoption.
+First, our analysis draws on surveys, interviews, and co-design workshops; these methods elicit rich qualitative insights but are not a substitute for longitudinal field data, thereby limiting the ability to observe how these factors may evolve over time.
+Self-reported behaviour and intentions may not directly translate to sustained real-world behaviour change.
+As generalizability was not an objective of this work, the focus on urban, middle- and upper-middle-class participants restricts the application of the findings to other populations, where different adoption barriers may exist.
+Finally, the reliance on self-reported data also introduces potential bias, as participants may inaccurately report their use of PHI, utility, and adoption for proactive care.
+The prototype testing was qualitative and short-term, serving as a proof-of-concept for design directions rather than evidence of real-world impact.
+The prototype requires field deployment and longitudinal measurement to validate the effect of the recommended features on proactivity rate and clinical outcomes.
 Building upon our observations and the strategic direction outlined, we propose the design and development of a proof-of-concept application aimed at addressing general health, wellbeing, and medical inquiries.
@@
 Leveraging advancements in artificial intelligence, particularly finetuned open-source large language models (LLMs), the application will integrate data from clinical documents (e.g., physician reports, laboratory results, and medication records) alongside information from wearable health devices.
-This multimodal data synthesis will enable the generation of personalized, holistic insights, thereby enhancing user engagement and health awareness.
+This multimodal data synthesis will enable the generation of personalized, holistic insights (similar to the objective of Moore et al.’s work (Moore et al., 2022)), thereby enhancing user engagement and health awareness (achieving observations akin to Lin et al.
+(Lin et al., 2024)).
 The interface will employ a conversational, chatbot-style design to facilitate intuitive interaction, mediate family sharing norms, and surface literacy-appropriate explanations.
-Beyond helping users understand their personal health information (PHI), the system will also provide access to evidence-based health, medical, and wellness content.
+Beyond helping users understand their PHI, the system will also provide access to evidence-based health, medical, and wellness content.
 By sourcing validated information from reputable sources, the application seeks to address the pervasive issue of health misinformation and contribute to the improvement of public health literacy.
 6. Conclusion
-This study comprehensively explores how Indian users integrate PHI wearables for health management.
-Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of PHI in India.
-By doing so, we emphasize the influence of sociocultural factors, usability challenges, and trust issues in their consistent/long-term use of PHI for health management.
-While PHI and specifically wearables hold significant potential to shift personal health management from a reactive to a proactive model, their integration is hindered by cultural norms, limited health literacy, and skepticism around data accuracy.
-To address these challenges, we propose a dual focus on improving health literacy and building trust through transparent communication, medical validation, and user-centered design.
-Two essential contributions of this research are identifying critical areas where designers and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants.
-The critical areas include expanding PHI functionalities to align with family-based caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information.
-As shown by the evaluated prototype, the HCI community, technology developers, and policymakers must work together to standardize core functionalities and integrate PHI into formal healthcare systems.
-By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive health management, supporting both individual users and familial care structures.
-This work offers a framework for future research and industry efforts, reinforcing the transformative potential of PHI within diverse cultural contexts.
-By addressing technological and sociocultural barriers, PHI can aid in preventive healthcare, fostering long-term engagement and broader adoption across India and beyond.
+This study comprehensively explores how Indian users integrate PHI wearables for proactive care.
+Through a mixed-methods approach, we identify the barriers and opportunities for improving the adoption and sustained use of this approach in India.
+By doing so, we emphasize the influence of sociocultural factors (collective care practices), usability challenges (fragmented ecosystems), and trust issues (lack of provenance) in their consistent/long-term use of PHI for proactive care.
+Two essential contributions of this research are identifying critical areas where designers, developers, and policymakers must focus, and a qualitative evaluation of a sharing prototype among the participants.
+The critical areas include expanding PHI functionalities to align with collective caregiving, improving usability through simplified interfaces and understandable explanations, and ensuring that systems offer accurate and actionable health information.
+As demonstrated by the evaluated prototype and the recommendations, the HCI community, technology developers, and policymakers must collaborate to standardize core functionalities and integrate PHI into formal healthcare systems.
+By doing so, PHI can evolve from auxiliary support tools into indispensable mechanisms for proactive care, supporting both individual users and collective care structures.
 Acknowledgements.
-We would like to thank the student Harsh Pala for his involvement in the development of Figma prototypes used for evaluating an interactive interface of the sharing preferences identified by the card sorting activity.References
+We would like to thank student Harsh Pala for his involvement in the development of Figma prototypes used to evaluate the interactive interface of the sharing preferences identified through the card sorting activity.References
 - (1)
@@
 https://doi.org/10.1016/j.smhs.2019.08.006
+- Aneja et al.
+(2025) Urvashi Aneja, Aarushi Gupta, and Aditya Vashistha.
+2025.
+Beyond Semantics: Examining Gender Bias in LLMs Deployed within Low-resource Contexts in India.
+In Proceedings of the 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25).
+Association for Computing Machinery, New York, NY, USA, 2784–2795.
+https://doi.org/10.1145/3715275.3732180
 - Ankrah et al.
@@
 https://doi.org/10.1016/j.jstrokecerebrovasdis.2021.105826
+- Bathelt et al.
+(2022) Franziska Bathelt, Ines Reinecke, Brita Sedlmayr, Sepp Höhne, Christian Gierschner, and Martin Sedlmayr.
+2022.
+Visualization of medical wearable-data using SMART-on-FHIR–A concept for an interdisciplinary complex practical course.
+In Digital Professionalism in Health and Care: Developing the Workforce, Building the Future.
+IOS Press, Amsterdam, Netherlands, 61–65.
 - Benes and Alperin (2022) Sarah Benes and Holly Alperin. 2022. The Essentials of Teaching Health Education: Curriculum, Instruction, and Assessment. Human Kinetics, United States.
@@
 https://doi.org/10.1145/3154862.3154895
+- Binda et al.
+(2018) Jomara Binda, Chien Wen Yuan, Natalie Cope, Hyehyun Park, Eun Kyoung Choe, and John M.
+Carroll.
+2018.
+Supporting Effective Sharing of Health Information among Intergenerational Family Members.
+In Proceedings of the 12th EAI International Conference on Pervasive Computing Technologies for Healthcare (New York, NY, USA) (PervasiveHealth ’18).
+Association for Computing Machinery, New York, NY, USA, 148–157.
+https://doi.org/10.1145/3240925.3240936
 - Blandford (2019) Ann Blandford.
@@
 - Cha et al.
+(2025) Yoon Jeong Cha, Jiongyu Chen, Yasemin Gunal, Qiying Zhu, Mark W Newman, and Sun Young Park.
+2025.
+Collaborative Health-Tracking Technologies for Children and Parents: A Review of Current Studies and Directions for Future Research.
+In Proceedings of the 2025 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’25).
+Association for Computing Machinery, New York, NY, USA, Article 1038, 13 pages.
+https://doi.org/10.1145/3706598.3713596
+- Cha et al.
 (2024) Yoon Jeong Cha, Yasemin Gunal, Alice Wou, Joyce Lee, Mark W Newman, and Sun Young Park.
@@
 https://doi.org/10.2196/22443
+- Chang et al.
+(2024) Zhaoxin Chang, Fusang Zhang, Xujun Ma, Pei Wang, Weiyan Chen, Duo Zhang, Badii Jouaber, and Daqing Zhang.
+2024.
+MmECare: Enabling Fine-grained Vital Sign Monitoring for Emergency Care with Handheld MmWave Radars.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 4, Article 207 (Nov.
+2024), 24 pages.
+https://doi.org/10.1145/3699766
+- Chen et al.
+(2024) Wenqiang Chen, Jiaxuan Cheng, Leyao Wang, Wei Zhao, and Wojciech Matusik.
+2024.
+Sensor2Text: Enabling Natural Language Interactions for Daily Activity Tracking Using Wearable Sensors.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 4, Article 192 (Nov.
+2024), 26 pages.
+https://doi.org/10.1145/3699747
 - Cherian et al.
@@
 https://doi.org/10.2196/51231
+- Cho et al.
+(2021) Sylvia Cho, Chunhua Weng, Michael G Kahn, and Karthik Natarajan.
+2021.
+Identifying Data Quality Dimensions for Person-Generated Wearable Device Data: Multi-Method Study.
+JMIR Mhealth Uhealth 9, 12 (23 Dec 2021), e31618.
+https://doi.org/10.2196/31618
+- Chopra et al.
+(2025) Shaan Chopra, Katherine Juarez, James Fogarty, and Sean A.
+Munson.
+2025.
+Engagements with Generative AI and Personal Health Informatics: Opportunities for Planning, Tracking, Reflecting, and Acting around Personal Health Data.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+9, 3, Article 75 (Sept.
+2025), 33 pages.
+https://doi.org/10.1145/3749503
+- Choube et al.
+(2025) Akshat Choube, Ha Le, Jiachen Li, Kaixin Ji, Vedant Das Swain, and Varun Mishra.
+2025.
+GLOSS: Group of LLMs for Open-ended Sensemaking of Passive Sensing Data for Health and Wellbeing.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+9, 3, Article 76 (Sept.
+2025), 32 pages.
+https://doi.org/10.1145/3749474
+- Chung et al.
+(2017) Chia-Fang Chung, Nanna Gorm, Irina A.
+Shklovski, and Sean Munson.
+2017.
+Finding the Right Fit: Understanding Health Tracking in Workplace Wellness Programs.
+In Proceedings of the 2017 CHI Conference on Human Factors in Computing Systems (Denver, Colorado, USA) (CHI ’17).
+Association for Computing Machinery, New York, NY, USA, 4875–4886.
+https://doi.org/10.1145/3025453.3025510
 - Costa Figueiredo and Chen (2021) Mayara Costa Figueiredo and Yunan Chen.
@@
 (2020) Daniel A.
+Epstein, Clara Caldeira, Mayara Costa Figueiredo, Xi Lu, Lucas M.
+Silva, Lucretia Williams, Jong Ho Lee, Qingyang Li, Simran Ahuja, Qiuer Chen, Payam Dowlatyari, Craig Hilby, Sazeda Sultana, Elizabeth V.
+Eikey, and Yunan Chen.
+2020.
+Mapping and Taking Stock of the Personal Informatics Literature.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+4, 4, Article 126 (Dec.
+2020), 38 pages.
+https://doi.org/10.1145/3432231
+- Epstein et al.
+(2015) Daniel A.
 Epstein, An Ping, James Fogarty, and Sean A.
@@
 https://doi.org/10.1145/3491102.3517513
+- Feustel et al.
+(2018) Clayton Feustel, Shyamak Aggarwal, Bongshin Lee, and Lauren Wilcox.
+2018.
+People Like Me: Designing for Reflection on Aggregate Cohort Data in Personal Informatics Systems.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+2, 3, Article 107 (Sept.
+2018), 21 pages.
+https://doi.org/10.1145/3264917
 - Fritz et al.
@@
 https://doi.org/10.1145/3613904.3642851
+- Gupta et al.
+(2022) Meghna Gupta, Devansh Mehta, Anandita Punj, and Indrani Medhi Thies.
+2022.
+Sophistication with Limitation: Understanding Smartphone Usage by Emergent Users in India.
+In Proceedings of the 5th ACM SIGCAS/SIGCHI Conference on Computing and Sustainable Societies (Seattle, WA, USA) (COMPASS ’22).
+Association for Computing Machinery, New York, NY, USA, 386–400.
+https://doi.org/10.1145/3530190.3534824
 - Gustafsson et al.
@@
 https://doi.org/10.1145/3613904.3642962
+- Harrison et al.
+(2015) Daniel Harrison, Paul Marshall, Nadia Bianchi-Berthouze, and Jon Bird.
+2015.
+Activity tracking: barriers, workarounds and customisation.
+In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Osaka, Japan) (UbiComp ’15).
+Association for Computing Machinery, New York, NY, USA, 617–621.
+https://doi.org/10.1145/2750858.2805832
 - Hoque and Bao (2015) Md.
@@
 https://doi.org/10.1145/3613904.3642793
+- Intille (2004) Stephen S Intille. 2004. Ubiquitous computing technology for just-in-time motivation of behavior change. In MEDINFO 2004. IOS Press, Amsterdam, Netherlands, 1434–1437.
+- Jiang et al.
+(2023) Zhihan Jiang, Lin Lin, Xinchen Zhang, Jianduo Luan, Running Zhao, Longbiao Chen, James Lam, Ka-Man Yip, Hung-Kwan So, Wilfred H.S.
+Wong, Patrick Ip, and Edith C.H.
+Ngai.
+2023.
+A Data-Driven Context-Aware Health Inference System for Children during School Closures.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+7, 1, Article 18 (March 2023), 26 pages.
+https://doi.org/10.1145/3580800
 - Jones et al.
@@
 Wearable Ubiquitous Technol.
-5, 1 (March 2021), 16:1–16:24.
+5, 1, Article 16 (March 2021), 24 pages.
 https://doi.org/10.1145/3448093
@@
 https://doi.org/10.1145/3610210
+- Kaur and Singh (2025) Jasmeet Kaur and Pushpendra Singh.
+2025.
+Exploring the Potential of Peer Support Group for Family Planning Needs of Women in Resource-constrained Settings in India.
+Proc.
+ACM Hum.-Comput.
+Interact.
+9, 2, Article CSCW080 (May 2025), 25 pages.
+https://doi.org/10.1145/3710978
 - Kemp et al.
@@
 https://doi.org/10.1002/hpja.387
+- Ketabdar and Polzehl (2009) Hamed Ketabdar and Tim Polzehl.
+2009.
+Fall and emergency detection with mobile phones.
+In Proceedings of the 11th International ACM SIGACCESS Conference on Computers and Accessibility (Pittsburgh, Pennsylvania, USA) (Assets ’09).
+Association for Computing Machinery, New York, NY, USA, 241–242.
+https://doi.org/10.1145/1639642.1639695
 - Kim et al.
@@
 https://doi.org/10.1145/3495243.3558250
+- Kirchner et al.
+(2021) Susanne Kirchner, Jessica Schroeder, James Fogarty, and Sean A.
+Munson.
+2021.
+“They don’t always think about that”: Translational Needs in the Design of Personal Health Informatics Applications.
+In Proceedings of the 2021 CHI Conference on Human Factors in Computing Systems (Yokohama, Japan) (CHI ’21).
+Association for Computing Machinery, New York, NY, USA, Article 210, 16 pages.
+https://doi.org/10.1145/3411764.3445587
 - Klasnja et al.
@@
 https://doi.org/10.1145/1978942.1979396
+- Ko et al.
+(2010) Jeonggil Ko, Jong Hyun Lim, Yin Chen, Rvăzvan Musvaloiu-E, Andreas Terzis, Gerald M.
+Masson, Tia Gao, Walt Destler, Leo Selavo, and Richard P.
+Dutton.
+2010.
+MEDiSN: Medical emergency detection in sensor networks.
+ACM Trans.
+Embed.
+Comput.
+Syst.
+10, 1, Article 11 (Aug.
+2010), 29 pages.
+https://doi.org/10.1145/1814539.1814550
 - Kreps (2017) Gary L. Kreps. 2017. The Relevance of Health Literacy to mHealth. Information Services & Use 37, 2 (Jan. 2017), 123–130. https://doi.org/10.3233/ISU-170828
@@
 https://doi.org/10.1145/2750858.2804288
+- Le et al.
+(2024) Ha Le, Rithika Lakshminarayanan, Jixin Li, Varun Mishra, and Stephen Intille.
+2024.
+Collecting Self-reported Physical Activity and Posture Data Using Audio-based Ecological Momentary Assessment.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 3, Article 111 (Sept.
+2024), 35 pages.
+https://doi.org/10.1145/3678584
 - Lee et al.
-(2024a) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee.
+(2024a) Hyunsoo Lee, Yugyeong Jung, Hei Yiu Law, Seolyeong Bae, and Uichin Lee.
 2024a.
+PriviAware: Exploring Data Visualization and Dynamic Privacy Control Support for Data Collection in Mobile Sensing Research.
+In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24).
+Association for Computing Machinery, New York, NY, USA, Article 787, 17 pages.
+https://doi.org/10.1145/3613904.3642815
+- Lee et al.
+(2024b) Hyunsoo Lee, Yugyeong Jung, Youwon Shin, Hyesoo Park, Woohyeok Choi, and Uichin Lee.
+2024b.
 FamilyScope: Visualizing Affective Aspects of Family Social Interactions Using Passive Sensor Data.
@@
 - Lee et al.
-(2024b) Jarrett G.W.
+(2024c) Jarrett G.W.
 Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe.
-2024b.
+2024c.
+Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 2, Article 51 (May 2024), 21 pages.
+https://doi.org/10.1145/3659585
+- Lee et al.
+(2024d) Jarrett G.W.
+Lee, Bongshin Lee, Soyoung Choi, JooYoung Seo, and Eun Kyoung Choe.
+2024d.
 Identify, Adapt, Persist: The Journey of Blind Individuals with Personal Health Technologies.
@@
 https://doi.org/10.1097/CIN.0000000000000854
+- Li et al.
+(2010) Ian Li, Anind Dey, and Jodi Forlizzi.
+2010.
+A stage-based model of personal informatics systems.
+In Proceedings of the SIGCHI Conference on Human Factors in Computing Systems (Atlanta, Georgia, USA) (CHI ’10).
+Association for Computing Machinery, New York, NY, USA, 557–566.
+https://doi.org/10.1145/1753326.1753409
+- Li et al.
+(2021) Qingyang Li, Clara Caldeira, Daniel A.
+Epstein, and Yunan Chen.
+2021.
+Supporting Caring among Intergenerational Family Members through Family Fitness Tracking.
+In Proceedings of the 14th EAI International Conference on Pervasive Computing Technologies for Healthcare (Atlanta, GA, USA) (PervasiveHealth ’20).
+Association for Computing Machinery, New York, NY, USA, 1–10.
+https://doi.org/10.1145/3421937.3422018
 - Lieber et al.
@@
 https://doi.org/10.1016/j.ssmph.2020.100572
+- Lin et al.
+(2025) Georgianna Lin, Minh Ngoc Le, Khai N.
+Truong, and Alex Mariakakis.
+2025.
+The Cognitive Strategies Behind Multimodal Health Sensemaking: A Menstrual Health Tracking Case Study.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+9, 3, Article 110 (Sept.
+2025), 27 pages.
+https://doi.org/10.1145/3749482
+- Lin et al.
+(2024) Georgianna Lin, Brenna Li, Jin Yi Li, Chloe Zhao, Khai Truong, and Alexander Mariakakis.
+2024.
+Users’ Perspectives on Multimodal Menstrual Tracking Using Consumer Health Devices.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 3, Article 116 (Sept.
+2024), 24 pages.
+https://doi.org/10.1145/3678575
+- Loerakker et al.
+(2024) Meagan B.
+Loerakker, Jasmin Niess, Marit Bentvelzen, and Paweł W.
+Woźniak.
+2024.
+Designing Data Visualisations for Self-Compassion in Personal Informatics.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+7, 4, Article 169 (Jan.
+2024), 22 pages.
+https://doi.org/10.1145/3631448
+- Lu et al.
+(2020) Lin Lu, Jiayao Zhang, Yi Xie, Fei Gao, Song Xu, Xinghuo Wu, and Zhewei Ye.
+2020.
+Wearable Health Devices in Health Care: Narrative Systematic Review.
+JMIR Mhealth Uhealth 8, 11 (9 Nov 2020), e18907.
+https://doi.org/10.2196/18907
 - Lu et al.
@@
 https://doi.org/10.2196/jmir.6349
+- Mandl et al.
+(2020) Kenneth D Mandl, Daniel Gottlieb, Joshua C Mandel, Vladimir Ignatov, Raheel Sayeed, Grahame Grieve, James Jones, Alyssa Ellis, and Adam Culbertson.
+2020.
+Push button population health: the SMART/HL7 FHIR bulk data access application programming interface.
+NPJ digital medicine 3, 1 (2020), 151.
 - Meyers et al.
@@
 - MoHFW (2005) MoHFW. 2005. School Health Programme.
+- Moore et al.
+(2022) Jimmy Moore, Pascal Goffin, Jason Wiese, and Miriah Meyer.
+2022.
+An Interview Method for Engaging Personal Data.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+5, 4, Article 173 (Dec.
+2022), 28 pages.
+https://doi.org/10.1145/3494964
+- Motahar and Wiese (2022) Tamanna Motahar and Jason Wiese.
+2022.
+A Review of Personal Informatics Research for People with Motor Disabilities.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+6, 2, Article 66 (July 2022), 31 pages.
+https://doi.org/10.1145/3534614
 - Murnane et al.
@@
 https://doi.org/10.1145/3589462.3589496
+- Niess et al.
+(2021) Jasmin Niess, Paweł W.
+Woźniak, Yomna Abdelrahman, Passant ElAgroudy, Yasmeen Abdrabou, Caroline Eckerth, Sarah Diefenbach, and Kristina Knaving.
+2021.
+‘I Don’t Need a Goal’: Attitudes and Practices in Fitness Tracking beyond WEIRD User Groups.
+In Proceedings of the 23rd International Conference on Mobile Human-Computer Interaction (Toulouse & Virtual, France) (MobileHCI ’21).
+Association for Computing Machinery, New York, NY, USA, Article 15, 14 pages.
+https://doi.org/10.1145/3447526.3472062
 - Ning et al.
@@
 https://doi.org/10.1145/3637348
+- Orzikulova et al.
+(2024) Adiba Orzikulova, Han Xiao, Zhipeng Li, Yukang Yan, Yuntao Wang, Yuanchun Shi, Marzyeh Ghassemi, Sung-Ju Lee, Anind K Dey, and Xuhai Xu.
+2024.
+Time2Stop: Adaptive and Explainable Human-AI Loop for Smartphone Overuse Intervention.
+In Proceedings of the 2024 CHI Conference on Human Factors in Computing Systems (Honolulu, HI, USA) (CHI ’24).
+Association for Computing Machinery, New York, NY, USA, Article 250, 20 pages.
+https://doi.org/10.1145/3613904.3642747
 - Pai and Alathur (2020) Rajesh R.
@@
 https://doi.org/10.1145/3389189.3397989
+- Park and Lee (2023) Joonyoung Park and Uichin Lee.
+2023.
+Understanding Disengagement in Just-in-Time Mobile Health Interventions.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+7, 2, Article 72 (June 2023), 27 pages.
+https://doi.org/10.1145/3596240
 - Pasi (2024) Suyash Pasi.
@@
 https://doi.org/10.3390/jcm11237019
-- Poole (2013) Erika S. Poole. 2013. HCI and Mobile Health Interventions. Translational Behavioral Medicine 3, 4 (Dec. 2013), 402–405. https://doi.org/10.1007/s13142-013-0214-3
+- Pina et al.
+(2017) Laura R.
+Pina, Sang-Wha Sien, Teresa Ward, Jason C.
+Yip, Sean A.
+Munson, James Fogarty, and Julie A.
+Kientz.
+2017.
+From Personal Informatics to Family Informatics: Understanding Family Practices around Health Monitoring.
+In Proceedings of the 2017 ACM Conference on Computer Supported Cooperative Work and Social Computing (Portland, Oregon, USA) (CSCW ’17).
+Association for Computing Machinery, New York, NY, USA, 2300–2315.
+https://doi.org/10.1145/2998181.2998362
 - Pulimeno et al.
@@
 https://doi.org/10.34172/hpp.2020.50
+- Raj et al.
+(2019) Shriti Raj, Joyce M.
+Lee, Ashley Garrity, and Mark W.
+Newman.
+2019.
+Clinical Data in Context: Towards Sensemaking Tools for Interpreting Personal Health Data.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+3, 1, Article 22 (March 2019), 20 pages.
+https://doi.org/10.1145/3314409
 - Rapp and Cena (2016) Amon Rapp and Federica Cena.
@@
 https://doi.org/10.1038/s41591-024-03180-7
+- Rey et al.
+(2023) Bradley Rey, Bongshin Lee, Eun Kyoung Choe, and Pourang Irani.
+2023.
+Investigating In-Situ Personal Health Data Queries on Smartwatches.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+6, 4, Article 179 (Jan.
+2023), 19 pages.
+https://doi.org/10.1145/3569481
+- Robinson et al.
+(2015) Charlotte Robinson, Clara Mancini, Janet van der Linden, Claire Guest, and Lydia Swanson.
+2015.
+Exploring assistive technology for assistance dog owners in emergency situations.
+In Proceedings of the 8th ACM International Conference on PErvasive Technologies Related to Assistive Environments (Corfu, Greece) (PETRA ’15).
+Association for Computing Machinery, New York, NY, USA, Article 92, 2 pages.
+https://doi.org/10.1145/2769493.2769576
 - Rooksby et al.
@@
 https://doi.org/10.3390/s23239498
+- Shen et al.
+(2025) Vivian Shen, Xiaoying Yang, Chris Harrison, and Yang Zhang.
+2025.
+Hapt-Aids: Self-Powered, On-Body Haptics for Activity Monitoring.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+9, 3, Article 129 (Sept.
+2025), 26 pages.
+https://doi.org/10.1145/3749468
 - Shin et al.
@@
 https://doi.org/10.1145/3585088.3593922
+- Sullivan and Lachman (2017) Alycia N Sullivan and Margie E Lachman.
+2017.
+Behavior change with fitness technology in sedentary adults: a review of the evidence for increasing physical activity.
+Frontiers in public health 4, Article 289 (2017), 16 pages.
 - Sulz et al.
@@
 https://doi.org/10.1145/3308561.3354628
+- Wang et al.
+(2024) Jianyang Wang, Dongheng Zhang, Binbin Zhang, Jinbo Chen, Yang Hu, and Yan Chen.
+2024.
+RF-GymCare: Introducing Respiratory Prior for RF Sensing in Gym Environments.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+8, 3, Article 135 (Sept.
+2024), 28 pages.
+https://doi.org/10.1145/3678568
 - Wani et al.
@@
 - Wisetsri and Vijai (2021) Worakamol Wisetsri and C. Vijai. 2021. Rise of Artificial Intelligence in Healthcare Startups in India. Advances in Management 14, 1 (March 2021), 48–52.
+- Wolf (2019) Christine T.
+Wolf.
+2019.
+Explainability scenarios: towards scenario-based XAI design.
+In Proceedings of the 24th International Conference on Intelligent User Interfaces (Marina del Ray, California) (IUI ’19).
+Association for Computing Machinery, New York, NY, USA, 252–257.
+https://doi.org/10.1145/3301275.3302317
 - Wozniak-O’Connor et al.
@@
 https://doi.org/10.1145/3638380.3638436
+- Yang et al.
+(2015) Rayoung Yang, Eunice Shin, Mark W.
+Newman, and Mark S.
+Ackerman.
+2015.
+When fitness trackers don’t ’fit’: end-user difficulties in the assessment of personal tracking device accuracy.
+In Proceedings of the 2015 ACM International Joint Conference on Pervasive and Ubiquitous Computing (Osaka, Japan) (UbiComp ’15).
+Association for Computing Machinery, New York, NY, USA, 623–634.
+https://doi.org/10.1145/2750858.2804269
+- Yfantidou et al.
+(2023) Sofia Yfantidou, Pavlos Sermpezis, Athena Vakali, and Ricardo Baeza-Yates.
+2023.
+Uncovering Bias in Personal Informatics.
+Proc.
+ACM Interact.
+Mob.
+Wearable Ubiquitous Technol.
+7, 3, Article 139 (Sept.
+2023), 30 pages.
+https://doi.org/10.1145/3610914
 - Zaher et al.
@@
 https://doi.org/10.1145/3613905.3650907
+Appendix A Thematic Analysis Coding Chart
+| ††: Thematic Analysis table. | ||
+| Themes | Sub-Theme/Code | Design implication |
+| Barriers to Proactivity | Ecosystem Fragmentation & Inter-op Problems Definition: Friction from multiple non-interoperable apps/brands and scattered records preventing longitudinal PHI use.
+Evidence (source): Interviews.
+| Design interoperable dashboards, import/export (EHR) connectors, and a single-view aggregator; prioritize Android-first flows based on market share.
+|
+| Low Digital Health Literacy Definition: Users cannot interpret metrics or know what action to take from raw analytics.
+Evidence (source): Field notes and Interviews.
+| Provide literacy-appropriate threshold-based plain-language recommendations and visual metaphors.
+| |
+| Trust & Sensor-Confidence Issues Definition: Doubts about the accuracy of advanced metrics; users cross-check multiple sources before acting.
+Evidence (source): Interviews and Co-design workshops | Add sensor-confidence indicators, provenance badges, cross-device comparison views, and clinician-validated thresholds.
+| |
+| Device & Form-factor Friction Definition: Physical discomfort, charging burden, and undesired form factors cause abandonment.
+Evidence (source): Interviews | Support multiple modalities (bands, rings, app-only), low-maintenance devices, battery-friendly modes, and non-wearable capture options.
+| |
+| Preference for Embodied/Traditional Cues & Reactive Norms Definition: Reliance on body-feeling and doctor visits; belief that reactive care is sufficient.
+Evidence (source): Interviews and Co-design workshops.
+| Design a hybrid onboarding that maps PHI outputs to familiar embodied cues; provide low-friction “check only when alerted” modes.
+| |
+| Enablers of Proactivity: Behaviours & Triggers | Stakeholder specific sharing Definition: PHI used as a shared family resource; younger family members’ recommendations create derived trust; Linking PHI to clinicians enables earlier action.
+Evidence (source): Interviews and health-sharing survey: 114 respondents — 20 users, 54 non-users, 40 unaware.
+| Create a user-friendly family-sharing UI with approval/request features.
+Include one-page PDF/HL7-style exports, customizable clinician summaries, and time-based trend highlights.
+|
+| Motivational Social Features Definition: Family challenges, badges, and friendly competition that encourage preventive routines.
+Evidence (source): Co-design workshops.
+| Add consensual family challenges, non-punitive badges, group goals, and shareable progress snapshots that respect privacy controls.
+| |
+| Actionable Explanations (literacy-appropriate) Definition: Plain-language recommendations tied to metric thresholds (what to do, not just numbers).
+Evidence (source): Interviews and Co-design workshops.
+| Convert numbers actions: “If X over Y days, do Z”; contextual links to vetted local resources; in-app “explain this metric” with examples.
+| |
+| Proactive Monitoring (behavior) Definition: Regular, routine checking of PHI and prevention-oriented actions (daily/weekly review, goal setting, early consultations).
+Evidence (source): 55/87 (60%) survey participants reported tracking with a wearable; interviews | Quantify proactive vs reactive segments; surface summarized trends (daily/weekly) and goal dashboards; surface family-monitoring affordances.
+| |
+| Triggers for Proactivity Definition: Specific events, alerts, family prompts or clinician requests that cause earlier action (alerts, family messages, prototype “ask to share”).
+Evidence (source): Co-design workshops | Implement configurable alerts, “ask-to-share” nudges, permissioned push messages, and simple export/share buttons for clinician workflows.
+|